### v2

Abstract
In recent years, the holographic duality between $T\bar{T}$-deformed conformal field theory (CFT) and Anti-de Sitter (AdS) space with finite radial truncation has received significant attention. The study of $T\bar{T}$ deformation within the framework of de Sitter (dS)/CFT duality has also progressed. This paper shows that the trace flow equation in dS spacetime can be analytically extended from its AdS counterpart through double Wick rotations. Additionally, we generalize the replica sphere method in both AdS and dS holography to derive a general expression for the entanglement entropy of arbitrary single spatial intervals. For both finite size and finite temperature systems, we obtain the entanglement entropy corrections due to $T\bar{T}$ deformation. Finally, using strong subadditivity and boosted strong subadditivity, we show that the dual field theory on the timelike boundary is non-local in dS holography.
Keywords:
Entanglement entropy, $T\bar{T}$ deformation, AdS holography, dS holography.
1 Introduction
The $T\bar{T}$ deformation Smirnov:2016lqw ; Cavaglia:2016oda of conformal field theory (CFT) is an irrelevant deformation that exhibits factorization properties in the large central charge limit Zamolodchikov:2004ce ; Jiang:2019tcq . For a two-dimensional CFT, the $T\bar{T}$ deformation is defined by the double trace operator
$\frac{dS}{d\lambda}=2\pi\int d^{2}x\sqrt{g}~{}T\bar{T},\quad T\bar{T}\equiv\frac{1}{8}(T^{ij}T_{ij}-(T_{i}^{i})^{2}).$ (1)
The strength of this deformation is controlled by the deformation parameter $\lambda$. If we treat the deformation as a perturbation and expand the deformed theory to first-order in the deformation parameter $\lambda$, we obtain
$S_{\mathrm{QFT}}\simeq S_{\mathrm{CFT}}+2\pi\lambda\int d^{2}x\sqrt{g}~{}T\bar{T}.$ (2)
From the perspective of the AdS/CFT duality Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj , the $T\bar{T}$-deformed CFT corresponds to the semi-classical gravity theory living in an AdS spacetime with finite cutoff McGough:2016lol ; Kraus:2018xrn , as shown in figure 1.
Figure 1: The holography $T\bar{T}$ deformation. The black solid circle represents the AdS boundary, the black dashed circle represents the cutoff at the asymptotic boundary $r=\frac{1}{\epsilon}$, and the red dashed circle represents the finite cutoff $r=r_{c}$.
In refs. McGough:2016lol ; Kraus:2018xrn , the authors calculated the propagation speed of the massless modes, energy spectrum, and partition functions in both the contexts of the $T\bar{T}$-deformed quantum field theory and AdS gravity with a finite cutoff. The generic holographic correlation functions have been established in the refs. He:2023hoj ; He:2023knl ; He:2024fdm ; He:2024xbi . These results support the duality depicted in figure 1. In addition to the aforementioned correspondences, another significant result of holographic principles is that the entanglement entropy of the boundary field theory equals the area of the minimal surface, also known as the Ryu-Takayanagi (RT) surface, inside the AdS spacetime Hubeny:2007xt ; Ryu:2006bv . The RT formula for holographic entanglement entropy serves as a criterion for validating the holographic principle, with $T\bar{T}$-deformed theories offering an additional test. For the holographic duality with $T\bar{T}$ deformation, this duality has been discussed in refs. Chen:2018eqk ; Jeong:2019ylz ; Jiang:2019epa ; Allameh:2021moy ; Setare:2022qls ; He:2023xnb ; Apolo:2023ckr . In ref. Chen:2018eqk , the authors investigated the entanglement entropy of the $T\bar{T}$-deformed field theory at the level of perturbation (2). They found that for the system at finite temperature, the entanglement entropy receives a first-order $\lambda$ correction from the $T\bar{T}$ deformation, but for a system of finite size, this first-order correction disappears. To explore the impact of $T\bar{T}$ deformation on the entanglement entropy in finite-size systems, ref. Jiang:2023ffu examined the timelike entanglement entropy, drawing similar conclusions to previous studies Li:2022tsv ; Doi:2023zaf . They found that the timelike entanglement entropy receives no first-order correction at finite temperature, but there is a first-order correction in the finite size case. Additionally, the reflected entropy Dutta:2019gen and pseudo entropy Nakata:2020luh of the $T\bar{T}$ deformed field theory have been studied in refs. Basu:2024bal ; He:2023wko . For more interesting work on $T\bar{T}$ deformation, please refer to refs. He:2019vzf ; Asrat:2020uib ; Khoeini-Moghaddam:2020ymm ; Li:2020pwa ; Jeong:2022jmp ; He:2023obo ; Tian:2023fgf ; Apolo:2023vnm ; Basu:2023bov ; Basu:2023aqz ; He:2024pbp ; Tsolakidis:2024wut ; Babaei-Aghbolagh:2024hti .
The holographic principle has been well established in the AdS spacetime with finite cutoff $r=r_{c}$. However, given that our real universe is more accurately described by a dS metric, it is crucial to explore whether the holographic principle extends to the dS spacetime Witten:2001kn ; Maldacena:2002vr ; Miyaji:2015yva ; Jiang:2023loq ; Chang:2023gkt ; Chen:2023prz ; Doi:2024nty . In the usual dS/CFT duality Strominger:2001pn ; Strominger:2001gp , the CFT resides on a future spacelike boundary, which renders the dual field theory non-unitary Maldacena:2002vr ; Doi:2022iyj . Due to the challenges of the dS/CFT duality Anninos:2011ui ; Ng:2012xp , there are only a few clearly established examples Hikida:2022ltr ; Hikida:2021ese ; Chen:2022ozy . In recent years, other holographic duality models in the dS spacetime, such as the dS/dS correspondence Alishahiha:2004md ; Alishahiha:2005dj ; Dong:2018cuv ; Geng:2019bnn ; Geng:2019ruz ; Geng:2020kxh ; Geng:2021wcq and half-dS holography Kawamoto:2023nki , have gained attention and been studied. In these two models, the holographic boundary is a timelike boundary of the dS spacetime, so the dual field theory is expected to be unitary. In the dS/dS correspondence, the cosmological constant term also appears in the trace flow equation of the $T\bar{T}$ deformation, thus the $T\bar{T}$ deformation of this model is referred to as $T\bar{T}+\Lambda_{2}$ deformation Gorbenko:2018oov ; Shyam:2021ciy ; Coleman:2021nor ; Torroba:2022jrk ; Batra:2024kjl .
We aim to study the impact of $T\bar{T}$ deformation on the entanglement entropy in the dS/CFT duality. However, the deformed field theory no longer retains conformal symmetry, rendering the original method for calculating the entanglement entropy inapplicable. Fortunately, when the deformed CFT is situated on a two-dimensional sphere, the entanglement entropy of a subregion with endpoints at antipodal points can be calculated via the replica method Calabrese:2009qy ; Donnelly:2018bef ; Caputa:2019pam ; Deng:2023pjs . In this context, by adjusting the spherical radius $r$, we can study the renormalization group flow Wilson:1973jj ; Polchinski:1983gv of the dual field theory under $T\bar{T}$ deformation, with $r$ functioning as the radial coordinate in the bulk spacetime. This method is an extension of the approach proposed in ref. Lewkowycz:2013nqa and was further generalized in ref. Lewkowycz:2019xse to compute the entanglement entropy of arbitrary spatial intervals on a two-dimensional plane. In ref. Chen:2023eic , the replica method was employed within the $T\bar{T}$-deformed dS/CFT duality to calculate the entanglement entropy for the subregion between two antipodal points on the sphere.
In this paper, we present a general expression for the entanglement entropy of arbitrary spatial intervals in both AdS and dS holography by combining the Casini-Huerta-Myers (CHM) map Casini:2011kv with the replica method on a sphere. The results are
  • AdS($0<r_{\text{eff}}<\infty$): $S=\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\text{eff}}^{2}}}}\right),$
  • dS($\ell_{\rm dS}<r_{\text{eff}}<\infty$): $S=-\frac{ic_{\rm dS}}{3}\mathrm{arccoth}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6},$
  • dS($0<r_{\text{eff}}<\ell_{\rm dS}$): $S=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}-1}}\right),$
where $\ell_{(\rm{A})dS}$ represents the (A)dS radius, $c_{\rm(A)dS}=\frac{3\ell_{(\rm{A})dS}}{2G}$ is the Brown-Henneaux central charge Brown1986 , $\lambda_{\rm(A)dS}=8G\ell_{(\rm{A})dS}$ is the $T\bar{T}$ deformation parameter [^1] As mentioned in ref. Donnelly:2018bef , we adopt the same unit system. The factor $r_{c}$, which distinguishes the CFT metric from the induced bulk metric, is absorbed into the effective radius $r_{\text{eff}}$ of the sphere. and $r_{\text{eff}}$ denotes the effective radius of the sphere, which is related to the size of the subsystem. In the case of dS holography, we find that the results for the holographic entanglement entropy can be summarized in a single formula:
$S=-\frac{ic_{\rm dS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}}}\right)\,.$ (3)
When the effective radius of the sphere is greater than the cosmological horizon radius ($\ell_{\rm dS}<r_{\text{eff}}<\infty$), this formula corresponds to the result of the $T\bar{T}$-deformed dS/CFT duality, where the dual field theory is non-unitary. When the effective radius of the sphere is less than the cosmological horizon radius ($0<r_{\text{eff}}<\ell_{\rm dS}$), this formula agrees with the results of dS/dS correspondence and half-dS holography, where the dual field theory is non-local Geng:2019ruz ; Geng:2020kxh ; Kawamoto:2023nki .
This paper is organized as follows: In Section 2, we review the results from ref. Kraus:2018xrn to derive the trace flow equations in AdS/CFT with Euclidean signature and in dS/CFT with Lorentzian signature. We then demonstrate the extension from one result to the other via a double Wick rotation. In Section 3, we review the calculation of the entanglement entropy of the $T\bar{T}$-deformed CFT, providing general expressions for both AdS and dS holographies. Based on these findings, we calculate the $C$-function and discuss the non-unitarity and non-locality of the dual field theory. In Section 4, we review and generalize the CHM method for calculating the entanglement entropy of arbitrary spatial intervals in the $T\bar{T}$-deformed theory. We explore the effects of the $T\bar{T}$ deformation in finite temperature and finite size systems. In Section 5, we first verify the calculation of the entanglement entropy in the AdS/dS correspondence. We then examine the non-locality of the dual field theory in dS/dS correspondence and half-dS holography by analyzing the strong subadditivity and the boosted strong subadditivity. Finally, we summarize our results and present a discussion in Section 6.
2 Holographic $T\bar{T}$ Flow
In this section, we review the holographic derivation of the trace flow equation for the $T\bar{T}$ deformation Shyam:2017znq ; Hartman:2018tkw . We then show that the trace flow equation in the dS/CFT duality can be obtained from its AdS counterpart via a double Wick rotation.
2.1 $T\bar{T}$ Flow in AdS/CFT
The action for Einstein gravity in Euclidean AdS3 (EAdS3) spacetime $\mathcal{M}$, with a boundary $\partial\mathcal{M}$ at a finite radial cutoff, is given by
$S=\frac{1}{16\pi G}\int_{\mathcal{M}}d^{3}x\sqrt{g}\left(\mathcal{R}^{(3)}+\frac{2}{\ell_{\rm AdS}^{2}}\right)+\frac{1}{8\pi G}\int_{\partial\mathcal{M}}d^{2}x\sqrt{\gamma}\left(K-\frac{1}{\ell_{\rm AdS}}\right),$ (4)
where $K$ denotes the trace of the extrinsic curvature of the timelike boundary $\partial\mathcal{M}$. The quasi-local stress tensor Balasubramanian:1999re is defined as the variation of the on-shell action with respect to the induced boundary metric $\gamma^{ij}$,
$T_{ij}=\frac{2}{\sqrt{\gamma}}\frac{\delta S}{\delta\gamma^{ij}}=\frac{1}{8\pi G}\left(K_{ij}-K\gamma_{ij}+\frac{1}{\ell_{\rm AdS}}\gamma_{ij}\right).$ (5)
The EAdS3 metric in Fefferman-Graham Fefferman1985 coordinates is
$ds^{2}=\frac{\ell_{\rm AdS}^{2}}{r^{2}}\left(dr^{2}+h_{ij}dx^{i}dx^{j}\right),\quad\gamma_{ij}=\frac{\ell_{\rm AdS}^{2}}{r_{c}^{2}}h_{ij}.$ (6)
In these coordinates, the spacetime is foliated by hypersurfaces labeled by the radial coordinate $r$. The extrinsic curvature of the hypersurface $\Sigma_{r_{c}}$ at $r=r_{c}$ is given by
$K_{ij}=\frac{1}{2}n^{r}\partial_{r}\left.\left(\frac{\ell_{\rm AdS}^{2}}{r^{2}}h_{ij}\right)\right|_{r=r_{c}}=\frac{r}{2\ell_{\rm AdS}}\partial_{r}\left.\left(\frac{\ell_{\rm AdS}^{2}}{r^{2}}h_{ij}\right)\right|_{r=r_{c}},$ (7)
where $n^{r}$ is the outward-pointing normal vector to the hypersurface $\Sigma_{r_{c}}$. The $(r,r)$ component of Einstein’s equations is
$K^{2}-K_{ij}K^{ij}=\mathcal{R}^{(2)}+\frac{2}{\ell_{\rm AdS}^{2}},$ (8)
where $\mathcal{R}^{(2)}$ is the Ricci scalar of the hypersurface $\Sigma_{r_{c}}$. Substituting the expression for $K_{ij}$ from eq. (5) into eq. (8) yields the trace flow equation
${T^{i}}_{i}=-\frac{\ell_{\rm AdS}}{16\pi G}\mathcal{R}^{(2)}-4\pi G\ell_{\rm AdS}\left[T^{ij}T_{ij}-({T^{i}}_{i})^{2}\right].$ (9)
Equation (8) is equivalent to the Hamiltonian constraint $H=0$ in the canonical form of Einstein’s gravity theory Arnowitt:1962hi , with the Hamiltonian given by
$H=\frac{16\pi G}{\sqrt{\gamma}}\left[\Pi^{ij}\Pi_{ij}-({\Pi^{i}}_{i})^{2}\right]+\frac{\sqrt{\gamma}}{16\pi G}\left(\mathcal{R}^{(2)}+\frac{2}{\ell_{\rm AdS}^{2}}\right)=0,$ (10)
where $\Pi^{ij}$ is the canonical momentum defined as
$\Pi^{ij}=\frac{\sqrt{\gamma}}{16\pi G}(K\gamma^{ij}-K^{ij}).$ (11)
In the canonical form of quantum gravity with Euclidean signature, the wavefunction of the induced metric $\gamma_{ij}$ on the hypersurface $\Sigma_{r_{c}}$ can be written as $\Psi\equiv\text{e}^{-S}$, where $S$ is the classical action of Einstein gravity with the Gibbons-Hawking-York boundary term. After renormalization by introducing a counterterm, the wavefunction takes the form Witten:2022xxp
$\Psi=\exp\left(-\frac{1}{8\pi G\ell_{\rm AdS}}\int_{\partial\mathcal{M}}d^{2}x\sqrt{\gamma}\right)\hat{\Psi}.$ (12)
This corresponds to performing a canonical transformation on the conjugate momentum
$\Pi^{ij}\to\Pi^{ij}+\frac{\sqrt{\gamma}}{16\pi G\ell_{\rm AdS}}\gamma^{ij}.$ (13)
The Wheeler-DeWitt equation $H\hat{\Psi}=0$ DeWitt:1967yk in terms of the original conjugate momentum becomes
$H\hat{\Psi}$ $=\left(\frac{16\pi G}{\sqrt{\gamma}}(\Pi^{ij}\Pi_{ij}-({\Pi^{i}}_{i})^{2})-\frac{2}{\ell_{\rm AdS}}\Pi+\frac{\sqrt{\gamma}}{16\pi G}\mathcal{R}^{(2)}\right)\hat{\Psi},$ (14)
$=\sqrt{\gamma}\left(4\pi G(T^{ij}T_{ij}-({T^{i}}_{i})^{2})+\frac{1}{\ell_{\rm AdS}}{T^{i}}_{i}+\frac{\mathcal{R}^{(2)}}{16\pi G}\right)\hat{\Psi}=0,$
where in the second line we have used the relation $T^{ij}=-\frac{2}{\sqrt{\gamma}}\Pi^{ij}$. This aligns with the trace flow equation (9). To match this with the CFT trace flow equation
${T^{i}}_{i}=-\frac{c}{24\pi}\mathcal{R}^{(2)}-4\pi\lambda T\bar{T},$ (15)
we require the central charge $c$ and the deformation parameter $\lambda$ to be
$\text{AdS}:\quad c=\frac{3\ell_{\rm AdS}}{2G}=c_{\rm AdS},\quad\lambda=8G\ell_{\rm AdS}=\lambda_{\rm AdS}.$ (16)
2.2 $T\bar{T}$ Flow in dS/CFT
In the dS/CFT duality, the holographic direction corresponds to time, so the dS metric in Fefferman-Graham coordinates is given by
$ds^{2}=\frac{\ell_{\rm dS}^{2}}{r^{2}}(-dr^{2}+h_{ij}dx^{i}dx^{j}),\quad\gamma_{ij}=\frac{\ell_{\rm dS}^{2}}{r_{c}^{2}}h_{ij},$ (17)
where $r$ is the holographic coordinate. The holographic boundary $\Sigma_{r_{c}}$, determined by $r=r_{c}$, is spacelike. The bulk geometry has a Lorentzian signature, and its gravitational partition function is assumed to be the Bunch-Davies wavefunction Hikida:2022ltr ; Doi:2022iyj . Thus, according to the holographic dictionary, the partition function of the CFT is
$\Psi_{\text{dS}}=Z_{\text{QFT}}=e^{iS_{\rm dS}},$ (18)
where $S_{\rm dS}$ is the on-shell action of Einstein gravity with the dS geometry,
$S_{\rm dS}=\frac{1}{16\pi G}\int_{\mathcal{M}}d^{3}x\sqrt{-g}\left(\mathcal{R}^{(3)}-\frac{2}{\ell_{\rm dS}^{2}}\right)+\frac{1}{8\pi G}\int_{\partial\mathcal{M}}d^{2}x\sqrt{\gamma}\left(K-\frac{1}{\ell_{\rm dS}}\right).$ (19)
The trace flow equation for the dS/CFT case can be derived using a method parallel to that of the AdS/CFT case. First, we define the renormalized wavefunction by introducing boundary counterterms, corresponding to a canonical transformation on the conjugate momentum
$\Psi_{\rm dS}=\exp\left(\frac{i}{8\pi G\ell_{\rm dS}}\int_{\partial\mathcal{M}}d^{2}x\sqrt{\gamma}\right)\hat{\Psi}_{\rm dS},\quad\Pi^{ij}_{\rm dS}\to\Pi^{ij}_{\rm dS}+\frac{\sqrt{\gamma}}{16\pi G\ell_{\rm dS}}\gamma^{ij}.$ (20)
The Wheeler-DeWitt equation for the renormalized wavefunction, expressed in terms of the original conjugate momentum, is
$H\hat{\Psi}_{\rm dS}$ $=\left[\frac{16\pi G}{\sqrt{\gamma}}(\Pi^{ij}\Pi_{ij}-({\Pi^{i}}_{i})^{2})-\frac{2}{\ell_{\rm dS}}\Pi-\frac{\sqrt{\gamma}}{16\pi G}\mathcal{R}^{(2)}\right]\hat{\Psi}_{\rm dS}$ (21)
$=\sqrt{\gamma}\left(-4\pi G(T^{ij}T_{ij}-({T^{i}}_{i})^{2})-\frac{i}{\ell_{\rm dS}}{T^{i}}_{i}-\frac{\mathcal{R}^{(2)}}{16\pi G}\right)\hat{\Psi}_{\rm dS}=0,$
where $T^{ij}=-\frac{2i}{\sqrt{\gamma}}\Pi^{ij}$. This leads to the $T\bar{T}$ trace flow equation for the dS/CFT is
${T^{i}}_{i}=\frac{i\ell_{\rm dS}}{16\pi G}\mathcal{R}^{(2)}+i4\pi G\ell_{\rm dS}\left(T^{ij}T_{ij}-({T^{i}}_{i})^{2}\right).$ (22)
By comparing this with the field theory equation, the correspondences between the Brown-Henneaux central charge, the deformation parameter, and the gravitational quantities for the dS spacetime Chen:2023eic are given by
$c=-i\frac{3\ell_{\rm AdS}}{2G}=-ic_{\rm dS}\,,\quad\lambda=-i8G\ell_{\rm AdS}=-i\lambda_{\rm dS}\,.$ (23)
2.3 Double Wick Rotation to dS/CFT
In this subsection, we demonstrate that the trace flow equation for the dS/CFT, given by eq. (22), can be derived by performing a double Wick rotation on the trace flow equation for the AdS/CFT, given by eq. (9). The double Wick rotation is defined by
$\ell_{\rm AdS}\to-i\ell_{\rm dS},\quad r\to-ir.$ (24)
The transformations of various components in the partition function under the double Wick rotation are summarized in Table 1222Note that for the general dS/CFT, where the CFT resides on the infinite spacelike boundary, the external curvature must be defined using the inward normal vector..
Table 1: Analytical continuation from EAdS to dS
EAdS dS Analytical Continuation
$n_{a}^{(\text{AdS})}=\frac{\ell_{\rm AdS}}{r}(dr)_{a}$ $n_{a}^{(\text{dS})}=\frac{\ell_{\rm dS}}{r}(dr)_{a}$ $n_{a}^{(\text{AdS})}\to-in_{a}^{(\text{dS})}$
$n^{a}_{(\text{AdS})}=\frac{r}{\ell_{\rm AdS}}\left(\frac{\partial}{\partial r}\right)^{a}$ $n^{a}_{(\text{dS})}=-\frac{r}{\ell_{\rm dS}}\left(\frac{\partial}{\partial r}\right)^{a}$ $n^{a}_{(\text{AdS})}\to-in^{a}_{(\text{dS})}$
$K_{ab}^{(\text{AdS})}=\frac{1}{2}n^{r}\partial_{r}g_{ab}$ $K_{ab}^{(\text{dS})}=-\frac{1}{2}n^{r}\partial_{r}g_{ab}$ $K_{ab}^{(\text{AdS})}\to iK_{ab}^{(\text{dS})}$
$\epsilon_{abc}^{(\text{AdS})}=\sqrt{g}\,(dr)_{a}\wedge(d\tau)_{b}\wedge(dx)_{c}$ $\epsilon_{abc}^{(\text{dS})}=\sqrt{-g}\,(dr)_{a}\wedge(d\tau)_{b}\wedge(dx)_{c}$ $\epsilon_{abc}^{(\text{AdS})}\to-i\epsilon_{abc}^{(\text{dS})}$
$\epsilon_{ab}^{(\text{AdS})}=n^{a}_{(\text{AdS})}\epsilon_{abc}^{(\text{AdS})}$ $\epsilon_{ab}^{(\text{dS})}=n^{a}_{(\text{dS})}\epsilon_{abc}^{(\text{dS})}$ $\epsilon_{ab}^{(\text{AdS})}\to-\epsilon_{ab}^{(\text{dS})}$
$\mathcal{R}_{(\text{AdS})}^{(3)}=-\frac{6}{\ell_{\rm AdS}^{2}}$ $\mathcal{R}_{(\text{dS})}^{(3)}=\frac{6}{\ell_{\rm dS}^{2}}$ $\mathcal{R}_{(\text{AdS})}^{(3)}\to\mathcal{R}_{(\text{dS})}^{(3)}$
$S_{ct}^{(\text{AdS})}$ $S_{ct}^{(\text{dS})}$ $S_{ct}^{(\text{AdS})}\to-iS_{ct}^{(\text{dS})}$
By applying the transformation rules in Table 1 to the on-shell action of AdS gravity (4), we obtain the on-shell action for dS gravity, multiplied by an imaginary factor $i$:
$S_{\rm AdS}$ $\rightarrow-\frac{i}{16\pi G}\int_{\mathcal{M}}d^{3}x\sqrt{-g}\left(\mathcal{R}^{(3)}-\frac{2}{\ell_{\rm dS}^{2}}\right)$
$\quad-\frac{1}{8\pi G}\int_{\partial\mathcal{M}}d^{2}x\sqrt{\gamma}\left(iK-\frac{i}{\ell_{\rm dS}}\right)$
$=-iS_{\rm dS}.$ (25)
The double Wick rotation implies that the partition function of AdS gravity corresponds to the Bunch-Davies wavefunction of dS gravity. Additionally, all quantities in the Wheeler-DeWitt equation for the AdS spacetime are mapped to their corresponding quantities in the Wheeler-DeWitt equation for the dS spacetime. As a result, eqs. (12) and (14) correspond to eqs. (20) and (21), respectively. Therefore, the trace flow equation in the dS/CFT can be directly derived by applying this rotation to the trace flow equation in the AdS/CFT.
3 The Replica Method for the $T\bar{T}$-Deformed Field Theory
This section reviews the replica trick for calculating the entanglement entropy of the subregion between two antipodal points on a two-dimensional sphere Donnelly:2018bef ; Chen:2023eic . The sphere serves as the background manifold for the $T\bar{T}$-deformed CFT, with its radius linked to the deformation parameter from a holographic perspective. In the dS/CFT correspondence, the CFT resides on a future spacelike boundary, which is also a two-dimensional sphere in global dS coordinates. Thus, the replica method applied in the AdS case is equally applicable in the dS case.
The metric that describes the boundary two-dimensional sphere in both AdS and dS cases is
$ds^{2}=\gamma_{ab}dx^{a}dx^{b}=r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right),$ (26)
where $r$ represents the effective radius of the boundary sphere. The replica trick is used to compute the entanglement entropy of the subregion $A$, defined between the antipodal points $\theta=0$ and $\theta=\pi$ at $\phi=\text{constant}$, as illustrated in figure 2.
Figure 2: The meridian at $\phi=\text{constant}$ with endpoints at antipodal points defines the subregion $A$.
By performing $n$ copies along the meridian at $\phi=\text{constant}$, the $n$-replica geometry is obtained Dong:2023bax ; Lewkowycz:2013nqa by
$ds_{n}^{2}=r^{2}\left(d\theta^{2}+n^{2}\sin^{2}\theta\,d\phi^{2}\right).$ (27)
The entanglement entropy of the subregion $A$ is computed by taking the limit $n\rightarrow 1$ of the conical entropy Headrick:2010zt ; Hung:2011nu ; Lewkowycz:2013nqa
$S_{A}=\lim_{n\to 1}\left(1-n\frac{\partial}{\partial n}\right)\ln Z_{n},$ (28)
where $Z_{n}$ is the partition function of the field theory on the $n$-replica geometry. Based on the definition of the stress tensor, the following expressions hold
$\frac{d}{dr}\ln Z_{n}$ $=-\int_{\mathbb{S}_{n}^{2}}d^{2}x\frac{\delta\mathcal{L}}{\delta g_{ab}}\frac{dg_{ab}}{dr}=-\frac{1}{r}\int_{\mathbb{S}_{n}^{2}}d^{2}x\sqrt{g}T^{ab}g_{ab},$ (29)
and
$\frac{\partial}{\partial n}\ln Z_{n}$ $=-\int_{\mathbb{S}_{n}^{2}}d^{2}x\frac{\delta\mathcal{L}}{\delta g_{\phi\phi}}\frac{\partial g_{\phi\phi}}{\partial n}=-\frac{1}{n}\int_{\mathbb{S}_{n}^{2}}d^{2}x\sqrt{g}T^{\phi\phi}g_{\phi\phi}.$ (30)
The trace flow equation and the conservation equation determine the stress tensor as follows:
$\begin{cases}{T^{a}}_{a}=-\frac{c}{24\pi}R^{(2)}-\frac{\pi\lambda}{2}\left[T^{ab}T_{ab}-({T^{a}}_{a})^{2}\right],\\ \nabla_{a}{T^{a}}_{b}=0.\end{cases}$ (31)
Due to the maximal symmetry of the geometry, the quasi-local stress tensor is assumed to take the form
$T_{ab}=\alpha\gamma_{ab},\quad{T^{\theta}}_{\theta}={T^{\phi}}_{\phi}=\frac{1}{2}\text{tr}\,T.$ (32)
Under these assumptions, it follows that $\frac{d\ln Z_{n}}{dr}\big{|}_{n=1}=-8\pi r\alpha$, and the entanglement entropy of the subregion $A$ is given by
$S_{A}=\lim_{n\to 1}\left(1-\frac{r}{2}\frac{d}{dr}\right)\ln Z_{n}.$ (33)
In this case, the parameter $\alpha$ determines the stress tensor, which is uniquely defined by the trace flow equations. This approach is specifically designed to calculate the entanglement entropy of the subregion between two antipodal points on the sphere. Future work aims to extend this framework to explore broader implications of the holographic principle, building on the refs. Grieninger:2019zts ; Lewkowycz:2019xse .
3.1 AdS case
By substituting the assumed form of the stress tensor $T_{ij}=\alpha\gamma_{ij}$ into eq. (9), the equation for $\alpha$ is obtained as
$8\pi G\ell_{\rm AdS}\alpha^{2}=2\alpha+\frac{\ell_{\rm AdS}}{16\pi G}\mathcal{R}^{(2)},$ (34)
where $\mathcal{R}^{(2)}=\frac{2}{r^{2}}$ for the two-dimensional sphere (26). The solution to this equation is given by
$\alpha=\frac{1}{\pi\lambda_{\rm AdS}}\left(1-\sqrt{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r^{2}}}\right).$ (35)
Substituting this expression for $\alpha$ into eq. (29), the $n$-replica partition function $Z_{n}$ can be computed by integrating with respect to $\alpha$:
$\ln Z_{n}=\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\right)+\frac{4r^{2}}{\lambda_{\rm AdS}}\left(-1+\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}\right).$ (36)
The entanglement entropy of the subregion $A$ is then obtained as
$S_{A}=\lim_{n\to 1}\left(1-\frac{r}{2}\frac{d}{dr}\right)\log Z_{n}=\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\right).$ (37)
The entanglement entropy between two antipodal points on a sphere reflects the number of degrees of freedom in the field theory defined on the sphere Casini:2012ei . Casini and Huerta, using the principles of unitarity and locality in quantum field theory (QFT), demonstrated that the $C$-function, which measures the degrees of freedom, exhibits a monotonically decreasing behavior along the renormalization group flow for unitary and Lorentz-invariant QFTs Casini:2004bw ; Barnes:2004jj ; Gukov:2015qea . This validates Zamolodchikov’s $C$-theorem Zamolodchikov:1986gt . Since the $T\bar{T}$ trace flow equation is equivalent to the holographic renormalization group flow equation, the renormalization group flow can be controlled by adjusting the radius of the sphere deBoer:1999tgo ; Verlinde:1999xm ; Heemskerk:2010hk ; Nishioka:2018khk ; Grieninger:2023knz ; Kamei:2024iei . In the AdS case, the $C$-function is given by
$C(r)=3rS_{A}^{\prime}(r)=c_{\rm AdS}\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\,.$ (38)
As illustrated in figure 3, the $C$-function decreases monotonically as $r$ decreases, consistent with the number of degrees of freedom decreases along the renormalization group flow. Additionally, as $r$ approaches infinity, the $C$-function converges to the central charge of the two-dimensional CFT, indicating that the renormalization group flow approaches the UV fixed point Nakayama:2013is . This behavior is in agreement with the $C$-theorem.
(a) $C(r)$
(b) $C(\lambda)$
Figure 3: The $C$-function of the AdS spacetime. We use $\lambda=\frac{\lambda_{\rm AdS}}{r^{2}}$ to represent the deformation parameter. As the boundary cutoff $r=r_{c}$ is pushed further into the bulk, the deformation parameter increases and the theory deviates from the CFT increasingly.
The $C$-theorem can be proven using locality and unitarity, but the converse is not true. Therefore, the locality of the field theory is next examined through the strong subadditivity inequality and the boosted strong subadditivity inequality. The strong subadditivity inequality is given by
$S(A)+S(B)-S(A\cup B)-S(A\cap B)\geq 0,$ (39)
where the subsystems $A$ and $B$ are shown in figure 4(a). In the limit where $A\rightarrow B$ and $|A|=|B|=L$, the strong subadditivity inequality can be expressed in its infinitesimal form as
$L^{2}S^{\prime\prime}(L)\leq 0,$ (40)
where $L$ represents the size of the subsystems.
(a) strong subadditivity
(b) boosted strong subadditivity
Figure 4: The left figure shows two spatial regions $A$ and $B$ at $t=0$. The right figure considers the regions $A$ and $B$ after a boost, resulting in two spacelike regions.
A stronger criterion for proving locality, compared to the strong subadditivity, is the boosted strong subadditivity, which can be derived by applying a boost transformation to subsystems $A$ and $B$, as shown in figure 4(b). The infinitesimal form of the boosted strong subadditivity inequality is expressed as
$L^{2}S^{\prime\prime}(L)+LS^{\prime}(L)\leq 0.$ (41)
Since the subsystems correspond to great arcs between antipodal points on a sphere, their size is proportional to the sphere’s radius, i.e., $L=\pi r$. Consequently, the derivatives with respect to $L$ in the strong subadditivity and the boosted strong subadditivity inequalities are equivalent to derivatives with respect to $r$. Taking the second derivative of eq. (37) with respect to $r$ yields the strong subadditivity inequality
$r^{2}\partial^{2}_{r}S_{A}(r)=-\frac{c_{\rm AdS}}{3\left(1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}\right)^{\frac{3}{2}}}<0.$ (42)
This result shows that the strong subadditivity inequality is always satisfied in the AdS case, while the boosted strong subadditivity inequality is violated. This violation is consistent within the framework of $T\bar{T}$-deformed theories, where the boundary field theory resides on a finitely truncated surface. In such scenarios, the entanglement wedge of the boosted subsystem $A$ is not fully contained within the entanglement wedge of $A\cup B$ Lewkowycz:2019xse .
3.2 dS case
In the context of the dS spacetime, there are two primary holographic models: dS/CFT and dS/dS. The focus here is on the dS/dS correspondence in Euclidean signature, where the holographic boundary is a 2-dimensional sphere. The metric of the dS spacetime in global coordinates is given by
$ds^{2}=\ell_{\rm dS}^{2}\left[-dt^{2}+\cosh^{2}t\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right)\right],$ (43)
where the radius of the boundary sphere is $r=\ell_{\rm dS}\cosh t$, which exceeds $\ell_{\rm dS}$ for any finite cutoff at $t=t_{c}$. In dS/dS coordinates, the metric of the dS spacetime is
$ds^{2}=\ell_{\rm dS}^{2}\left[dw^{2}+\sin^{2}w\left(-d\tau^{2}+\cosh^{2}\tau\,d\phi^{2}\right)\right],$ (44)
where the effective sphere radius is $r=\ell_{\rm dS}\sin w<\ell_{\rm dS}$. The point $r=\ell_{\rm dS}$ corresponds to the cosmological horizon of the dS spacetime. When $r>\ell_{\rm dS}$, the hypersurface $r=r_{c}$ is spacelike, which aligns with the standard dS/CFT duality. Conversely, when $r<\ell_{\rm dS}$, the hypersurface $r=r_{c}$ becomes timelike, corresponding to the dS/dS correspondence, as illustrated in figure 5.
Figure 5: The Penrose diagram for the dS spacetime. The orange solid line represents future infinity, while the orange dashed line represents the finite cutoff at $t=t_{c}$. The red squares represent the boundaries at $w=0,\pi$ in dS/dS coordinates. The blue solid line represents $w=\pi/2$, and the blue dashed line represents the finite cutoff at $w=w_{c}$.
For the dS/CFT duality, the assumed form of the quasi-local stress tensor (32) is substituted into eq. (22) to determine the value of $\alpha$:
$8\pi G\ell_{\rm dS}\alpha^{2}=2i\alpha+\frac{\ell_{\rm dS}}{16\pi G}\mathcal{R}^{(2)},\quad\alpha=\frac{i}{\pi\lambda_{\rm dS}}\left(1-\sqrt{1-\frac{\lambda_{\rm dS}c_{\rm dS}}{12r^{2}}}\right).$ (45)
By integrating $\alpha$ with respect to $r$, the $n$-replica partition function is obtained as
$\ln Z_{n}=-\frac{ic_{\rm dS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)+\frac{i4r^{2}}{\lambda_{\rm dS}}\left(-1+\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}\right).$ (46)
The entanglement entropy between of $A$ is then given by
$S_{A}=\lim_{n\to 1}\left(1-\frac{r}{2}\frac{d}{dr}\right)\ln Z_{n}=-\frac{ic_{\rm dS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right).$ (47)
In this case, the argument of the inverse hyperbolic function is greater than one, making $\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)$ a complex value. The positive branch of the inverse hyperbolic function is chosen to match the real part of the entanglement entropy in the dS/CFT. The final expression for the entanglement entropy in the dS/CFT correspondence is
$S_{A}=-\frac{ic_{\rm dS}}{3}\mathrm{arccoth}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6},$ (48)
where the real part agrees with the results in refs. Hikida:2022ltr ; Doi:2023zaf ; Chen:2023eic . As discussed in ref. Doi:2022iyj , the entanglement entropy in the dS/CFT can be interpreted as the pseudo entropy, which is viewed as the origin of the non-unitarity of the boundary field theory in the dS/CFT. Additionally, we still want to use the $C$-function to measure the degree of freedom of the field theory. However, due to the non-unitarity of the dual field theory, the $C$-function is also a purely imaginary number.
$C(r)=3rS_{A}^{\prime}(r)=-ic_{\rm dS}\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\,.$ (49)
For the undeformed CFT, we have $C(r\to\infty)=-ic_{\rm dS}=c$. But when $r\to\ell_{\rm dS}$, the imaginary part of the $C$-function diverges, as shown in figure 6.
(a) $C(r)$
(b) $C(\lambda)$
Figure 6: The $C$-function of the dS spacetime at $r>\ell_{\rm dS}$, and $\lambda=\frac{\lambda_{\rm dS}}{r^{2}}$. The red dashed line represents $r=\ell_{\text{dS}}$.
In the dS/dS case, the entanglement entropy and $C$-function have been analyzed in refs. Gorbenko:2018oov ; Dong:2018cuv , with the results given as
$S_{A}=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}\right),\quad C(r)=3rS_{A}^{\prime}(r)=c_{\rm dS}\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}.$ (50)
It is noted that when the effective radius is expressed as $r=\ell_{\rm dS}\sin w<\ell_{\rm dS}$, these results can be directly obtained through analytic continuation from the corresponding entanglement entropy and $C$-function in the dS/CFT correspondence. The $C$-function does decrease monotonically as $r$ decreases, and it still diverges at $r=\ell_{\rm dS}$, as shown in figure 7.
(a) $C(r)$
(b) $C(\lambda)$
Figure 7: The $C$-function of the dS spacetime at $r<\ell_{\rm dS}$.
By calculating the second derivative of the entanglement entropy with respect to the sphere radius $r$, it can be found that when the dual field theory resides on a timelike boundary, the entanglement entropy violates the strong subadditivity inequality,
$r^{2}\partial^{2}_{r}S_{A}(r)=\frac{1}{\left(\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1\right)^{\frac{3}{2}}}>0.$ (51)
Therefore, it implies that the dual field theory is non-local Kawamoto:2023nki ; Geng:2019ruz when the effective radius of the sphere is smaller than the cosmological horizon.
4 Entanglement entropy for arbitrary single spatial intervals
In section 3, the replica method is employed to calculate the entanglement entropy for a subsystem between two antipodal points on a sphere under $T\bar{T}$ deformation. This method applies only to subsystems along semicircular arcs between antipodal points. To analyze the entanglement entropy for subsystems in arbitrary single spatial intervals on a two-dimensional plane, ref. Deng:2023pjs introduced the CHM mapping, which transforms the subsystem into a semicircular arc on a two-dimensional sphere. The entanglement entropy can then be determined using the results obtained for the spherical case. For subsystems on more general manifolds, they can first be mapped onto the two-dimensional plane, after which the method from ref. Deng:2023pjs can be applied. The replica method is generalized in the following section to calculate the entanglement entropy for arbitrary single spatial intervals on the two-dimensional plane with $T\bar{T}$ deformation and extended to general manifolds, including cylinders and spheres.
4.1 The generalization of the replica method
The goal is to calculate the entanglement entropy of the $T\bar{T}$-deformed quantum field theory in a spatial subregion $A$ on an arbitrary two-dimensional manifold $\mathcal{M}$, where $A$ is defined by $t=0$ and $x\in[x_{a},x_{b}]$ in a chosen coordinate system. When $\mathcal{M}$ is a two-dimensional sphere and $A$ is specified by $\phi=\text{constant}$ and $\theta\in(0,\pi)$, the entanglement entropy can be computed using eqs. (29) and (30).
The entanglement entropy of $A$ can be obtained through the $C$-function. Differentiating $S_{A}$ with respect to $r$, we have
$r\frac{d}{dr}S_{A}(r)$ $=r\frac{d}{dr}\lim_{n\rightarrow 1}(1-n\partial_{n})\ln Z_{n}=\lim_{n\rightarrow 1}(1-n\partial_{n})r\frac{d}{dr}\ln Z_{n}$
$=-\lim_{n\rightarrow 1}(1-n\partial_{n})2\pi nr^{2}\int_{0}^{\pi}\sin\theta d\theta\,\text{tr}\,T$
$=2\pi r^{2}\lim_{n\rightarrow 1}n^{2}\int_{0}^{\pi}\sin\theta d\theta\,\partial_{n}\text{tr}\,T,$ (52)
where, in the last two steps, it is assumed that the metric (27) is independent of the azimuthal angle $\phi$, and thus $T_{ab}$ is also independent of $\phi$. Solving the trace flow and conservation equations without assuming spherical symmetry must determine the energy-momentum tensor $T_{ab}$. For the $n$-replica sphere (27), the solution for the energy-momentum tensor is known Lewkowycz:2019xse and is given by
${T^{\theta}}_{\theta}$ $=\frac{1}{\pi\lambda}\left(1-\sqrt{1+\frac{\lambda c}{12r^{2}}+\frac{\lambda c}{12r^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}\right),$
${T^{\phi}}_{\phi}$ $=\frac{1}{\pi\lambda}\left(1-\frac{1+\frac{\lambda c}{12r^{2}}}{\sqrt{1+\frac{\lambda c}{12r^{2}}+\frac{\lambda c}{12r^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}}\right).$ (53)
Substituting this into the integral in eq. (52), in the limit $n\rightarrow 1$, the integrand becomes
$n\partial_{n}\text{tr}\,T=\frac{c}{3}\frac{\lambda c}{48\pi}\frac{1-n^{2}}{n^{4}r^{4}\sin^{4}\theta}\frac{1}{\left[1+\frac{\lambda c}{12r^{2}}\left(1+\frac{2(1-n)}{\sin^{2}\theta}\right)\right]^{3/2}},$ (54)
which tends to zero for $\theta\neq 0,\pi$. However, near the branch points, the integrand has a non-zero contribution. Thus, the main contribution to the integral in eq. (52) comes from small neighborhoods near the branch points
$rS^{\prime}(r)=2\pi r^{2}\lim_{n\rightarrow 1}\left(\int_{\theta\in B(0,\epsilon)}+\int_{\theta\in B(\pi,\epsilon)}\right)\sin\theta d\theta\,\partial_{n}\text{tr}\,T,$ (55)
where $B(p,\epsilon)$ is a small neighborhood centered at $p$ with radius $\epsilon$.
For a general manifold $\mathcal{M}$, we first map $\mathcal{M}$ to $\mathbb{S}^{2}$, ensuring that the image of the subsystem $A$ under this mapping is a meridian on $\mathbb{S}^{2}$. The mapped metric is conformally equivalent to the sphere, denoted as $\Omega\mathbb{S}^{2}$. After applying the $n$-replica trick, we assume the metric takes the form
$ds^{2}=r^{2}\Omega^{2}(\theta,\phi)\left(d\theta^{2}+n^{2}\sin^{2}\theta\,d\phi^{2}\right),$ (56)
where the conformal factor $\Omega(\theta,\phi)$ is constant at the branch points $\theta=0$ and $\theta=\pi$, denoted as $\Omega_{0}$ and $\Omega_{\pi}$, respectively. For this metric, eqs. (29) and (30) remain valid
$\frac{d}{dr}\ln Z_{n}$ $=-\int_{\Omega\mathbb{S}_{n}^{2}}d^{2}x\frac{\delta\mathcal{L}}{\delta g_{ab}}\frac{dg_{ab}}{dr}=-\frac{1}{r}\int_{\Omega\mathbb{S}_{n}^{2}}d^{2}x\sqrt{g}T^{ab}g_{ab}.$ (57)
Thus, we have
$r\frac{d}{dr}S(r)$ $=r\frac{d}{dr}\lim_{n\rightarrow 1}(1-n\partial_{n})\ln Z_{n}=\lim_{n\rightarrow 1}(1-n\partial_{n})r\frac{d}{dr}\ln Z_{n}$
$=-\lim_{n\rightarrow 1}(1-n\partial_{n})nr^{2}\int_{\Omega\mathbb{S}_{n}^{2}}\Omega^{2}(\theta,\phi)\sin\theta\,d\theta d\phi\,\text{tr}\,T$
$=r^{2}\lim_{n\rightarrow 1}n^{2}\int_{\Omega\mathbb{S}_{n}^{2}}\Omega^{2}(\theta,\phi)\sin\theta\,d\theta d\phi\,\partial_{n}\text{tr}\,T.$ (58)
Since the conformal factor $\Omega(\theta,\phi)$ depends on the azimuthal angle $\phi$, $T_{ab}$ cannot be assumed to depend solely on $\theta$. Therefore, the $\phi$ coordinate is not integrated in the final two steps. In standard CFT calculations, a conformal transformation is often applied to map $\Omega\mathbb{S}^{2}$ to the spherical metric, and conformal symmetry ensures that the results for $\Omega\mathbb{S}^{2}$ and $\mathbb{S}^{2}$ differ only by a function of the conformal factor. However, for the $T\bar{T}$-deformed field theory, conformal symmetry is absent, and the relationship between quantities before and after the conformal transformation is not determined. As discussed in ref. Lewkowycz:2019xse , for the partition function on the sphere (52), in the limit $n\rightarrow 1$, the main contribution to the integral (52) comes from a small neighborhood near the branch points, with radius $\epsilon\sim\sqrt{n-1}$. This property is expected to hold for any manifold $\mathcal{M}$, so the integral (58) is also performed only in small neighborhoods near the branch points. In these regions, the conformal factor $\Omega$ is constant, and the metric (56) can be approximated by the $n$-replica sphere with radius $r\Omega$. As a result, the conformal transformation can be regarded as an identity transformation. Therefore, partition functions on any manifold $\mathcal{M}$ can be calculated using the results on $\mathbb{S}^{2}$ for non-conformal field theories. Finally, we can get
$rS^{\prime}(r)$ $=-\lim_{n\rightarrow 1}(1-n\partial_{n})\int_{\Omega\mathbb{S}_{n}^{2}}d^{2}x\sqrt{g}\,\text{tr}\,T$
$=\lim_{n\rightarrow 1}2\pi(r\Omega)^{2}\left(\int_{0}^{\theta_{0}}+\int_{\pi-\theta_{0}}^{\pi}\right)\sin\theta\,d\theta\,\partial_{n}\text{tr}\,T,$ (59)
where $\text{tr}\,T$ corresponds to the solution on the sphere with radius $r\Omega$.
When Poincaré coordinates are chosen in EAdS3, the holographic boundary corresponds to a two-dimensional complex plane with the metric
$ds_{2}^{2}=\frac{\ell_{\rm AdS}^{2}}{z_{c}^{2}}(dT^{2}+dX^{2}).$ (60)
The subregion $A$ is defined by $T=0$ and $X\in(-X_{a},X_{a})$. To compute the entanglement entropy for $A$, the complex plane is mapped onto a two-dimensional sphere such that the image of $A$ corresponds to the meridian on the sphere at $\phi=0$, as illustrated in figure 8. This mapping, known as the CHM mapping, is given by
$X=\frac{X_{a}\cos\theta}{1+\sin\theta\cos\phi},\quad T=\frac{X_{a}\sin\theta\sin\phi}{1+\sin\theta\cos\phi}.$ (61)
Under this transformation, the boundary plane in $(\theta,\phi)$ coordinates has the metric
$ds_{2}^{2}=\frac{\ell_{\rm AdS}^{2}X_{a}^{2}}{z_{c}^{2}(1+\cos\phi\sin\theta)^{2}}(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}),$ (62)
which is conformally equivalent to a sphere with an effective radius $r_{\mathrm{eff}}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}$, and the conformal factor $\Omega(\theta,\phi)=(1+\cos\phi\sin\theta)^{-1}$.
Figure 8: Mapping from the plane to the sphere. The north and south poles of the sphere correspond to the endpoints of the interval $A$ on the complex plane. The subregion $A$ corresponds to the meridian at $\phi=0$, while the locally time-evolved subregions (shown as blue or yellow lines) correspond to other meridians on the sphere.
At this stage, the energy-momentum tensor for the metric (62) must be determined using the equations in (31), followed by evaluating the integral in (52). As previously noted, when computing the partition function on a general manifold $\mathcal{M}$, the primary contribution to the integral (52) comes from an infinitesimal neighborhood near the branch points. In these regions, the conformal factor $\Omega(\theta,\phi)$ can be approximated as 1, implying that the metric (62) effectively describes a sphere with radius $r_{\mathrm{eff}}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}$. Thus, within the infinitesimal neighborhoods around the endpoints, transforming (60) to the sphere involves only a conformal transformation with a conformal factor of 1, effectively an identity transformation. Consequently, the partition function remains the same as that of the sphere, with the radius $r$ replaced by the effective radius $r_{\mathrm{eff}}$.
For the AdS/CFT, the deformation parameter $\lambda$ and the central charge $c$ in eq. (31) are $c_{\rm AdS}=\frac{3\ell_{\rm AdS}}{2G}$ and $\lambda_{\rm AdS}=8G\ell_{\rm AdS}$, respectively. Therefore, the solution for the energy-momentum tensor on the $n$-replica sphere is
${T^{\theta}}_{\theta}$ $=\frac{1}{\pi\lambda_{\rm AdS}}\left(1-\sqrt{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}\right),$
${T^{\phi}}_{\phi}$ $=\frac{1}{\pi\lambda_{\rm AdS}}\left(1-\frac{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}}{\sqrt{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}}\right)\,,$ (63)
where the effective radius is $r_{\rm eff}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}$. Substituting this into eq. (52) and performing the integration near the endpoints yields the $C$-function as
$C(r)=\frac{c_{\rm AdS}}{3}\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\rm eff}^{2}}}}\,.$ (64)
Using the $C$-function, one can get the entanglement entropy of the subregion A as
$S_{A}=\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\rm eff}^{2}}}}\right)\simeq\frac{c_{\mathrm{AdS}}}{3}\log\frac{2X_{a}}{z_{c}}+\frac{c_{\mathrm{AdS}}}{12}\frac{z_{c}^{2}}{X_{a}^{2}}.$ (65)
In this expression, the first term represents the entanglement entropy of the undeformed field theory in the interval $(-X_{a},X_{a})$ on the plane. The second term accounts for the correction introduced by the $T\bar{T}$ deformation. This result is consistent with the findings reported in ref. He:2023xnb .
For the dS/CFT, the deformation parameter $\lambda$ and the central charge $c$ in eq. (31) are given by $c\equiv-ic_{\rm dS}=-i\frac{3\ell_{\rm dS}}{2G}$ and $\lambda\equiv-i\lambda_{\rm dS}=-i8G\ell_{\rm dS}$, respectively. The solution for the energy-momentum tensor on the $n$-replica sphere
${T^{\theta}}_{\theta}$ $=\frac{i}{\pi\lambda_{\rm dS}}\left(1-\sqrt{1-\frac{\lambda_{\rm dS}c_{\rm dS}}{12r_{\rm eff}^{2}}-\frac{\lambda_{\rm dS}c_{\rm dS}}{12r_{\rm eff}^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}\right),$
${T^{\phi}}_{\phi}$ $=\frac{i}{\pi\lambda_{\rm dS}}\left(1-\frac{1-\frac{\lambda_{\rm dS}c_{\rm dS}}{12r_{\rm eff}^{2}}}{\sqrt{1-\frac{\lambda_{\rm dS}c_{\rm dS}}{12r_{\rm eff}^{2}}-\frac{\lambda_{\rm dS}c_{\rm dS}}{12r_{\rm eff}^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}}\right)\,,$ (66)
where the effective radius is $r_{\rm eff}=\frac{\ell_{\rm dS}X_{a}}{\eta_{c}}$. Substituting this into eq. (52), we obtain the $C$-function is
$C(r)=-\frac{ic_{\rm dS}}{3}\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}}}\,.$ (67)
The entanglement entropy is then given by
$S_{A}=-\frac{ic_{\rm dS}}{3}\mathrm{arccoth}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6}\simeq-\frac{ic_{\rm dS}}{3}\log\frac{2X_{a}}{\eta_{c}}+\frac{ic_{\rm dS}\eta_{c}^{2}}{12X_{a}^{2}}+\frac{\pi c_{\rm dS}}{6}\,.$ (68)
In Appendix A.1, this quantity is computed using holographic methods by evaluating the area of the RT surface with a finite cutoff $\eta_{c}$. The result is consistent with eq. (68).
To generalize the method for computing subsystems’ entanglement entropy on arbitrary holographic boundaries, we note that any two-dimensional manifold is locally conformally equivalent to the plane. Thus, its line element can be expressed as
$ds^{2}=\omega(T,X)^{2}(dT^{2}+dX^{2})\,.$ (69)
Suppose the subregion $A$ in the coordinate system $(T,X)$ is a line segment with endpoints at $(T_{a},-X_{a})$ and $(T_{a},X_{a})$. To calculate the entanglement entropy of this subsystem, we map it onto the sphere using the transformation
$T$ $\rightarrow\frac{X_{a}\sin\theta\sin\phi}{1+\sin\theta\cos\phi}+T_{a}\,,\quad X\rightarrow\frac{X_{a}\cos\theta}{1+\sin\theta\cos\phi}\,.$ (70)
This mapping sends the endpoints $(T_{a},-X_{a})$ and $(T_{a},X_{a})$ on the plane to the antipodal points on the sphere. The line element (69) in these new coordinates becomes
$ds^{2}=\frac{\omega(T(\theta,\phi),X(\theta,\phi))^{2}X_{a}^{2}}{(1+\cos\phi\sin\theta)^{2}}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)\,,$ (71)
where $\theta(T_{a},X_{a})=0$ and $\theta(T_{a},-X_{a})=\pi$. We then perform a conformal transformation to eliminate the conformal factor $\frac{\omega(T(\theta,\phi),X(\theta,\phi))}{1+\cos\phi\sin\theta}$, resulting in a sphere with an effective radius $r_{\rm eff}$. If $\omega(T(\theta,\phi),X(\theta,\phi))$ is constant at both endpoints, this factor only contributes to the effective radius. At the endpoints $\theta=0$ and $\pi$, the factor $1+\cos\phi\sin\theta$ is 1. Thus, near these endpoints, the metric (71) approximates that of a sphere, with an effective radius of $\omega(T_{a},X_{a})X_{a}$. Once again, assuming that the integrand in (52) is non-zero only in small neighborhoods near the endpoints, the conformal transformation is the identity map in these neighborhoods. Thus, we can directly use the partition function on the sphere with radius $r_{\rm eff}$ as the partition function on the manifold $\mathcal{M}$ within the subregion $A$. Within these small neighborhoods, we can use the solution for the energy-momentum tensor on the $n$-sphere to compute the integral, with the effective radius in the solutions (63) or (66) replaced by $r_{\mathrm{eff}}=\omega(T_{a},X_{a})X_{a}$.
If $\omega(T,X)$ takes different constant values at the two endpoints, denoted as $\omega_{a}$ and $\omega_{-a}$, respectively. The contributions to the integral (52) from different endpoints correspond to spheres with radii $\omega_{a}X_{a}$ and $\omega_{-a}X_{a}$. Therefore, these integrals should be replaced by the partition function on spheres with radii $\omega_{a}X_{a}$ and $\omega_{-a}X_{a}$, respectively.
4.2 System with finite size
In the previous subsection, we reviewed the CHM mapping method for calculating the entanglement entropy of the $T\bar{T}$-deformed CFT in spatial intervals on a two-dimensional plane. Specifically, we computed the entanglement entropy for a subsystem on the boundary of the dS spacetime in Poincaré coordinates. We now extend this method to two-dimensional cylindrical and spherical surfaces to examine whether the $T\bar{T}$ deformation introduces corrections to the entanglement entropy for finite-size systems. The corresponding holographic models for these geometries are AdS and dS spacetimes in global coordinates, respectively. The entanglement entropy calculation for general spatial intervals in curved space has been discussed in ref. Grieninger:2019zts .
4.2.1 AdS case
The metric of EAdS3 in global coordinates is
$ds_{\mathrm{EAdS}_{3}}^{2}=\ell_{\rm AdS}^{2}\left(d\rho^{2}+\cosh^{2}\rho\,dt^{2}+\sinh^{2}\rho\,dx^{2}\right)\,,$ (72)
where $x\sim x+2\pi$. At the boundary $\rho=\rho_{c}$, we can get the induced metric as
$ds_{\mathrm{cyl}}^{2}=\ell_{\rm AdS}^{2}\sinh^{2}\rho_{c}\left(d\tilde{t}^{2}+dx^{2}\right)\,,$ (73)
where $\tilde{t}=\frac{\cosh\rho_{c}}{\sinh\rho_{c}}t$. To compute the entanglement entropy of a subsystem in the subregion $A$ defined by the interval $(-x_{a},x_{a})$ on the time slice $t=0$, we first map the cylinder onto the complex plane using the following transformations
$\tilde{t}\rightarrow\frac{1}{2}\ln(T^{2}+X^{2})\,,\quad x\rightarrow\frac{1}{2i}\ln\left(\frac{T+iX}{T-iX}\right)\,.$ (74)
Under this transformation, the endpoints $(0,-x_{a})$ and $(0,x_{a})$ of $A$ are mapped to $(T_{a},-X_{a})=(\cos x_{a},-\sin x_{a})$ and $(T_{a},X_{a})=(\cos x_{a},\sin x_{a})$, respectively. The line element in this coordinate becomes
$ds^{2}_{\mathrm{pl}}=\ell_{\mathrm{AdS}}^{2}\sinh^{2}\rho_{c}\frac{dT^{2}+dX^{2}}{T^{2}+X^{2}}\,.$ (75)
Next, we apply the CHM transformation to map the plane onto a two-dimensional sphere, with the coordinate transformation
$T\rightarrow\frac{\sin x_{a}\sin\theta\sin\phi}{1+\sin\theta\cos\phi}+\cos x_{a}\,,\quad X\rightarrow\frac{\sin x_{a}\cos\theta}{1+\sin\theta\cos\phi}\,.$ (76)
This transformation maps the endpoints of the interval on the plane to antipodal points on the sphere. In these new coordinates, the metric becomes
$ds^{2}_{\Omega\mathbb{S}^{2}}$ $=\frac{\ell_{\rm AdS}^{2}\sinh^{2}\rho_{c}\sin^{2}x_{a}}{(1+\cos(2x_{a}-\phi)\sin\theta)(1+\cos\phi\sin\theta)}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$
$=\Omega^{2}(\theta,\phi)r_{\rm eff}^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)\,,$ (77)
where the effective radius is $r_{\mathrm{eff}}=\ell_{\mathrm{AdS}}\sinh\rho_{c}\sin x_{a}$, and the conformal factor is $\Omega^{-2}(\theta,\phi)=(1+\cos(2x_{a}-\phi)\sin\theta)(1+\cos\phi\sin\theta)$. The mapping from the cylinder to the sphere is illustrated in figure 9.
Figure 9: The CHM transformation for the AdS case. After applying two mappings, an arc on the cylinder with endpoints $(-x_{a},x_{a})$ is mapped to a meridian on the sphere at $\phi=x_{a}$. The meridian is then replicated $n$ times by cutting it along the $\phi$ direction. Using the replica trick, the entanglement entropy of the subregion $A$ is calculated.
At the antipodal points $\theta=0$, $\pi$, the conformal factor simplifies to $\Omega=1$, allowing us to directly compute the entanglement entropy. The resulting entanglement entropy for the subregion $A$ is given by
$S_{A}=$ $\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\rm eff}^{2}}}}\right)$
$\simeq$ $\frac{c_{\rm AdS}}{3}\log(2\sinh\rho_{c}\sin x_{a})+\frac{c_{\rm AdS}}{12\sinh^{2}\rho_{c}\sin^{2}x_{a}}\,,$ (78)
where the first term of the last line represents the entanglement entropy of the undeformed field theory for the interval $(-x_{a},x_{a})$ on the cylinder, with the UV cutoff $\epsilon=\frac{1}{2\sinh\rho_{c}}$. The second term accounts for the contribution from the $T\bar{T}$ deformation. This expression aligns with the perturbative results found in ref. Chen:2018eqk .
4.2.2 dS case
For the global coordinates of the dS spacetime, the metric is given by
$ds^{2}=\ell_{\rm dS}^{2}(-dt^{2}+\cosh^{2}t\,(d\Theta^{2}+\sin^{2}\Theta\,d\Phi^{2}))\,.$ (79)
After the $T\bar{T}$ deformation, its holographic boundary is a 2-dimensional sphere determined by $t=t_{c}$. The subregion $A$ is defined by the arc $\Theta=\pi/2$, $\Phi\in(-\Phi_{a},\Phi_{a})$ on this 2-dimensional sphere. As done previously, we begin by mapping the sphere to the complex plane using the transformation
$\Theta=2\arctan\sqrt{T^{2}+X^{2}}\,,\quad\Phi=\frac{1}{2i}\ln\frac{T+iX}{T-iX}\,.$ (80)
In this coordinate system, the endpoints of $A$ are $(\cos\Phi_{a},-\sin\Phi_{a})$ and $(\cos\Phi_{a},\sin\Phi_{a})$, and the line element becomes
$ds^{2}=\frac{4\ell_{\rm dS}^{2}\cosh^{2}t_{c}}{1+X^{2}+T^{2}}(dT^{2}+dX^{2})\,.$ (81)
As shown in figure 10, after the above mapping, we map an arc on the sphere to an arc on the plane. We then apply the CHM transformation
$T\rightarrow\frac{\sin\Phi_{a}\sin\theta\sin\phi}{1+\sin\theta\cos\phi}+\cos\Phi_{a}\,,\quad X\rightarrow\frac{\sin\Phi_{a}\cos\theta}{1+\sin\theta\cos\phi}\,,$ (82)
which maps this subregion to the meridian $\phi=\Phi_{a}$ on the sphere.
Figure 10: The CHM transformation of the global dS case.
The transformed line element is
$ds^{2}$ $=\frac{\ell_{\rm dS}^{2}\cosh^{2}t_{c}\sin^{2}\Phi_{a}}{(1+\cos(\phi-\Phi_{a})\cos(\Phi_{a})\sin(\theta))^{2}}(d\theta^{2}+\sin^{2}\theta d\phi^{2})$
$=\Omega(\theta,\phi)^{2}r_{\rm eff}^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\,,$ (83)
where $r_{\rm eff}=\ell_{\rm dS}\cosh t_{c}\sin\Phi_{a}$ is the effective radius of the sphere, and the conformal factor is
$\Omega(\theta,\phi)=\frac{1}{1+\cos(\phi-\Phi_{a})\cos(\Phi_{a})\sin(\theta)}\,.$ (84)
At the antipodal points $\theta=0,\pi$, the conformal factor $\Omega(\theta,\phi)=1$, allowing us to directly use the results in eq. (66) to calculate the entanglement entropy. The final result is
$S_{A}$ $=-\frac{ic_{\rm dS}}{3}\mathrm{arccoth}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6}$
$\simeq-\frac{ic_{\rm dS}}{3}\log(2\cosh t_{c}\sin\Phi_{a})+\frac{ic_{\rm dS}}{12\cosh^{2}t_{c}\sin^{2}\Phi_{a}}+\frac{\pi c_{\rm dS}}{6}.$ (85)
This result is consistent with the area of the RT surface calculated in Appendix A.2, and also matches the findings of ref. Chen:2023eic when $\Phi_{a}=\frac{\pi}{2}$.
4.3 System with finite temperature
[…]
$ds^{2}=\frac{R^{2}-R_{0}^{2}}{\ell_{\mathrm{AdS}}^{2}}dt^{2}+\frac{\ell_{\mathrm{AdS}}^{2}}{R^{2}-R_{0}^{2}}dR^{2}+R^{2}dx^{2}\,,$ (86)
where $x$ is the coordinate of the non-compact direction, and $t$ is periodic with $t\sim t+\beta$, where $\beta=\frac{2\pi\ell_{\mathrm{AdS}}^{2}}{R_{0}}$ is the inverse Hawking temperature. The induced metric on the holographic boundary $R=R_{c}$ is
$ds^{2}=\frac{R_{c}^{2}-R_{0}^{2}}{\ell_{\mathrm{AdS}}^{2}}\left(dt^{2}+\frac{R_{c}^{2}\ell_{\mathrm{AdS}}^{2}}{R_{c}^{2}-R_{0}^{2}}dx^{2}\right)\coloneqq\frac{R_{c}^{2}-R_{0}^{2}}{\ell_{\mathrm{AdS}}^{2}}\left(dt^{2}+d\tilde{x}^{2}\right)\,,$ (87)
where $\tilde{x}=\frac{R_{c}\ell_{\mathrm{AdS}}}{\sqrt{R_{c}^{2}-R_{0}^{2}}}x$. We aim to compute the entanglement entropy of the subsystem within the interval $x\in(-x_{a},x_{a})$ at $t=0$. The length of the interval is $L=\frac{2x_{a}R_{c}\ell_{\rm AdS}}{\sqrt{R_{c}^{2}-R_{0}^{2}}}$. To simplify the calculation, we redefine the coordinates by
$R=R_{0}\cosh\rho,\quad t=\frac{\ell_{\rm AdS}^{2}}{R_{0}}\phi=\frac{\beta}{2\pi}\phi,\quad x=\frac{\ell_{\rm AdS}}{R_{0}}\tau$ (88)
to transform the metric (86) to the global coordinate of EAdS3
$ds^{2}=\ell_{\rm AdS}^{2}(\sinh^{2}\rho\,d\phi^{2}+d\rho^{2}+\cosh^{2}\rho\,d\tau^{2}).$ (89)
At this point, the chosen subregion corresponds to the interval $\tau\in(-\tau_{a},\tau_{a})$ at $\phi=0$, where $\tau_{a}=\frac{R_{0}}{\ell_{\rm AdS}}x_{a}$. The holographic boundary $R=R_{c}$ corresponds to $\rho=\rho_{c}$ with $\cosh\rho_{c}=\frac{R_{c}}{R_{0}}$. The induced metric on this boundary is
$ds^{2}=\ell_{\rm AdS}^{2}\cosh^{2}\rho_{c}(d\tau^{2}+d\phi^{2}).$ (90)
Up to this point, the setup remains similar to that in sec 4.2.1. According to eq. (74), the endpoints of the interval in the complex plane are $(T,X)=(e^{-\tau_{a}},0)$ and $(e^{\tau_{a}},0)$, and the line element in this coordinate becomes
$ds^{2}_{\mathrm{pl}}=\ell_{\rm AdS}^{2}\cosh^{2}\rho_{c}\frac{dT^{2}+dX^{2}}{T^{2}+X^{2}}\,.$ (91)
For the interval $X=0$, with $T_{a}\in(e^{-\tau_{a}},e^{\tau_{a}})$, the CHM map takes the form
$X\rightarrow\frac{\sinh\tau_{a}\sin\theta\sin\phi}{1+\sin\theta\cos\phi}\,,\quad T\rightarrow\frac{\sinh\tau_{a}\cos\theta}{1+\sin\theta\cos\phi}+\cosh\tau_{a}\,.$ (92)
Figure 11: The CHM transformation of the BTZ black hole. In this case, the effective radius at the two endpoints is different.
As depicted in figure 11, under the CHM transformation, the interval along the cylindrical axis is mapped to the meridian at $\phi=0$ on the sphere. The resulting metric in these coordinates is given by
$ds^{2}=\frac{\ell_{\rm AdS}^{2}\cosh^{2}\rho_{c}\sinh^{2}\tau_{a}(d\theta^{2}+\sin^{2}\theta d\phi^{2})}{(1+\cos\phi\sin\theta)(\cosh 2\tau_{a}+\cos\phi\sin\theta+\cos\theta\sinh 2\tau_{a})}.$ (93)
In contrast to the previous situation, the effective radius $r_{\rm eff}$ at the endpoints $\theta=0$ and $\theta=\pi$ is not equal. We can regard the constant factor $\ell_{\rm AdS}\cosh\rho_{c}\sinh\tau_{a}$ in the metric (93) as the role of radius “$r$” in eq. (58), with the remaining angle-dependent part serving as the conformal factor $\Omega(\theta,\phi)$. Accordingly, the effective radii corresponding to the two endpoints are
$r_{{\rm eff},\theta=0}$ $=r\Omega(0,\phi)=\ell_{\rm AdS}\cosh\rho_{c}\sinh\tau_{a}e^{-2\tau_{a}}=\ell_{\rm AdS}\frac{R_{c}}{R_{0}}e^{\frac{-2\pi L\sqrt{1-\left(\frac{R_{0}}{R_{c}}\right)^{2}}}{\beta}}\sinh\frac{\pi L\sqrt{1-\left(\frac{R_{0}}{R_{c}}\right)^{2}}}{\beta}\,,$
$r_{{\rm eff},\theta=\pi}$ $=r\Omega(\pi,\phi)=\ell_{\rm AdS}\cosh\rho_{c}\sinh\tau_{a}e^{2\tau_{a}}=\ell_{\rm AdS}\frac{R_{c}}{R_{0}}e^{\frac{2\pi L\sqrt{1-\left(\frac{R_{0}}{R_{c}}\right)^{2}}}{\beta}}\sinh\frac{\pi L\sqrt{1-\left(\frac{R_{0}}{R_{c}}\right)^{2}}}{\beta}\,.$ (94)
As discussed at the end of Section 4.1, we assume that the CHM mapping method is still applicable. However, the energy-momentum tensor in the integrals near different endpoints corresponds to solutions on the sphere with different radii. According to eq. (59), we have
$r\frac{d}{dr}S_{A}(r)$ $=-\lim_{n\rightarrow 1}2\pi\left[r^{2}\Omega^{2}(0,\phi)\int_{0}^{\theta_{0}}\sin\theta d\theta\partial_{n}\mathrm{Tr}T+r^{2}\Omega^{2}(\pi,\phi)\int_{\pi-\theta_{0}}^{\pi}\sin\theta d\theta\partial_{n}\mathrm{Tr}T\right]$
$=\frac{c}{6}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}\Omega^{2}(0,\phi)}}}+\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}\Omega^{2}(\pi,\phi)}}}\right)\,.$ (95)
Integrating the above function with respect to $r$, the entanglement entropy is given by
$S_{A}=\frac{c_{\rm AdS}}{6}\left[\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{{\rm eff},\theta=0}^{2}}}}\right)+\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{{\rm eff},\theta=\pi}^{2}}}}\right)\right]\,.$ (96)
Substituting the radii $r_{{\rm eff},\theta}$ into this expression and expanding the result to second order in $R_{0}/R_{c}$,
$S_{A}$ $\simeq\frac{c_{\rm AdS}}{3}\left[\ln\left(\frac{\beta R_{c}\sinh\frac{\pi L}{\beta}}{\pi\ell_{\rm AdS}^{2}}\right)-\frac{2\pi^{3}\ell_{\rm AdS}^{4}L\coth\left(\frac{\pi L}{\beta}\right)}{R_{c}^{2}\beta^{3}}+\frac{\pi^{2}\ell_{\rm AdS}^{4}\left(1+\coth\left(\frac{2\pi L}{\beta}\right)\right)}{R_{c}^{2}\beta^{2}}\right]\,,$ (97)
where we have replaced $R_{0}$ with $\frac{2\pi\ell_{\rm AdS}^{2}}{\beta}$. In the high-temperature limit $\beta\ll 1$, the last term is negligible compared to the second term. This result is consistent with the findings in refs. Chen:2018eqk ; Apolo:2023ckr by identifying $R_{c}^{2}=\frac{6\ell_{\rm AdS}^{2}}{\mu\pi c_{\rm AdS}}$.
5 (A)dS/dS case and non-locality
In this section, we apply the methods from the previous section to the AdS/dS, dS/dS, and half-dS holography cases. For all three, the holographic boundary is a timelike hypersurface. In dS/dS and half-dS holography, previous studies Kawamoto:2023nki ; Geng:2019ruz suggest that the dual field theory may exhibit non-locality. Unlike the examples discussed earlier, the boundary-induced metrics in this section are time-dependent. As a result, the properties of entanglement entropy may vary significantly over time, potentially influencing the determination of the strong subadditivity or the boosted strong subadditivity. This issue will be discussed in detail in this section.
5.1 AdS/dS
For the AdS/dS correspondence, the bulk metric is expressed as
$ds^{2}=\ell_{\rm AdS}^{2}(dw^{2}+\sinh^{2}w(-dt^{2}+\cosh^{2}tdx^{2})),$ (98)
where $x\sim x+2\pi$, and the holographic boundary is determined by $w=w_{c}\to\infty$. We choose the boundary subregion $A$ as $x\in(-x_{a},x_{a})$ at $t=t_{0}$. In the Euclidean signature, after rescaling the time coordinate, the boundary metric can be written as
$ds^{2}=\ell_{\rm AdS}^{2}\frac{\sinh^{2}w_{c}}{\cosh^{2}\tilde{t}}\left(d\tilde{t}^{2}+dx^{2}\right)\,,$ (99)
where $\tilde{t}$ is related to $t$ by the relation
$\tilde{t}=\ln\frac{\cot\frac{t}{2}+1}{\cot\frac{t}{2}-1}\,.$ (100)
This metric is conformally equivalent to the cylinder. As we have done in the last section, we can map it to the complex plane through the following transformations
$\tilde{t}\rightarrow\frac{1}{2}\ln(T^{2}+X^{2})\,,\quad x\rightarrow\frac{1}{2i}\ln\left(\frac{T+iX}{T-iX}\right)\,.$ (101)
and the resulting conformal flat metric is
$ds^{2}=\ell_{\rm AdS}^{2}\sinh^{2}w_{c}\frac{4\left(dT^{2}+dX^{2}\right)}{\left(1+T^{2}+X^{2}\right)^{2}}\,.$ (102)
The two endpoints of the subregion $A$ in the $(T,X)$ coordinates are
$(T_{1},X_{1})=(e^{\tilde{t}_{0}}\cos x_{a},-e^{\tilde{t}_{0}}\sin x_{a}),\quad(T_{2},X_{2})=(e^{\tilde{t}_{0}}\cos x_{a},e^{\tilde{t}_{0}}\sin x_{a})\,.$ (103)
Using the CHM map,
$T\rightarrow\frac{e^{\tilde{t}_{0}}\sin x_{a}\sin\theta\sin\phi}{1+\sin\theta\cos\phi}+e^{\tilde{t}_{0}}\cos x_{a}\,,\quad X\rightarrow\frac{e^{\tilde{t}_{0}}\sin x_{a}\cos\theta}{1+\sin\theta\cos\phi}\,,$ (104)
this complex plane can be mapped onto the sphere, yielding the conformal spherical metric
$ds^{2}=\frac{4e^{2\tilde{t}_{0}}\ell_{\rm AdS}^{2}\sinh^{2}w_{c}\sin^{2}x_{a}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)}{\left(1+\sin\theta\cos\phi+e^{2\tilde{t}_{0}}\sin\theta\cos(\phi-2x_{a})+e^{2\tilde{t}_{0}}\right)^{2}}\,.$ (105)
It is easy to verify that the conformal factor at the endpoints $\theta=0,\pi$ is the same constant, which can be absorbed into $r_{\rm eff}$. Therefore, the effective radius is
$r_{\text{eff}}=\ell_{\rm AdS}\sinh w_{c}\cos t_{0}\sin x_{a}\,,$ (106)
or, in Lorentz signature, $r_{\text{eff}}=\ell_{\rm AdS}\sinh w_{c}\cosh t_{0}\sin x_{a}$. Finally, the entanglement entropy of the subsystem is
$S_{A}\simeq\frac{c_{\rm AdS}}{3}\log(2\sinh w_{c}\cosh t_{0}\sin x_{a})+\frac{c_{\rm AdS}}{12\sinh^{2}w_{c}\cosh^{2}t_{0}\sin^{2}x_{a}}\,.$ (107)
If we choose the UV cutoff as $\sinh w_{c}=1/\epsilon$, we find that eq.(107) agrees with the result in ref. Kawamoto:2023nki . Alternatively, if we select the subregion as $t=t_{0}$, $x\in[-\frac{\pi}{2},\frac{\pi}{2}]$, the entanglement entropy is $S\simeq\frac{c_{\rm AdS}}{3}w_{c}$, which matches the result in ref. Donnelly:2018bef . For verification, we holographically calculate the area of the RT surface, which is expressed as
$D_{12}$ $=\ell_{\rm AdS}\mathrm{arccosh}(1+2\sinh^{2}w_{c}\cosh^{2}t_{0}\sin^{2}x)$
$\simeq 2\ell_{\rm AdS}\log(2\sinh w_{c}\cosh t_{0}\sin x_{a})+\frac{\ell_{\rm AdS}}{2\sinh^{2}w_{c}\cosh^{2}t_{0}\sin^{2}x_{a}}.$ (108)
After dividing $D_{12}$ by $4G$ and substituting the relation $G=3\ell_{\rm AdS}/2c_{\rm AdS}$, it is consistent with (107) up to the second-order correction of the $T\bar{T}$ deformation.
5.2 dS/dS
For the dS/dS correspondence, the bulk metric is given by
$ds^{2}=\ell_{\rm dS}^{2}(dw^{2}+\sin^{2}w(-dt^{2}+\cosh^{2}td\phi^{2}))\,,$ (109)
with the holographic boundary specified at $w=w_{c}$. The subregion $A$ is choosen as $t=t_{0}$, $\phi\in(-\phi_{a},\phi_{a})$. The boundary induced metric is
$ds^{2}=\ell_{\rm dS}^{2}\sin^{2}w_{c}(-dt^{2}+\cosh^{2}td\phi^{2})\,.$ (110)
Comparing this with the metric in eq. (98), the only difference between the induced metrics for the dS and AdS cases is the replacement of the conformal factor $\sinh^{2}w$ with $\sin^{2}w$. As a result, the subsequent calculations are identical to those in the AdS case, except for replacing $\sinh^{2}w_{c}$ with $\sin^{2}w_{c}$. Consequently, the effective radius is given by $r_{\rm eff}=\ell_{\rm dS}\sin w_{c}\cosh t_{0}\sin\phi_{a}$. For dS/dS, the entanglement entropy expression for a general spatial interval is given by eq. (50). In this case, the effective radius must be less than $\ell_{\rm dS}$, which implies
$\sin w_{c}\cosh t_{0}\sin\phi_{a}<1.$ (111)
Finally, we can obtain the entanglement entropy of subregion $A$ as
$S_{A}=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{1}{(\sin w_{c}\cosh t_{0}\sin\phi_{a})^{2}}-1}}\right)\,.$ (112)
For the specific case $t=0$, we can derive the relation
$\frac{1}{(\sin w_{c}\sin\phi_{a})^{2}}-1=\cot^{2}\phi_{a}+\frac{\cot^{2}w_{c}}{\sin^{2}\phi_{a}}\,.$ (113)
Substituting this into (112), the entanglement entropy at $t_{0}=0$ becomes
$S_{A}=\frac{c_{\rm dS}}{3}\mathrm{arccot}\left(\sqrt{\cot^{2}\phi_{a}+\frac{\cot^{2}w_{c}}{\sin^{2}\phi_{a}}}\right)\,,$ (114)
which is consistent with the holographic calculation of the area of the RT surface in ref. Geng:2019ruz . For the case where the subsystem is defined as $\phi\in(-\frac{\pi}{2},\frac{\pi}{2})$, the entanglement entropy is given by $S_{A}=\frac{c_{\rm dS}}{3}w_{c}$, which agrees with the result in ref. Gorbenko:2018oov .
The next task is to determine whether the system’s entanglement entropy satisfies the strong subadditivity inequality. To achieve this, we compute the second-order derivative of the entanglement entropy with respect to the subsystem scale $\Delta\phi$, where $\Delta\phi=2\phi_{a}$. The result is
$\partial^{2}_{\Delta\phi}S_{A}(\Delta\phi)=-\frac{c_{\rm dS}}{12\sin^{2}\frac{\Delta\phi}{2}}\frac{\frac{1}{\sin^{2}w_{c}\cosh^{2}t_{0}}-1}{\left(\frac{1}{\sin^{2}\frac{\Delta\phi}{2}\sin^{2}w_{c}\cosh^{2}t_{0}}-1\right)^{3/2}}.$ (115)
We find that when $\sin w_{c}\cosh t_{0}>1$, the entanglement entropy always violates the strong subadditivity inequality. However, when $\sin w_{c}\cosh t_{0}<1$, for any $\Delta\phi$, the condition $r_{\rm eff}<\ell_{\rm dS}$ is always satisfied, and the entanglement entropy satisfies the strong subadditivity inequality, as illustrated in figure 12(a) and figure 12(b).
(a) $w_{c}=\frac{\pi}{2}$.
(b) $t_{0}=0$.
Figure 12: The second-order derivative of the entanglement entropy. 12(a) we take $w_{c}=\frac{\pi}{2}$, with $r_{\text{eff}}=\ell_{\rm dS}\cosh t_{0}\sin\frac{\Delta\phi}{2}$. In this case, we need to always ensure that $\cosh t_{0}\sin\frac{\Delta\phi}{2}<1$. 12(b) we take $t_{0}=0$, with $r_{\text{eff}}=\ell_{\rm dS}\sin w_{c}\sin\frac{\Delta\phi}{2}<\ell_{\rm dS}$.
For the two cases described above, the boundary positions are illustrated in figure. 13. The orange region corresponds to $\sin w_{c}\cosh t_{0}>1$, where the dual field theory is nonlocal. The blue region corresponds to $\sin w_{c}\cosh t_{0}<1$, where the dual field theory satisfies the strong subadditivity inequality. However, it can be proven that in this case, the entanglement entropy does not satisfy the boosted strong subadditivity inequality. In the dS2 background, the infinitesimal form of the boosted strong subadditivity inequality is given by Geng:2019ruz
$\text{BSSA}(\Delta\phi)=\partial^{2}_{\Delta\phi}S_{A}(\Delta\phi)+\frac{1}{\sin\Delta\phi}\partial_{\Delta\phi}S_{A}(\Delta\phi)\leq 0.$ (116)
Figure 13: The Penrose diagram of dS/dS correspondence. The black and blue dashed lines represent the codimension-one hypersurfaces at $t=0$ and $w=\frac{\pi}{2}$, respectively. The black curve represents the spacelike hypersurface at $t=t_{0}$, and the orange and blue curves represent the dS slices at $w=w_{c}$, where the dual field theory is located. The intersection of these two hypersurfaces corresponds to the critical case that saturates the strong subadditivity inequality, where $\sin w_{c}\cosh t_{0}=1$.
By substituting eq. (114) into the left-hand side of the boosted strong subadditivity inequality (116), we obtain
$\text{BSSA}(\Delta\phi)=\frac{c_{\rm dS}}{48}\frac{\sin^{2}\Delta\phi}{\sin^{6}\frac{\Delta\phi}{2}\sin^{2}w_{c}\left(\frac{1}{\sin^{2}\frac{\Delta\phi}{2}\sin^{2}w_{c}}-1\right)^{\frac{3}{2}}}>0,$ (117)
which indicates that, in this case, the dual field theory remains nonlocal. In particular, at the ultraviolet boundary where $w_{c}=\frac{\pi}{2}$, there is a significant violation of the boosted strong subadditivity inequality, as indicated by the purple line in figure 14.
Figure 14: The boosted strong subadditivity inequality violate in the dS/dS holography.
5.3 The half-dS holography
By introducing a timelike boundary $\Theta=\Theta_{0}$ in dS global spacetime, the half-dS holographic model was obtained in ref. Kawamoto:2023nki . The dual field theory resides on the timelike boundary, and the bulk spacetime is confined to the region where $0<\Theta<\Theta_{0}$. The induced metric on the boundary is given by
$ds^{2}=\ell_{\rm dS}^{2}(-dt^{2}+\cosh^{2}t\sin^{2}\Theta_{0}d\Phi^{2}).$ (118)
By rescaling the coordinate $t$ in a manner analogous to that discussed in Section 5.1, we obtain the conformal cylindrical metric
$ds^{2}=\ell_{\rm dS}^{2}\frac{\sin^{2}\Theta_{0}}{\cosh^{2}\left(\tilde{t}\sin\Theta_{0}\right)}\left(d\tilde{t}^{2}+d\Phi^{2}\right),$ (119)
where the coordinate transformation between $\tilde{t}$ and $t$ is
$\tilde{t}=\frac{1}{\sin\Theta_{0}}\ln\frac{\cot\frac{t}{2}+1}{\cot\frac{t}{2}-1}.$ (120)
Mapping from the cylinder to the plane as in eq. (101), the metric transforms into a conformally flat form
$ds^{2}=\frac{\ell_{\rm dS}^{2}\sin^{2}\Theta_{0}}{\cosh^{2}\left(\ln(\sqrt{T^{2}+X^{2}})\sin\Theta_{0}\right)}\frac{dT^{2}+dX^{2}}{T^{2}+X^{2}}.$ (121)
Next, we calculate the entanglement entropy for the subregion interval $(-\Phi_{a},\Phi_{a})$ at $t=t_{0}$. After the mapping, the coordinates of the subregion endpoints are
$(T_{1},X_{1})=\left(e^{\tilde{t}_{0}}\cos\Phi_{a},-e^{\tilde{t}_{0}}\sin\Phi_{a}\right),\quad(T_{2},X_{2})=\left(e^{\tilde{t}_{0}}\cos\Phi_{a},e^{\tilde{t}_{0}}\sin\Phi_{a}\right).$ (122)
Using the CHM mapping,
$T\rightarrow\frac{e^{\tilde{t}_{0}}\sin\Phi_{a}\sin\theta\sin\phi}{1+\sin\theta\cos\phi}+e^{\tilde{t}_{0}}\cos\Phi_{a}\,,\quad X\rightarrow\frac{e^{\tilde{t}_{0}}\sin\Phi_{a}\cos\theta}{1+\sin\theta\cos\phi}\,,$ (123)
we obtain the conformal spherical metric
$ds^{2}=\frac{2\Omega^{2}\ell_{\rm dS}^{2}\sin^{2}\Theta_{0}\sin^{2}\Phi_{a}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)}{(1+\sin\theta\cos\phi)(2+\sin(\theta+\phi-2\Phi_{a})+\sin(\theta-\phi+2\Phi_{a}))},$ (124)
where
$\Omega^{-1}=\cosh\left(\frac{1}{2}\ln\left(\frac{e^{2\tilde{t}_{0}}\left(2+\sin(\theta+\phi-2\Phi_{a})+\sin(\theta-\phi+2\Phi_{a})\right)}{2+2\sin\theta\cos\phi}\right)\sin\Theta_{0}\right).$ (125)
At the endpoints ($\theta=0,\pi$), the conformal factor becomes $\cosh^{-1}\tilde{t}_{0}$. Thus, the effective radius of the sphere in the Lorentzian signature is
$r_{\text{eff}}=\ell_{\rm dS}\cosh t_{0}\sin\Theta_{0}\sin\Phi_{a}.$ (126)
Therefore, the entanglement entropy is
$S=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{1}{(\cosh t_{0}\sin\Theta_{0}\sin\Phi_{a})^{2}}-1}}\right).$ (127)
In the case of $t_{0}=0$ and $\Theta_{0}=\frac{\pi}{2}$, the radius in the spherical partition function is $r_{\text{eff}}=\ell_{\rm dS}\sin\frac{\Delta\Phi}{2}$, where $\Delta\Phi=|-\Phi_{a}-\Phi_{a}|=2\Phi_{a}$. The entanglement entropy is
$S=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{1}{\sin^{2}\frac{\Delta\Phi}{2}}-1}}\right)=\frac{c_{\rm dS}}{6}\Delta\Phi.$ (128)
By replacing the field theory parameter $c_{\rm dS}$ with the gravitational parameter $\frac{3\ell_{\rm dS}}{2G}$, we find that the above result matches the holographic calculation presented in ref. Kawamoto:2023nki , confirming the validity of the half-dS holography. Additionally, the requirement that the effective radius of the sphere be smaller than the cosmological horizon aligns precisely with the condition given in eq. (3.6) of ref. Kawamoto:2023nki to ensure the existence of spacelike geodesics:
$r_{\text{eff}}=\ell_{\rm dS}\cosh t_{0}\sin\Theta_{0}\sin\frac{\Delta\Phi}{2}\leq\ell_{\rm dS}.$ (129)
In ref. Kawamoto:2023nki , the dual field theory in the half-dS holography was proposed to exhibit non-locality. It was concluded that when $\cosh t_{0}\leq\frac{1}{\sin\Theta_{0}}$, the entanglement entropy satisfies the strong subadditivity inequality, whereas when $\cosh t_{0}>\frac{1}{\sin\Theta_{0}}$, the strong subadditivity inequality is violated. As shown in figure 15, the orange line segment represents the region where the strong subadditivity inequality is violated, while the blue line segment represents the region where the strong subadditivity inequality is satisfied. Next, we demonstrate that even in the blue region where the strong subadditivity inequality is satisfied, the boosted strong subadditivity inequality is still violated.
Figure 15: The half-dS holography. The dual field theory resides on a timelike boundary at $\Theta=\Theta_{0}$, while the bulk spacetime covers the region $0<\Theta<\Theta_{0}$. Here we set $\cos T=\frac{1}{\cosh t}$. The blue intervals indicate $\cosh t\leq\frac{1}{\sin\Theta_{0}}$, i.e., $\Theta_{0}-\frac{\pi}{2}\leq T\leq\frac{\pi}{2}-\Theta_{0}$.
By observing eqs. (112) and (127), we find that the results from calculating the spatial interval entanglement entropy in the dS/dS holography and half-dS holography are similar. However, due to the difference in boundary-induced metrics, the boosted strong subadditivity inequality receives corresponding corrections. In the half-dS holography, the infinitesimal form of the boosted strong subadditivity inequality is
$\text{BSSA}(\Delta\tilde{\Phi})=\sin^{2}\Theta_{0}\left(\partial^{2}_{\Delta\tilde{\Phi}}S(\Delta\tilde{\Phi})+\frac{1}{\sin\Delta\tilde{\Phi}}\partial_{\Delta\tilde{\Phi}}S(\Delta\tilde{\Phi})\right)\leq 0,$ (130)
where $\Delta\tilde{\Phi}=\sin\Theta_{0}\Delta\Phi$. Substituting the result for the entanglement entropy into the above expression, we can get
$\text{BSSA}(\Delta\phi)=\frac{c_{\rm dS}}{48}\frac{\sin^{2}(\Delta\phi\sin\Theta_{0})}{\sin^{6}(\frac{1}{2}\Delta\phi\sin\Theta_{0})(\frac{1}{\sin^{2}\Theta_{0}\sin^{2}(\frac{1}{2}\Delta\phi\sin\Theta_{0})}-1)^{\frac{3}{2}}}>0.$ (131)
The result is illustrated in figure 16, which indicates that, in the half-dS holography, the boundary field theory still does not satisfy the boosted strong subadditivity inequality.
Figure 16: The boosted strong subadditivity inequality violates in the half-dS holography.
We have calculated the entanglement entropy of subsystem $A$ using the generalized spherical partition function method within the dS/dS correspondence and half-dS holography frameworks. Furthermore, we have reaffirmed the conclusions from Section 3.2 by applying the boosted strong subadditivity inequality and the strong subadditivity inequality: in dS3 spacetime, when the effective radius of the sphere is smaller than the cosmological horizon, the dual field theory exhibits non-locality.
6 Conclusion and Discussion
This paper investigated the effects of the $T\bar{T}$ deformation on various holographic dualities, including AdS/CFT, dS/CFT, and dS/dS, by analyzing the entanglement entropy of the boundary field theory. When the boundary is spherical, and the subsystem is a semicircular arc, the entanglement entropy can be derived from the classical partition function on the $n$-replica manifold, with the partition function governed by the trace flow equation.
To study the entanglement entropy in the dS case, we first derived the trace flow equation for the dS/CFT by applying a double Wick rotation to the AdS case. Then, we considered a simple scenario where the subsystem is a semicircular arc on the sphere, which corresponds to global coordinates in the dS/CFT, as well as the $t=0$ moment in both AdS/dS and dS/dS. Our findings are as follows:
  • In dS/CFT, the entanglement entropy is complex, indicating that the dual field theory is non-unitary.
  • In dS/dS, the entanglement entropy is real but violates the strong subadditivity inequality, suggesting that the dual field theory is unitary but non-local.
We then examined a more complicated scenario where the boundary-induced metric is time-dependent, covering AdS/dS, dS/dS, and half-dS holography. In these cases, since the expressions for the strong subadditivity inequality and the boosted strong subadditivity inequality depend on the specific form of the metric, and the boundary-induced metric is time-dependent, the locality of the boundary field theory may vary across different time periods, as discussed in ref. Kawamoto:2023nki . To address this, we further generalized the replica method on the sphere to accommodate more general manifolds and subsystems. We first verified this method in the global coordinates of AdS/CFT, dS/CFT, and the planar BTZ case, where the results are consistent with existing perturbative calculations Chen:2018eqk ; Chen:2023eic . We then applies it to the cases where the boundary-induced metrics are time-dependent. Our results show that while the subsystem’s entanglement entropy may satisfy the strong subadditivity inequality at certain time periods in dS/dS and half-dS holographies, the boosted strong subadditivity inequality is consistently violated, indicating that the dual field theories remain non-local throughout.
Acknowledgements.
We would like to thank Shan-Ming Ruan for reading our draft and providing comments. We also thank Yang He, and Shan-Ping Wu for their helpful discussion. This work was supported by the National Natural Science Foundation of China (Grants No. 12475056, No. 12235016, No. 12075101, No. 12475053, and No. 12247101), the 111 Project under (Grant No. B20063), the Major Science and Technology Projects of Gansu Province, and Lanzhou City’s scientific research funding subsidy to Lanzhou University. S.H. would appreciate the financial support from the Fundamental Research Funds for the Central Universities and Max Planck Partner Group and the Natural Science Foundation of China (NSFC) Grants No. 12075101 and No. 12235016. L.Z. is supported by the Science and Technology Development Plan Project of Jilin Province, China (No. 20240101326JC).
Appendix A The gravity calculation
A.1 The dS/Poincaré case
In the dS Poincaré spacetime,
$ds^{2}=\frac{\ell_{\rm dS}^{2}}{\eta^{2}}(-d\eta^{2}+dX^{2}+dT^{2})\,,$ (132)
the geodesic distance $D_{12}$ between two points “1” $(\eta,X,T)=(\eta_{c},-X_{a},0)$ and “2” $(\eta_{c},-X_{a},0)$ can be desolved by
$\cos\frac{D_{12}}{\ell_{\rm dS}}=-V_{1}V_{2}+U_{1}U_{2}+Y_{1}Y_{2}+W_{1}W_{2}.$ (133)
The uppercase coordinates “$U$, $Y$, $W$, $V$” denote the coordinates of the embedded spacetime
$ds^{2}=-dV^{2}+dW^{2}+dU^{2}+dY^{2}.$ (134)
The dS spacetime is a codimensional one hypersurface in the embedded spacetime
$-V^{2}+W^{2}+U^{2}+Y^{2}=\ell_{\rm dS}^{2},$ (135)
and they adhere to the following transformation relations with respect to the coordinates of the dS spacetime itself
$U$ $=\frac{\ell_{\rm dS}}{\eta}X,\quad V=\ell_{\rm dS}\sinh(\log(\frac{\ell_{\rm dS}}{\eta}))+\frac{1}{\eta}\frac{X^{2}+T^{2}}{2},$ (136)
$Y$ $=\frac{\ell_{\rm dS}}{\eta}T,\quad W=\ell_{\rm dS}\cosh(\log(\frac{\ell_{\rm dS}}{\eta}))-\frac{1}{\eta}\frac{X^{2}+T^{2}}{2}.$
Finally we can get the geodesic distance is
$D_{12}=\ell_{\rm dS}\arccos\left(\ell_{\rm dS}^{2}\left(1-\frac{2X_{a}^{2}}{\eta_{c}^{2}}\right)\right)\simeq-2i\ell_{\rm dS}\log\frac{2X_{a}}{\eta_{c}}+i\ell_{\rm dS}\frac{\eta_{c}^{2}}{2X_{a}^{2}}+\pi\ell_{\rm dS}.$ (137)
So the entanglement entropy of the subregion is
$S=\frac{D_{12}}{4G}\simeq-\frac{ic_{\rm dS}}{3}\log\frac{2X_{a}}{\eta_{c}}+\frac{ic_{\rm dS}\eta_{c}^{2}}{12X_{a}^{2}}+\frac{\pi c_{\rm dS}}{6}.$ (138)
This is consistent with eq. (68).
A.2 The global dS case
For the dS global coordinates, the coordinate transformation in embedding spacetime is
$U$ $=\ell_{\rm dS}\cosh t\sin\Theta\cos\Phi,\quad V=\ell_{\rm dS}\sinh t,$ (139)
$Y$ $=\ell_{\rm dS}\cosh t\sin\Theta\sin\Phi,\quad W=\ell_{\rm dS}\cosh t\cos\Theta.$
So the geodesic distance of two points $(t,\Theta,\Phi)$=$(t_{1},\Theta_{1},\Phi_{1})$, $(t_{2},\Theta_{2},\Phi_{2})$ in the dS global coordinates is
$\cos\frac{D_{12}}{\ell_{\rm dS}}=\left(\cos\Theta_{1}\cos\Theta_{2}+\sin\Theta_{1}\sin\Theta_{2}\cos(\Phi_{1}-\Phi_{2})\right)\cosh t_{1}\cosh t_{2}-\sinh t_{1}\sinh t_{2}.$ (140)
Now we can select the endpoints of the subsystem as $(t,\Theta,\Phi)$=$(t_{c},\frac{\pi}{2},-\Phi_{a})$, $(t_{c},\frac{\pi}{2},\Phi_{a})$, so we have
$D_{12}=\ell_{\rm dS}\cos^{-1}\left(\cos(2\Phi_{a})\cosh^{2}t_{c}-\sinh^{2}t_{c}\right)=\ell_{\rm dS}\cos^{-1}\left(1-2\cosh^{2}t_{c}\sin^{2}\Phi_{a}\right).$ (141)
Considering that the finite cutoff $t_{c}$ tends to infinity, we expand the above results to the second-order
$D_{12}\simeq-2i\ell_{\rm dS}\log(2\cosh t_{c}\sin\Phi_{a})+\frac{i\ell_{\rm dS}}{2\cosh^{2}t_{c}\sin^{2}\Phi_{a}}+\pi\ell_{\rm dS}.$ (142)
Finally, we have the entanglement entropy
$S=\frac{D_{12}}{4G}\simeq-\frac{ic_{\rm dS}}{3}\log(2\cosh t_{c}\sin\Phi_{a})+\frac{ic_{\rm dS}}{12\cosh^{2}t_{c}\sin^{2}\Phi_{a}}+\frac{\pi c_{\rm dS}}{6}.$ (143)
We can get the same result as eq. (4.2.2).

### v3

@@
 Abstract
-In recent years, the holographic duality between $T\bar{T}$-deformed conformal field theory (CFT) and Anti-de Sitter (AdS) space with finite radial truncation has received significant attention.
+In recent years, the holographic duality between $T\bar{T}$-deformed conformal field theory (CFT) and Anti-de Sitter (AdS) spacetime with finite radial cutoff has received significant attention.
 The study of $T\bar{T}$ deformation within the framework of de Sitter (dS)/CFT duality has also progressed.
 This paper shows that the trace flow equation in dS spacetime can be analytically extended from its AdS counterpart through double Wick rotations.
-Additionally, we generalize the replica sphere method in both AdS and dS holography to derive a general expression for the entanglement entropy of arbitrary single spatial intervals.
-For both finite size and finite temperature systems, we obtain the entanglement entropy corrections due to $T\bar{T}$ deformation.
-Finally, using strong subadditivity and boosted strong subadditivity, we show that the dual field theory on the timelike boundary is non-local in dS holography.
+Meanwhile, we generalize the replica method in both AdS and dS holography to derive a general expression for the entanglement entropy of arbitrary single spatial intervals within the $T\bar{T}$-deformed framework.
+For both finite size and finite temperature systems, we obtain the analytical expression for the entanglement entropy after $T\bar{T}$ deformation.
+Finally, in dS/dS holography and half-dS holography, we find that the dual field theory exhibits non-locality by analyzing the strong subadditivity and boosted strong subadditivity of entanglement entropy.
 Keywords:
@@
 $S_{\mathrm{QFT}}\simeq S_{\mathrm{CFT}}+2\pi\lambda\int d^{2}x\sqrt{g}~{}T\bar{T}.$ (2)
-From the perspective of the AdS/CFT duality Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj , the $T\bar{T}$-deformed CFT corresponds to the semi-classical gravity theory living in an AdS spacetime with finite cutoff McGough:2016lol ; Kraus:2018xrn , as shown in figure 1.
+From the perspective of the AdS/CFT duality Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj , the $T\bar{T}$-deformed CFT corresponds to the semiclassical gravity theory living in an AdS spacetime with finite cutoff McGough:2016lol ; Kraus:2018xrn , as shown in figure 1.
 Figure 1: The holography $T\bar{T}$ deformation.
@@
 For the holographic duality with $T\bar{T}$ deformation, this duality has been discussed in refs.
-Chen:2018eqk ; Jeong:2019ylz ; Jiang:2019epa ; Allameh:2021moy ; Setare:2022qls ; He:2023xnb ; Apolo:2023ckr .
-In ref.
-Chen:2018eqk , the authors investigated the entanglement entropy of the $T\bar{T}$-deformed field theory at the level of perturbation (2).
-They found that for the system at finite temperature, the entanglement entropy receives a first-order $\lambda$ correction from the $T\bar{T}$ deformation, but for a system of finite size, this first-order correction disappears.
-To explore the impact of $T\bar{T}$ deformation on the entanglement entropy in finite-size systems, ref.
-Jiang:2023ffu examined the timelike entanglement entropy, drawing similar conclusions to previous studies Li:2022tsv ; Doi:2023zaf .
-They found that the timelike entanglement entropy receives no first-order correction at finite temperature, but there is a first-order correction in the finite size case.
-Additionally, the reflected entropy Dutta:2019gen and pseudo entropy Nakata:2020luh of the $T\bar{T}$ deformed field theory have been studied in refs.
+Chen:2018eqk ; Jeong:2019ylz ; Jiang:2019epa ; Allameh:2021moy ; Setare:2022qls ; He:2023xnb ; Apolo:2023ckr ; Apolo:2023vnm .
+Meanwhile, the reflected entropy Dutta:2019gen and pseudo entropy Nakata:2020luh of the $T\bar{T}$-deformed field theory have been studied in refs.
 Basu:2024bal ; He:2023wko .
 For more interesting work on $T\bar{T}$ deformation, please refer to refs.
-He:2019vzf ; Asrat:2020uib ; Khoeini-Moghaddam:2020ymm ; Li:2020pwa ; Jeong:2022jmp ; He:2023obo ; Tian:2023fgf ; Apolo:2023vnm ; Basu:2023bov ; Basu:2023aqz ; He:2024pbp ; Tsolakidis:2024wut ; Babaei-Aghbolagh:2024hti .
-The holographic principle has been well established in the AdS spacetime with finite cutoff $r=r_{c}$.
-However, given that our real universe is more accurately described by a dS metric, it is crucial to explore whether the holographic principle extends to the dS spacetime Witten:2001kn ; Maldacena:2002vr ; Miyaji:2015yva ; Jiang:2023loq ; Chang:2023gkt ; Chen:2023prz ; Doi:2024nty .
-In the usual dS/CFT duality Strominger:2001pn ; Strominger:2001gp , the CFT resides on a future spacelike boundary, which renders the dual field theory non-unitary Maldacena:2002vr ; Doi:2022iyj .
-Due to the challenges of the dS/CFT duality Anninos:2011ui ; Ng:2012xp , there are only a few clearly established examples Hikida:2022ltr ; Hikida:2021ese ; Chen:2022ozy .
-In recent years, other holographic duality models in the dS spacetime, such as the dS/dS correspondence Alishahiha:2004md ; Alishahiha:2005dj ; Dong:2018cuv ; Geng:2019bnn ; Geng:2019ruz ; Geng:2020kxh ; Geng:2021wcq and half-dS holography Kawamoto:2023nki , have gained attention and been studied.
-In these two models, the holographic boundary is a timelike boundary of the dS spacetime, so the dual field theory is expected to be unitary.
-In the dS/dS correspondence, the cosmological constant term also appears in the trace flow equation of the $T\bar{T}$ deformation, thus the $T\bar{T}$ deformation of this model is referred to as $T\bar{T}+\Lambda_{2}$ deformation Gorbenko:2018oov ; Shyam:2021ciy ; Coleman:2021nor ; Torroba:2022jrk ; Batra:2024kjl .
-We aim to study the impact of $T\bar{T}$ deformation on the entanglement entropy in the dS/CFT duality.
+He:2019vzf ; Asrat:2020uib ; Khoeini-Moghaddam:2020ymm ; Li:2020pwa ; Jeong:2022jmp ; He:2023obo ; Tian:2023fgf ; Basu:2023bov ; Basu:2023aqz ; He:2024pbp ; Tsolakidis:2024wut ; Babaei-Aghbolagh:2024hti ; He:2025ppz ; He:2025fdz .
+Although the holographic principle has been well established in the AdS spacetime with finite cutoff $r=r_{c}$, it remains crucial to investigate whether the holographic principle extends to dS spacetime Witten:2001kn ; Maldacena:2002vr ; Anninos:2012ft ; Miyaji:2015yva ; Narayan:2017xca ; Hikida:2021ese ; Chen:2022ozy ; Chang:2023gkt ; Chen:2023prz ; Doi:2024nty , since our universe is more accurately described by a dS metric.
+In the standard dS/CFT duality Strominger:2001pn ; Strominger:2001gp , the CFT resides at a future spacelike infinite, which leads to some exotic characteristics Anninos:2011ui ; Ng:2012xp ; Hikida:2022ltr ; Doi:2022iyj .
+For example, the central charge is imaginary.
+Based on these exotic characteristics, a precise construction of the dual CFT in dS spacetime remains elusive.
+For the holographic entanglement entropy in dS/CFT, by calculating the extreme value of the area of the RT surface anchored at the future spacelike infinite boundary on the gravitational side, refs.
+Narayan:2015vda ; Narayan:2015oka ; Narayan:2016xwq ; Narayan:2022afv ; Narayan:2023zen have revealed the presence of complex-valued saddle points, which coincides with the imaginary central charge.
+At the same time, this observation aligns with results obtained from the analytic continuation of holographic entanglement entropy in Euclidean AdS spacetime, thereby offering a potential avenue to explore the dS/CFT duality through double Wick rotation Sato:2015tta .
+In addition, other holographic duality models in the dS spacetime, such as the dS/dS holography Alishahiha:2004md ; Alishahiha:2005dj ; Dong:2018cuv ; Geng:2019bnn ; Geng:2019ruz ; Geng:2020kxh ; Geng:2021wcq and half-dS holography Kawamoto:2023nki , have also gained attention and been studied.
+In these two models, the dual field theory resides on the timelike boundary of the dS spacetime.
+The primary purpose of this paper is to study the impact of the $T\bar{T}$ deformation on holographic entanglement entropy in the dS spacetime, including three holographic models: dS/CFT, dS/dS, and half-dS.
+For the $T\bar{T}$-deformed dS/CFT Chen:2023eic , the dual field theory is on the spacelike slice $t=t_{c}$, which coincides with the idea of Cauchy slice holography Araujo-Regado:2022gvw ; Araujo-Regado:2022jpj .
+In the $T\bar{T}$-deformed dS/dS, the cosmological constant term also appears in the trace flow equation of the $T\bar{T}$ deformation, thus the $T\bar{T}$ deformation of this model is referred to as $T\bar{T}+\Lambda_{2}$ deformation Gorbenko:2018oov ; Shyam:2021ciy ; Coleman:2021nor ; Torroba:2022jrk ; Batra:2024kjl .
 However, the deformed field theory no longer retains conformal symmetry, rendering the original method for calculating the entanglement entropy inapplicable.
-Fortunately, when the deformed CFT is situated on a two-dimensional sphere, the entanglement entropy of a subregion with endpoints at antipodal points can be calculated via the replica method Calabrese:2009qy ; Donnelly:2018bef ; Caputa:2019pam ; Deng:2023pjs .
-In this context, by adjusting the spherical radius $r$, we can study the renormalization group flow Wilson:1973jj ; Polchinski:1983gv of the dual field theory under $T\bar{T}$ deformation, with $r$ functioning as the radial coordinate in the bulk spacetime.
+Fortunately, when the $T\bar{T}$-deformed CFT is situated on a two-dimensional sphere, the entanglement entropy of a subregion with endpoints at antipodal points can be calculated by replicating the spherical geometry Calabrese:2009qy ; Donnelly:2018bef ; Caputa:2019pam ; Deng:2023pjs ; Grieninger:2019zts .
+In this context, by adjusting the spherical radius $r$, we can also study the renormalization group flow Wilson:1973jj ; Polchinski:1983gv of the dual field theory under $T\bar{T}$ deformation, with $r$ functioning as the radial coordinate in the bulk spacetime.
 This method is an extension of the approach proposed in ref.
@@
 Chen:2023eic , the replica method was employed within the $T\bar{T}$-deformed dS/CFT duality to calculate the entanglement entropy for the subregion between two antipodal points on the sphere.
-In this paper, we present a general expression for the entanglement entropy of arbitrary spatial intervals in both AdS and dS holography by combining the Casini-Huerta-Myers (CHM) map Casini:2011kv with the replica method on a sphere.
+In the $T\bar{T}$-deformed theories, non-perturbative calculations of entanglement entropy have been achieved using the replica trick, yielding exact results on spherical and flat manifolds Donnelly:2018bef ; Lewkowycz:2019xse .
+However, the non-perturbative calculation on general manifolds is still lacking.
+In this paper, we present a general expression for the entanglement entropy of arbitrary spatial intervals in general manifolds by combining the Casini-Huerta-Myers (CHM) map Casini:2011kv with the replica method on a sphere.
 The results are
-  • AdS($0<r_{\text{eff}}<\infty$): $S=\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\text{eff}}^{2}}}}\right),$
-  • dS($\ell_{\rm dS}<r_{\text{eff}}<\infty$): $S=-\frac{ic_{\rm dS}}{3}\mathrm{arccoth}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6},$
-  • dS($0<r_{\text{eff}}<\ell_{\rm dS}$): $S=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}-1}}\right),$
+  • AdS($0<r_{\text{eff}}<\infty$): $S=\frac{c_{\rm AdS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\text{eff}}^{2}}}}\right),$
+  • dS($\ell_{\rm dS}<r_{\text{eff}}<\infty$): $S=-\frac{ic_{\rm dS}}{3}\coth^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6},$
+  • dS($0<r_{\text{eff}}<\ell_{\rm dS}$): $S=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}-1}}\right),$
 where $\ell_{(\rm{A})dS}$ represents the (A)dS radius, $c_{\rm(A)dS}=\frac{3\ell_{(\rm{A})dS}}{2G}$ is the Brown-Henneaux central charge Brown1986 , $\lambda_{\rm(A)dS}=8G\ell_{(\rm{A})dS}$ is the $T\bar{T}$ deformation parameter [^1] As mentioned in ref.
@@
 In the case of dS holography, we find that the results for the holographic entanglement entropy can be summarized in a single formula:
-$S=-\frac{ic_{\rm dS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}}}\right)\,.$ (3)
-When the effective radius of the sphere is greater than the cosmological horizon radius ($\ell_{\rm dS}<r_{\text{eff}}<\infty$), this formula corresponds to the result of the $T\bar{T}$-deformed dS/CFT duality, where the dual field theory is non-unitary.
-When the effective radius of the sphere is less than the cosmological horizon radius ($0<r_{\text{eff}}<\ell_{\rm dS}$), this formula agrees with the results of dS/dS correspondence and half-dS holography, where the dual field theory is non-local Geng:2019ruz ; Geng:2020kxh ; Kawamoto:2023nki .
-This paper is organized as follows: In Section 2, we review the results from ref.
-Kraus:2018xrn to derive the trace flow equations in AdS/CFT with Euclidean signature and in dS/CFT with Lorentzian signature.
-We then demonstrate the extension from one result to the other via a double Wick rotation.
-In Section 3, we review the calculation of the entanglement entropy of the $T\bar{T}$-deformed CFT, providing general expressions for both AdS and dS holographies.
-Based on these findings, we calculate the $C$-function and discuss the non-unitarity and non-locality of the dual field theory.
-In Section 4, we review and generalize the CHM method for calculating the entanglement entropy of arbitrary spatial intervals in the $T\bar{T}$-deformed theory.
-We explore the effects of the $T\bar{T}$ deformation in finite temperature and finite size systems.
-In Section 5, we first verify the calculation of the entanglement entropy in the AdS/dS correspondence.
-We then examine the non-locality of the dual field theory in dS/dS correspondence and half-dS holography by analyzing the strong subadditivity and the boosted strong subadditivity.
-Finally, we summarize our results and present a discussion in Section 6.
+$S=-\frac{ic_{\rm dS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}}}\right)\,.$ (3)
+When the effective radius of the sphere is greater than the cosmological horizon radius ($\ell_{\rm dS}<r_{\text{eff}}<\infty$), this formula corresponds to the result of the $T\bar{T}$-deformed dS/CFT duality.
+When the effective radius of the sphere is less than the cosmological horizon radius ($0<r_{\text{eff}}<\ell_{\rm dS}$), this formula agrees with the results of dS/dS holography and half-dS holography.
+Through an analysis of entanglement entropy compliance with strong subadditivity (SSA) and boosted strong subadditivity (BSSA) in both dS/dS holography and half-dS holography, we demonstrate that the dual field theory--residing on a timelike slice in dS spacetime--exhibits nonlocality in these holographic frameworks Geng:2019ruz ; Geng:2020kxh ; Kawamoto:2023nki .
+This paper is organized as follows: In Section 2, we review the derivation of trace flow equations in AdS/CFT with Euclidean signature and in dS/CFT with Lorentzian signature.
+Subsequently, we demonstrate the analytic continuation from the AdS/CFT result to the dS/CFT counterpart through double Wick rotation.
+In Section 3, we review the computation of entanglement entropy for $T\bar{T}$-deformed CFTs, presenting general expressions applicable to both AdS and dS holography.
+In Section 4, we calculate the non-perturbative results for the entanglement entropy of arbitrary single spatial intervals in $T\bar{T}$-deformed theories on general manifolds by employing the CHM map.
+We further investigate the impact of $T\bar{T}$ deformations in finite temperature and finite size systems.
+In Section 5, we examine the non-locality of the dual field theory in dS/dS and half-dS holography by scrutinizing SSA and BSSA.
+Finally, we summarize our findings and discuss potential future directions in Section 6.
 2 Holographic $T\bar{T}$ Flow
-In this section, we review the holographic derivation of the trace flow equation for the $T\bar{T}$ deformation Shyam:2017znq ; Hartman:2018tkw .
+In this section, we review the holographic derivation of the trace flow equation for the $T\bar{T}$ deformation Shyam:2017znq ; Hartman:2018tkw ; Kraus:2018xrn .
 We then show that the trace flow equation in the dS/CFT duality can be obtained from its AdS counterpart via a double Wick rotation.
@@
 $T_{ij}=\frac{2}{\sqrt{\gamma}}\frac{\delta S}{\delta\gamma^{ij}}=\frac{1}{8\pi G}\left(K_{ij}-K\gamma_{ij}+\frac{1}{\ell_{\rm AdS}}\gamma_{ij}\right).$ (5)
-The EAdS3 metric in Fefferman-Graham Fefferman1985 coordinates is
+The EAdS3 metric in Fefferman-Graham FeffermanGraham1985 coordinates is
 $ds^{2}=\frac{\ell_{\rm AdS}^{2}}{r^{2}}\left(dr^{2}+h_{ij}dx^{i}dx^{j}\right),\quad\gamma_{ij}=\frac{\ell_{\rm AdS}^{2}}{r_{c}^{2}}h_{ij}.$ (6)
@@
 $=\sqrt{\gamma}\left(-4\pi G(T^{ij}T_{ij}-(T^{i}_{i})^{2})-\frac{i}{\ell_{\rm dS}}T^{i}_{i}-\frac{\mathcal{R}^{(2)}}{16\pi G}\right)\hat{\Psi}_{\rm dS}=0,$
-where $T^{ij}=-\frac{2i}{\sqrt{\gamma}}\Pi^{ij}$. This leads to the $T\bar{T}$ trace flow equation for the dS/CFT is
+where $T^{ij}=-\frac{2i}{\sqrt{\gamma}}\Pi^{ij}$. This leads to the $T\bar{T}$ trace flow equation for the dS/CFT being
 $T^{i}_{i}=\frac{i\ell_{\rm dS}}{16\pi G}\mathcal{R}^{(2)}+i4\pi G\ell_{\rm dS}\left(T^{ij}T_{ij}-(T^{i}_{i})^{2}\right).$ (22)
-By comparing this with the field theory equation, the correspondences between the Brown-Henneaux central charge, the deformation parameter, and the gravitational quantities for the dS spacetime Chen:2023eic are given by
-$c=-i\frac{3\ell_{\rm AdS}}{2G}=-ic_{\rm dS}\,,\quad\lambda=-i8G\ell_{\rm AdS}=-i\lambda_{\rm dS}\,.$ (23)
+By comparing this with the field theory equation, the correspondences between the Brown-Henneaux central charge, the deformation parameter, and the gravitational quantities for the dS spacetime are given by
+$c=-i\frac{3\ell_{\rm dS}}{2G}=-ic_{\rm dS}\,,\quad\lambda=-i8G\ell_{\rm dS}=-i\lambda_{\rm dS}\,.$ (23)
 2.3 Double Wick Rotation to dS/CFT
@@
 By applying the transformation rules in Table 1 to the on-shell action of AdS gravity (4), we obtain the on-shell action for dS gravity, multiplied by an imaginary factor $i$:
-$S_{\rm AdS}$ $\rightarrow-\frac{i}{16\pi G}\int_{\mathcal{M}}d^{3}x\sqrt{-g}\left(\mathcal{R}^{(3)}-\frac{2}{\ell_{\rm dS}^{2}}\right)$
-$\quad-\frac{1}{8\pi G}\int_{\partial\mathcal{M}}d^{2}x\sqrt{\gamma}\left(iK-\frac{i}{\ell_{\rm dS}}\right)$
+$S_{\rm AdS}$ $\rightarrow-\frac{i}{16\pi G}\int_{\mathcal{M}}d^{3}x\sqrt{-g}\left(\mathcal{R}^{(3)}-\frac{2}{\ell_{\rm dS}^{2}}\right)-\frac{1}{8\pi G}\int_{\partial\mathcal{M}}d^{2}x\sqrt{\gamma}\left(iK-\frac{i}{\ell_{\rm dS}}\right)$
 $=-iS_{\rm dS}.$ (25)
-The double Wick rotation implies that the partition function of AdS gravity corresponds to the Bunch-Davies wavefunction of dS gravity.
-Additionally, all quantities in the Wheeler-DeWitt equation for the AdS spacetime are mapped to their corresponding quantities in the Wheeler-DeWitt equation for the dS spacetime.
+Under the saddle point approximation, the double Wick rotation establishes that the wavefunction of AdS gravity, given by $\Psi=e^{-S_{\text{AdS}}}$, corresponds to the Bunch-Davies wavefunction of dS gravity, expressed as $\Psi_{\text{dS}}=e^{iS_{\text{dS}}}$.
 As a result, eqs.
@@
 (20) and (21), respectively.
-Therefore, the trace flow equation in the dS/CFT can be directly derived by applying this rotation to the trace flow equation in the AdS/CFT.
+This means that the trace flow equation (22) in the dS/CFT can be directly derived by applying the double Wick rotation to the trace flow equation (9) in the AdS/CFT.
+At the same time, it should be noted that although we have linked the WDW constraints in dS spacetime with the WDW constraints in AdS spacetime through the double Wick rotation, the Hilbert space of dS quantum gravity may not be derived by the analytical continuation from that of AdS [^3] Thanks to the referee for pointing out the limitations of the analytical continuation method here..
+This is because the radial wave functional in AdS spacetime obeys regularity at $r=0$, fixing the asymptotic form at $r=\infty$ to the phase factor times the partition function of the CFT.
+This constraint does not apply in dS spacetimeChakraborty:2023yed ; Cianfrani:2013oja ; Chakraborty:2023los .
 3 The Replica Method for the $T\bar{T}$-Deformed Field Theory
@@
 $ds_{n}^{2}=r^{2}\left(d\theta^{2}+n^{2}\sin^{2}\theta\,d\phi^{2}\right).$ (27)
-The entanglement entropy of the subregion $A$ is computed by taking the limit $n\rightarrow 1$ of the conical entropy Headrick:2010zt ; Hung:2011nu ; Lewkowycz:2013nqa
+The entanglement entropy of the subregion $A$ is computed by taking the limit $n\rightarrow 1$ of the conical entropy Headrick:2010zt ; Hung:2011nu
 $S_{A}=\lim_{n\to 1}\left(1-n\frac{\partial}{\partial n}\right)\ln Z_{n},$ (28)
@@
 The trace flow equation and the conservation equation determine the stress tensor as follows:
-$\begin{cases}{T^{a}}_{a}=-\frac{c}{24\pi}R^{(2)}-\frac{\pi\lambda}{2}\left[T^{ab}T_{ab}-({T^{a}}_{a})^{2}\right],\\ \nabla_{a}{T^{a}}_{b}=0.\end{cases}$ (31)
+$T^{a}_{a}=-\frac{c}{24\pi}R^{(2)}-\frac{\pi\lambda}{2}\left[T^{ab}T_{ab}-(T^{a}_{a})^{2}\right],$ (31)
+$\nabla_{a}T^{a}_{b}=0.$ (32)
 Due to the maximal symmetry of the geometry, the quasi-local stress tensor is assumed to take the form
@@
 In this case, the parameter $\alpha$ determines the stress tensor, which is uniquely defined by the trace flow equations.
-This approach is specifically designed to calculate the entanglement entropy of the subregion between two antipodal points on the sphere.
-Future work aims to extend this framework to explore broader implications of the holographic principle, building on the refs.
-Grieninger:2019zts ; Lewkowycz:2019xse .
-3.1 AdS case
+3.1 AdS spacetime
 By substituting the assumed form of the stress tensor $T_{ij}=\alpha\gamma_{ij}$ into eq. (9), the equation for $\alpha$ is obtained as
@@
 Substituting this expression for $\alpha$ into eq. (29), the $n$-replica partition function $Z_{n}$ can be computed by integrating with respect to $\alpha$:
-$\ln Z_{n}=\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\right)+\frac{4r^{2}}{\lambda_{\rm AdS}}\left(-1+\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}\right).$ (36)
+$\ln Z_{n}=\frac{c_{\rm AdS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\right)+\frac{4r^{2}}{\lambda_{\rm AdS}}\left(-1+\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}\right).$ (37)
 The entanglement entropy of the subregion $A$ is then obtained as
-$S_{A}=\lim_{n\to 1}\left(1-\frac{r}{2}\frac{d}{dr}\right)\log Z_{n}=\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\right).$ (37)
+$S_{A}=\lim_{n\to 1}\left(1-\frac{r}{2}\frac{d}{dr}\right)\log Z_{n}=\frac{c_{\rm AdS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\right).$ (38)
 The entanglement entropy between two antipodal points on a sphere reflects the number of degrees of freedom in the field theory defined on the sphere Casini:2012ei .
@@
 Since the $T\bar{T}$ trace flow equation is equivalent to the holographic renormalization group flow equation, the renormalization group flow can be controlled by adjusting the radius of the sphere deBoer:1999tgo ; Verlinde:1999xm ; Heemskerk:2010hk ; Nishioka:2018khk ; Grieninger:2023knz ; Kamei:2024iei .
-In the AdS case, the $C$-function is given by
-$C(r)=3rS_{A}^{\prime}(r)=c_{\rm AdS}\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\,.$ (38)
-As illustrated in figure 3, the $C$-function decreases monotonically as $r$ decreases, consistent with the number of degrees of freedom decreases along the renormalization group flow.
+In the AdS case, the $C$-function can be constructed through the entanglement entropy.
+$C(r)\equiv 3rS_{A}^{\prime}(r)=c_{\rm AdS}\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}}}\,.$ (39)
+As illustrated in figure 3, the $C$-function decreases monotonically as $r$ decreases.
 Additionally, as $r$ approaches infinity, the $C$-function converges to the central charge of the two-dimensional CFT, indicating that the renormalization group flow approaches the UV fixed point Nakayama:2013is .
@@
 As the boundary cutoff $r=r_{c}$ is pushed further into the bulk, the deformation parameter increases and the theory deviates from the CFT increasingly.
-The $C$-theorem can be proven using locality and unitarity, but the converse is not true.
-Therefore, the locality of the field theory is next examined through the strong subadditivity inequality and the boosted strong subadditivity inequality.
-The strong subadditivity inequality is given by
-$S(A)+S(B)-S(A\cup B)-S(A\cap B)\geq 0,$ (39)
-where the subsystems $A$ and $B$ are shown in figure 4(a). In the limit where $A\rightarrow B$ and $|A|=|B|=L$, the strong subadditivity inequality can be expressed in its infinitesimal form as
-$L^{2}S^{\prime\prime}(L)\leq 0,$ (40)
-where $L$ represents the size of the subsystems.
-(a) strong subadditivity
-(b) boosted strong subadditivity
-Figure 4: The left figure shows two spatial regions $A$ and $B$ at $t=0$. The right figure considers the regions $A$ and $B$ after a boost, resulting in two spacelike regions.
-A stronger criterion for proving locality, compared to the strong subadditivity, is the boosted strong subadditivity, which can be derived by applying a boost transformation to subsystems $A$ and $B$, as shown in figure 4(b).
-The infinitesimal form of the boosted strong subadditivity inequality is expressed as
-$L^{2}S^{\prime\prime}(L)+LS^{\prime}(L)\leq 0.$ (41)
-Since the subsystems correspond to great arcs between antipodal points on a sphere, their size is proportional to the sphere’s radius, i.e., $L=\pi r$.
-Consequently, the derivatives with respect to $L$ in the strong subadditivity and the boosted strong subadditivity inequalities are equivalent to derivatives with respect to $r$.
-Taking the second derivative of eq.
-(37) with respect to $r$ yields the strong subadditivity inequality
-$r^{2}\partial^{2}_{r}S_{A}(r)=-\frac{c_{\rm AdS}}{3\left(1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r^{2}}\right)^{\frac{3}{2}}}<0.$ (42)
-This result shows that the strong subadditivity inequality is always satisfied in the AdS case, while the boosted strong subadditivity inequality is violated.
-This violation is consistent within the framework of $T\bar{T}$-deformed theories, where the boundary field theory resides on a finitely truncated surface.
-In such scenarios, the entanglement wedge of the boosted subsystem $A$ is not fully contained within the entanglement wedge of $A\cup B$ Lewkowycz:2019xse .
-3.2 dS case
-In the context of the dS spacetime, there are two primary holographic models: dS/CFT and dS/dS.
-The focus here is on the dS/dS correspondence in Euclidean signature, where the holographic boundary is a 2-dimensional sphere.
-The metric of the dS spacetime in global coordinates is given by
+3.2 dS spacetime
+In the context of dS spacetime, two primary holographic models are considered: dS/CFT and dS/dS holography, as illustrated in figures 4(a) and 4(b).
+The dS/CFT duality is conventionally formulated in global dS coordinates, whose metric is expressed as
 $ds^{2}=\ell_{\rm dS}^{2}\left[-dt^{2}+\cosh^{2}t\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right)\right],$ (40)
-where the radius of the boundary sphere is $r=\ell_{\rm dS}\cosh t$, which exceeds $\ell_{\rm dS}$ for any finite cutoff at $t=t_{c}$. In dS/dS coordinates, the metric of the dS spacetime is
-$ds^{2}=\ell_{\rm dS}^{2}\left[dw^{2}+\sin^{2}w\left(-d\tau^{2}+\cosh^{2}\tau\,d\phi^{2}\right)\right],$ (44)
-where the effective sphere radius is $r=\ell_{\rm dS}\sin w<\ell_{\rm dS}$.
-The point $r=\ell_{\rm dS}$ corresponds to the cosmological horizon of the dS spacetime.
-When $r>\ell_{\rm dS}$, the hypersurface $r=r_{c}$ is spacelike, which aligns with the standard dS/CFT duality.
-Conversely, when $r<\ell_{\rm dS}$, the hypersurface $r=r_{c}$ becomes timelike, corresponding to the dS/dS correspondence, as illustrated in figure 5.
-Figure 5: The Penrose diagram for the dS spacetime.
-The orange solid line represents future infinity, while the orange dashed line represents the finite cutoff at $t=t_{c}$.
-The red squares represent the boundaries at $w=0,\pi$ in dS/dS coordinates.
-The blue solid line represents $w=\pi/2$, and the blue dashed line represents the finite cutoff at $w=w_{c}$.
+where the $T\bar{T}$-deformed CFT resides on a boundary at a finite cutoff $t=t_{c}$ and the radius of the boundary sphere, $r=\ell_{\rm dS}\cosh t_{c}$, exceeds $\ell_{\rm dS}$ for any finite cutoff except $t_{c}=0$.
+For dS/dS holography, the metric in dS/dS coordinates is
+$ds^{2}=\ell_{\rm dS}^{2}\left[dw^{2}+\sin^{2}w\left(-dt^{2}+\cosh^{2}t\,d\phi^{2}\right)\right].$ (41)
+In the Euclidean signature, the dual field theory is defined on a two-dimensional sphere with a radius $r=\ell_{\rm dS}\sin w_{c}$.
+Notably, for any finite radial cutoff $w_{c}$, the sphere’s radius $r$ remains strictly less than $\ell_{\rm dS}$.
+The critical condition $r=\ell_{\rm dS}$ corresponds to the central slices of dS spacetime, as depicted by the orange and blue dashed lines in Figure 4.
+According to the surface/state correspondence Miyaji:2015yva , since the central slice in dS spacetime is an extremal surface devoid of internal entanglement, the dual field theory on it should exhibit non-locality, i.e., we expect the $C$-function to diverge.
+(a) dS/CFT
+(b) dS/dS
+Figure 4: The Penrose diagram for de Sitter spacetime.
+4(a)The orange solid line represents the finite cutoff at $t=t_{c}$, while the orange dashed line represents $t=0$.
+4(b)The green squares represent the boundaries at $w=0,\pi$ in dS/dS coordinates.
+The blue dashed line represents $w=\pi/2$, and the blue solid line represents the finite cutoff at $w=w_{c}$.
+dS/CFT:
 For the dS/CFT duality, the assumed form of the quasi-local stress tensor (33) is substituted into eq. (22) to determine the value of $\alpha$:
@@
 By integrating $\alpha$ with respect to $r$, the $n$-replica partition function is obtained as
-$\ln Z_{n}=-\frac{ic_{\rm dS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)+\frac{i4r^{2}}{\lambda_{\rm dS}}\left(-1+\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}\right).$ (46)
+$\ln Z_{n}=-\frac{ic_{\rm dS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)+\frac{i4r^{2}}{\lambda_{\rm dS}}\left(-1+\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}\right).$ (43)
 The entanglement entropy between of $A$ is then given by
-$S_{A}=\lim_{n\to 1}\left(1-\frac{r}{2}\frac{d}{dr}\right)\ln Z_{n}=-\frac{ic_{\rm dS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right).$ (47)
-In this case, the argument of the inverse hyperbolic function is greater than one, making $\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)$ a complex value.
+$S_{A}=\lim_{n\to 1}\left(1-\frac{r}{2}\frac{d}{dr}\right)\ln Z_{n}=-\frac{ic_{\rm dS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right).$ (44)
+In this case, the argument of the inverse hyperbolic function is greater than one, making $\tanh^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)$ a complex value.
 The positive branch of the inverse hyperbolic function is chosen to match the real part of the entanglement entropy in the dS/CFT.
 The final expression for the entanglement entropy in the dS/CFT correspondence is
-$S_{A}=-\frac{ic_{\rm dS}}{3}\mathrm{arccoth}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6},$ (48)
+$S_{A}=-\frac{ic_{\rm dS}}{3}\coth^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6},$ (45)
 where the real part agrees with the results in refs.
-Hikida:2022ltr ; Doi:2023zaf ; Chen:2023eic .
+Hikida:2022ltr ; Doi:2023zaf ; Narayan:2022afv ; Narayan:2023zen .
 As discussed in ref.
@@
 However, due to the non-unitarity of the dual field theory, the $C$-function is also a purely imaginary number.
-$C(r)=3rS_{A}^{\prime}(r)=-ic_{\rm dS}\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\,.$ (49)
-For the undeformed CFT, we have $C(r\to\infty)=-ic_{\rm dS}=c$. But when $r\to\ell_{\rm dS}$, the imaginary part of the $C$-function diverges, as shown in figure 6.
+$C(r)=-ic_{\rm dS}\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}}}\,.$ (46)
+For the undeformed CFT, we have $C(r\to\infty)=-ic_{\rm dS}=c$. But as $r\to\ell_{\rm dS}$, the imaginary part of the $C$-function diverges, as illustrated in figure 5, as anticipated.
 (a) $C(r)$
 (b) $C(\lambda)$
-Figure 6: The $C$-function of the dS spacetime at $r>\ell_{\rm dS}$, and $\lambda=\frac{\lambda_{\rm dS}}{r^{2}}$. The red dashed line represents $r=\ell_{\text{dS}}$.
-In the dS/dS case, the entanglement entropy and $C$-function have been analyzed in refs. Gorbenko:2018oov ; Dong:2018cuv , with the results given as
-$S_{A}=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}\right),\quad C(r)=3rS_{A}^{\prime}(r)=c_{\rm dS}\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}.$ (50)
-It is noted that when the effective radius is expressed as $r=\ell_{\rm dS}\sin w<\ell_{\rm dS}$, these results can be directly obtained through analytic continuation from the corresponding entanglement entropy and $C$-function in the dS/CFT correspondence.
-The $C$-function does decrease monotonically as $r$ decreases, and it still diverges at $r=\ell_{\rm dS}$, as shown in figure 7.
+Figure 5: The $C$-function of the dS spacetime at $r>\ell_{\rm dS}$, and $\lambda=\frac{\lambda_{\rm dS}}{r^{2}}$. The red dashed line represents $r=\ell_{\text{dS}}$, that is $t_{c}=0$.
+dS/dS:
+In the dS/dS holography, the entanglement entropy and $C$-function have been analyzed in refs. Gorbenko:2018oov , with the results given as
+$S_{A}=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}\right),\quad C(r)=c_{\rm dS}\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}.$ (47)
+These two results can also be directly derived from expressions (44) and (46) within the dS/CFT duality by recognizing that the sphere’s radius is smaller than the de Sitter radius $\ell_{\rm dS}$.
+As anticipated, the $C$-function diverges at the ultraviolet boundary $w_{c}=\frac{\pi}{2}$, as illustrated in figure 6.
 (a) $C(r)$
 (b) $C(\lambda)$
-Figure 7: The $C$-function of the dS spacetime at $r<\ell_{\rm dS}$.
-By calculating the second derivative of the entanglement entropy with respect to the sphere radius $r$, it can be found that when the dual field theory resides on a timelike boundary, the entanglement entropy violates the strong subadditivity inequality,
-$r^{2}\partial^{2}_{r}S_{A}(r)=\frac{1}{\left(\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1\right)^{\frac{3}{2}}}>0.$ (51)
-Therefore, it implies that the dual field theory is non-local Kawamoto:2023nki ; Geng:2019ruz when the effective radius of the sphere is smaller than the cosmological horizon.
+Figure 6: The $C$-function of the dS spacetime at $r<\ell_{\rm dS}$. The red dashed line indicates the ultraviolet boundary at $w_{c}=\frac{\pi}{2}$.
+Half-dS holography:
+Recently, a novel holographic framework, half-dS holography, has been proposed Kawamoto:2023nki .
+This model introduces a timelike boundary into global dS spacetime.
+The induced metric on this boundary is given by
+$ds^{2}=\ell_{\rm dS}^{2}(-dt^{2}+\cosh^{2}t\sin^{2}\theta_{0}d\phi^{2}).$ (48)
+The dual field theory is defined on this timelike boundary, characterized by the condition $\theta=\theta_{0}$.
+To ensure that spacelike geodesics can probe deep into the bulk spacetime, the bulk region is restricted to $0<\theta<\theta_{0}$, as indicated by the shaded area in figure 7.
+Reference Kawamoto:2023nki proposes that the dual field theory in this model also exhibits non-locality.
+Specifically, when the timelike boundary is located at $\theta_{0}=\frac{\pi}{2}$, the subadditivity is consistently violated for $t\neq 0$, except for the subadditivity saturation at $t=0$.
+This position coincides with the ultraviolet boundary position $w_{c}=\frac{\pi}{2}$ in dS/dS holography, which corresponds to the black dashed line in figure 7 and the blue dashed line in figure 4(b), respectively.
+Figure 7: Half-dS holography.
+Here, we set $\cos T=\frac{1}{\cosh t}$.
+The dual field theory is defined on the timelike boundary $\theta=\theta_{0}$, and the bulk spacetime is represented by the shaded region.
+The dual field theories in dS/dS holography and half-dS holography are both defined on timelike boundaries within dS spacetime and exhibit non-locality at the central slice.
+To this end, we will extend the replica method for calculating entanglement entropy between antipodal points on a sphere to compute the entanglement entropy between any spatial region and its complement on a general manifold in Section 4.
+Subsequently, we will apply the replica method to dS/dS holography and half-dS holography in Section 5 to discuss the similarities and differences between these two holographic models.
 4 Entanglement entropy for arbitrary single spatial intervals
@@
 This method applies only to subsystems along semicircular arcs between antipodal points.
-To analyze the entanglement entropy for subsystems in arbitrary single spatial intervals on a two-dimensional plane, ref.
-Deng:2023pjs introduced the CHM mapping, which transforms the subsystem into a semicircular arc on a two-dimensional sphere.
-The entanglement entropy can then be determined using the results obtained for the spherical case.
-For subsystems on more general manifolds, they can first be mapped onto the two-dimensional plane, after which the method from ref.
-Deng:2023pjs can be applied.
-The replica method is generalized in the following section to calculate the entanglement entropy for arbitrary single spatial intervals on the two-dimensional plane with $T\bar{T}$ deformation and extended to general manifolds, including cylinders and spheres.
+While the entanglement entropy of a single interval subjected to $T\bar{T}$ deformation on a planar manifold can be determined through CHM mapping and endpoint integration of the entanglement region Lewkowycz:2019xse ; Deng:2023pjs , a comprehensive calculation of the entanglement entropy for an arbitrary single interval under $T\bar{T}$ deformation on spherical and cylindrical manifolds remains absent [^4] In ref.
+Grieninger:2019zts , the authors proposed an entanglement entropy expression for a general interval, formatted analogously to the entanglement entropy expression between antipodal points on a spherical manifold, by introducing an effective radius.
+However, a rigorous proof of this proposition is lacking..
+In this section, we will extend the replica method to general manifolds, including spherical and cylindrical manifolds, to compute the $T\bar{T}$ corrections to the entanglement entropy for any arbitrary single interval.
 4.1 The generalization of the replica method
-The goal is to calculate the entanglement entropy of the $T\bar{T}$-deformed quantum field theory in a spatial subregion $A$ on an arbitrary two-dimensional manifold $\mathcal{M}$, where $A$ is defined by $t=0$ and $x\in[x_{a},x_{b}]$ in a chosen coordinate system.
-When $\mathcal{M}$ is a two-dimensional sphere and $A$ is specified by $\phi=\text{constant}$ and $\theta\in(0,\pi)$, the entanglement entropy can be computed using eqs.
-(29) and (30).
-The entanglement entropy of $A$ can be obtained through the $C$-function. Differentiating $S_{A}$ with respect to $r$, we have
-$r\frac{d}{dr}S_{A}(r)$ $=r\frac{d}{dr}\lim_{n\rightarrow 1}(1-n\partial_{n})\ln Z_{n}=\lim_{n\rightarrow 1}(1-n\partial_{n})r\frac{d}{dr}\ln Z_{n}$
-$=-\lim_{n\rightarrow 1}(1-n\partial_{n})2\pi nr^{2}\int_{0}^{\pi}\sin\theta d\theta\,\text{tr}\,T$
-$=2\pi r^{2}\lim_{n\rightarrow 1}n^{2}\int_{0}^{\pi}\sin\theta d\theta\,\partial_{n}\text{tr}\,T,$ (52)
-where, in the last two steps, it is assumed that the metric (27) is independent of the azimuthal angle $\phi$, and thus $T_{ab}$ is also independent of $\phi$.
-Solving the trace flow and conservation equations without assuming spherical symmetry must determine the energy-momentum tensor $T_{ab}$.
-For the $n$-replica sphere (27), the solution for the energy-momentum tensor is known Lewkowycz:2019xse and is given by
-${T^{\theta}}_{\theta}$ $=\frac{1}{\pi\lambda}\left(1-\sqrt{1+\frac{\lambda c}{12r^{2}}+\frac{\lambda c}{12r^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}\right),$
-${T^{\phi}}_{\phi}$ $=\frac{1}{\pi\lambda}\left(1-\frac{1+\frac{\lambda c}{12r^{2}}}{\sqrt{1+\frac{\lambda c}{12r^{2}}+\frac{\lambda c}{12r^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}}\right).$ (53)
-Substituting this into the integral in eq. (52), in the limit $n\rightarrow 1$, the integrand becomes
-$n\partial_{n}\text{tr}\,T=\frac{c}{3}\frac{\lambda c}{48\pi}\frac{1-n^{2}}{n^{4}r^{4}\sin^{4}\theta}\frac{1}{\left[1+\frac{\lambda c}{12r^{2}}\left(1+\frac{2(1-n)}{\sin^{2}\theta}\right)\right]^{3/2}},$ (54)
-which tends to zero for $\theta\neq 0,\pi$.
-However, near the branch points, the integrand has a non-zero contribution.
-Thus, the main contribution to the integral in eq.
-(52) comes from small neighborhoods near the branch points
-$rS^{\prime}(r)=2\pi r^{2}\lim_{n\rightarrow 1}\left(\int_{\theta\in B(0,\epsilon)}+\int_{\theta\in B(\pi,\epsilon)}\right)\sin\theta d\theta\,\partial_{n}\text{tr}\,T,$ (55)
-where $B(p,\epsilon)$ is a small neighborhood centered at $p$ with radius $\epsilon$.
-For a general manifold $\mathcal{M}$, we first map $\mathcal{M}$ to $\mathbb{S}^{2}$, ensuring that the image of the subsystem $A$ under this mapping is a meridian on $\mathbb{S}^{2}$.
-The mapped metric is conformally equivalent to the sphere, denoted as $\Omega\mathbb{S}^{2}$.
-After applying the $n$-replica trick, we assume the metric takes the form
-$ds^{2}=r^{2}\Omega^{2}(\theta,\phi)\left(d\theta^{2}+n^{2}\sin^{2}\theta\,d\phi^{2}\right),$ (56)
-where the conformal factor $\Omega(\theta,\phi)$ is constant at the branch points $\theta=0$ and $\theta=\pi$, denoted as $\Omega_{0}$ and $\Omega_{\pi}$, respectively.
-For this metric, eqs.
-(29) and (30) remain valid
-$\frac{d}{dr}\ln Z_{n}$ $=-\int_{\Omega\mathbb{S}_{n}^{2}}d^{2}x\frac{\delta\mathcal{L}}{\delta g_{ab}}\frac{dg_{ab}}{dr}=-\frac{1}{r}\int_{\Omega\mathbb{S}_{n}^{2}}d^{2}x\sqrt{g}T^{ab}g_{ab}.$ (57)
-Thus, we have
-$r\frac{d}{dr}S(r)$ $=r\frac{d}{dr}\lim_{n\rightarrow 1}(1-n\partial_{n})\ln Z_{n}=\lim_{n\rightarrow 1}(1-n\partial_{n})r\frac{d}{dr}\ln Z_{n}$
-$=-\lim_{n\rightarrow 1}(1-n\partial_{n})nr^{2}\int_{\Omega\mathbb{S}_{n}^{2}}\Omega^{2}(\theta,\phi)\sin\theta\,d\theta d\phi\,\text{tr}\,T$
-$=r^{2}\lim_{n\rightarrow 1}n^{2}\int_{\Omega\mathbb{S}_{n}^{2}}\Omega^{2}(\theta,\phi)\sin\theta\,d\theta d\phi\,\partial_{n}\text{tr}\,T.$ (58)
-Since the conformal factor $\Omega(\theta,\phi)$ depends on the azimuthal angle $\phi$, $T_{ab}$ cannot be assumed to depend solely on $\theta$.
-Therefore, the $\phi$ coordinate is not integrated in the final two steps.
-In standard CFT calculations, a conformal transformation is often applied to map $\Omega\mathbb{S}^{2}$ to the spherical metric, and conformal symmetry ensures that the results for $\Omega\mathbb{S}^{2}$ and $\mathbb{S}^{2}$ differ only by a function of the conformal factor.
-However, for the $T\bar{T}$-deformed field theory, conformal symmetry is absent, and the relationship between quantities before and after the conformal transformation is not determined.
-As discussed in ref.
-Lewkowycz:2019xse , for the partition function on the sphere (52), in the limit $n\rightarrow 1$, the main contribution to the integral (52) comes from a small neighborhood near the branch points, with radius $\epsilon\sim\sqrt{n-1}$.
-This property is expected to hold for any manifold $\mathcal{M}$, so the integral (58) is also performed only in small neighborhoods near the branch points.
-In these regions, the conformal factor $\Omega$ is constant, and the metric (56) can be approximated by the $n$-replica sphere with radius $r\Omega$.
-As a result, the conformal transformation can be regarded as an identity transformation.
-Therefore, partition functions on any manifold $\mathcal{M}$ can be calculated using the results on $\mathbb{S}^{2}$ for non-conformal field theories.
-Finally, we can get
-$rS^{\prime}(r)$ $=-\lim_{n\rightarrow 1}(1-n\partial_{n})\int_{\Omega\mathbb{S}_{n}^{2}}d^{2}x\sqrt{g}\,\text{tr}\,T$
-$=\lim_{n\rightarrow 1}2\pi(r\Omega)^{2}\left(\int_{0}^{\theta_{0}}+\int_{\pi-\theta_{0}}^{\pi}\right)\sin\theta\,d\theta\,\partial_{n}\text{tr}\,T,$ (59)
-where $\text{tr}\,T$ corresponds to the solution on the sphere with radius $r\Omega$.
-When Poincaré coordinates are chosen in EAdS3, the holographic boundary corresponds to a two-dimensional complex plane with the metric
-$ds_{2}^{2}=\frac{\ell_{\rm AdS}^{2}}{z_{c}^{2}}(dT^{2}+dX^{2}).$ (60)
-The subregion $A$ is defined by $T=0$ and $X\in(-X_{a},X_{a})$.
-To compute the entanglement entropy for $A$, the complex plane is mapped onto a two-dimensional sphere such that the image of $A$ corresponds to the meridian on the sphere at $\phi=0$, as illustrated in figure 8.
-This mapping, known as the CHM mapping, is given by
-$X=\frac{X_{a}\cos\theta}{1+\sin\theta\cos\phi},\quad T=\frac{X_{a}\sin\theta\sin\phi}{1+\sin\theta\cos\phi}.$ (61)
-Under this transformation, the boundary plane in $(\theta,\phi)$ coordinates has the metric
-$ds_{2}^{2}=\frac{\ell_{\rm AdS}^{2}X_{a}^{2}}{z_{c}^{2}(1+\cos\phi\sin\theta)^{2}}(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}),$ (62)
-which is conformally equivalent to a sphere with an effective radius $r_{\mathrm{eff}}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}$, and the conformal factor $\Omega(\theta,\phi)=(1+\cos\phi\sin\theta)^{-1}$.
+In ref.
+Lewkowycz:2019xse , it is proposed that in the $T\bar{T}$-deformed theory, all dependence of the entanglement entropy on $\lambda$ originates from the contributions of the endpoints of the entanglement region.
+For the reader’s convenience, we provide a concise review of the relevant derivation in Appendix A and apply it to a planar manifold via the CHM map, yielding the expression for the entanglement entropy of a single interval in AdS Poincaré spacetime.
+Drawing inspiration from this, we can employ analogous steps in dS Poincaré spacetime to derive the entanglement entropy expression for a single spatial interval in the $T\bar{T}$-deformed CFT.
+For dS Poincaré spacetime, the metric is given by
+$ds^{2}=\frac{\ell_{\rm dS}^{2}}{\eta^{2}}(-d\eta^{2}+dX^{2}+dT^{2})\,.$ (49)
+In this context, the $T\bar{T}$-deformed CFT resides on the two-dimensional complex plane $\eta=\eta_{c}$.
+The subregion $A$ is defined by $X\in[-X_{a},X_{a}]$ at $T=0$.
+It is mapped via the CHM transformation to a meridian $\Phi=0$ on a sphere with an effective radius $r_{\rm eff}=\frac{\ell_{\rm dS}X_{a}}{\eta_{c}}$, as illustrated in figure 8.
 Figure 8: Mapping from the plane to the sphere.
 The north and south poles of the sphere correspond to the endpoints of the interval $A$ on the complex plane.
-The subregion $A$ corresponds to the meridian at $\phi=0$, while the locally time-evolved subregions (shown as blue or yellow lines) correspond to other meridians on the sphere.
-At this stage, the energy-momentum tensor for the metric (62) must be determined using the equations in (31), followed by evaluating the integral in (52).
-As previously noted, when computing the partition function on a general manifold $\mathcal{M}$, the primary contribution to the integral (52) comes from an infinitesimal neighborhood near the branch points.
-In these regions, the conformal factor $\Omega(\theta,\phi)$ can be approximated as 1, implying that the metric (62) effectively describes a sphere with radius $r_{\mathrm{eff}}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}$.
-Thus, within the infinitesimal neighborhoods around the endpoints, transforming (60) to the sphere involves only a conformal transformation with a conformal factor of 1, effectively an identity transformation.
-Consequently, the partition function remains the same as that of the sphere, with the radius $r$ replaced by the effective radius $r_{\mathrm{eff}}$.
-For the AdS/CFT, the deformation parameter $\lambda$ and the central charge $c$ in eq.
-(31) are $c_{\rm AdS}=\frac{3\ell_{\rm AdS}}{2G}$ and $\lambda_{\rm AdS}=8G\ell_{\rm AdS}$, respectively.
-Therefore, the solution for the energy-momentum tensor on the $n$-replica sphere is
-${T^{\theta}}_{\theta}$ $=\frac{1}{\pi\lambda_{\rm AdS}}\left(1-\sqrt{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}\right),$
-${T^{\phi}}_{\phi}$ $=\frac{1}{\pi\lambda_{\rm AdS}}\left(1-\frac{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}}{\sqrt{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}}\right)\,,$ (63)
-where the effective radius is $r_{\rm eff}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}$. Substituting this into eq. (52) and performing the integration near the endpoints yields the $C$-function as
-$C(r)=\frac{c_{\rm AdS}}{3}\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\rm eff}^{2}}}}\,.$ (64)
-Using the $C$-function, one can get the entanglement entropy of the subregion A as
-$S_{A}=\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\rm eff}^{2}}}}\right)\simeq\frac{c_{\mathrm{AdS}}}{3}\log\frac{2X_{a}}{z_{c}}+\frac{c_{\mathrm{AdS}}}{12}\frac{z_{c}^{2}}{X_{a}^{2}}.$ (65)
-In this expression, the first term represents the entanglement entropy of the undeformed field theory in the interval $(-X_{a},X_{a})$ on the plane.
-The second term accounts for the correction introduced by the $T\bar{T}$ deformation.
-This result is consistent with the findings reported in ref.
-He:2023xnb .
-For the dS/CFT, the deformation parameter $\lambda$ and the central charge $c$ in eq.
-(31) are given by $c\equiv-ic_{\rm dS}=-i\frac{3\ell_{\rm dS}}{2G}$ and $\lambda\equiv-i\lambda_{\rm dS}=-i8G\ell_{\rm dS}$, respectively.
-The solution for the energy-momentum tensor on the $n$-replica sphere
+The subregion $A$ corresponds to the meridian at $\Phi=0$.
+The solution for the energy-momentum tensor on the $n$-replica sphere is given by
 $T^{\Theta}_{\Theta}$ $=\frac{i}{\pi\lambda_{\rm dS}}\left(1-\sqrt{1-\frac{\lambda_{\rm dS}c_{\rm dS}}{12r_{\rm eff}^{2}}-\frac{\lambda_{\rm dS}c_{\rm dS}}{12r_{\rm eff}^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\Theta}}\right),$
 $T^{\Phi}_{\Phi}$ $=\frac{i}{\pi\lambda_{\rm dS}}\left(1-\frac{1-\frac{\lambda_{\rm dS}c_{\rm dS}}{12r_{\rm eff}^{2}}}{\sqrt{1-\frac{\lambda_{\rm dS}c_{\rm dS}}{12r_{\rm eff}^{2}}-\frac{\lambda_{\rm dS}c_{\rm dS}}{12r_{\rm eff}^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\Theta}}}\right)\,,$ (50)
-where the effective radius is $r_{\rm eff}=\frac{\ell_{\rm dS}X_{a}}{\eta_{c}}$. Substituting this into eq. (52), we obtain the $C$-function is
-$C(r)=-\frac{ic_{\rm dS}}{3}\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}}}\,.$ (67)
-The entanglement entropy is then given by
-$S_{A}=-\frac{ic_{\rm dS}}{3}\mathrm{arccoth}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6}\simeq-\frac{ic_{\rm dS}}{3}\log\frac{2X_{a}}{\eta_{c}}+\frac{ic_{\rm dS}\eta_{c}^{2}}{12X_{a}^{2}}+\frac{\pi c_{\rm dS}}{6}\,.$ (68)
-In Appendix A.1, this quantity is computed using holographic methods by evaluating the area of the RT surface with a finite cutoff $\eta_{c}$. The result is consistent with eq. (68).
-To generalize the method for computing subsystems’ entanglement entropy on arbitrary holographic boundaries, we note that any two-dimensional manifold is locally conformally equivalent to the plane.
+where the effective radius is $r_{\rm eff}=\frac{\ell_{\rm dS}X_{a}}{\eta_{c}}$. So we can obtain the $C$-function of the dS Poincaré spacetime
+$C(r)=-\frac{ic_{\rm dS}}{3}\frac{1}{\sqrt{1-\frac{\eta_{c}^{2}}{X_{a}^{2}}}}\,.$ (51)
+Finally, the entanglement entropy is given by
+$S_{A}=-\frac{ic_{\rm dS}}{3}\coth^{-1}\left(\frac{1}{\sqrt{1-\frac{\eta_{c}^{2}}{X_{a}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6}\simeq-\frac{ic_{\rm dS}}{3}\log\frac{2X_{a}}{\eta_{c}}+\frac{ic_{\rm dS}\eta_{c}^{2}}{12X_{a}^{2}}+\frac{\pi c_{\rm dS}}{6}\,.$ (52)
+In Appendix B.1, we obtain consistent results by computing the area of the RT surface in the bulk spacetime with a finite cutoff $\eta_{c}$.
+To generalize the replica method for arbitrary holographic boundaries, we note that any two-dimensional manifold is locally conformally equivalent to the plane.
 Thus, its line element can be expressed as
@@
 Within these small neighborhoods, we can use the solution for the energy-momentum tensor on the $n$-sphere to compute the integral, with the effective radius in the solutions (124) or (50) replaced by $r_{\mathrm{eff}}=\omega(T_{a},X_{a})X_{a}$.
-If $\omega(T,X)$ takes different constant values at the two endpoints, denoted as $\omega_{a}$ and $\omega_{-a}$, respectively.
-The contributions to the integral (52) from different endpoints correspond to spheres with radii $\omega_{a}X_{a}$ and $\omega_{-a}X_{a}$.
-Therefore, these integrals should be replaced by the partition function on spheres with radii $\omega_{a}X_{a}$ and $\omega_{-a}X_{a}$, respectively.
+If $\omega(T,X)$ takes distinct constant values at the two endpoints, denoted $\omega_{a}$ and $\omega_{-a}$, respectively, the contributions to the integral (113) from these endpoints correspond to spheres with radii $\omega_{a}X_{a}$ and $\omega_{-a}X_{a}$.
+These integrals are therefore replaced by the partition function on spheres with radii $\omega_{a}X_{a}$ and $\omega_{-a}X_{a}$, respectively.
 4.2 System with finite size
-In the previous subsection, we reviewed the CHM mapping method for calculating the entanglement entropy of the $T\bar{T}$-deformed CFT in spatial intervals on a two-dimensional plane.
-Specifically, we computed the entanglement entropy for a subsystem on the boundary of the dS spacetime in Poincaré coordinates.
-We now extend this method to two-dimensional cylindrical and spherical surfaces to examine whether the $T\bar{T}$ deformation introduces corrections to the entanglement entropy for finite-size systems.
-The corresponding holographic models for these geometries are AdS and dS spacetimes in global coordinates, respectively.
-The entanglement entropy calculation for general spatial intervals in curved space has been discussed in ref.
-Grieninger:2019zts .
+In the preceding subsection, we reviewed the CHM mapping method for computing the entanglement entropy of $T\bar{T}$-deformed CFTs within spatial intervals on a two-dimensional plane.
+Specifically, we calculated the entanglement entropy of a subsystem in dS Poincaré spacetime.
+We now extend this methodology to two-dimensional cylindrical and spherical manifolds to investigate whether the $T\bar{T}$ deformation induces corrections to the entanglement entropy in finite-size systems.
 4.2.1 AdS case
@@
 The resulting entanglement entropy for the subregion $A$ is given by
-$S_{A}=$ $\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\rm eff}^{2}}}}\right)$
+$S_{A}=$ $\frac{c_{\rm AdS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\rm eff}^{2}}}}\right)$
 $\simeq$ $\frac{c_{\rm AdS}}{3}\log(2\sinh\rho_{c}\sin x_{a})+\frac{c_{\rm AdS}}{12\sinh^{2}\rho_{c}\sin^{2}x_{a}}\,,$ (62)
@@
 $ds^{2}=\ell_{\rm dS}^{2}(-dt^{2}+\cosh^{2}t\,(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}))\,.$ (63)
-After the $T\bar{T}$ deformation, its holographic boundary is a 2-dimensional sphere determined by $t=t_{c}$.
-The subregion $A$ is defined by the arc $\Theta=\pi/2$, $\Phi\in(-\Phi_{a},\Phi_{a})$ on this 2-dimensional sphere.
+After the $T\bar{T}$ deformation, its holographic boundary is a two-dimensional sphere determined by $t=t_{c}$.
+The subregion $A$ is defined by the arc $\theta=\pi/2$, $\phi\in(-\phi_{a},\phi_{a})$ on this two-dimensional sphere.
 As done previously, we begin by mapping the sphere to the complex plane using the transformation
-$\Theta=2\arctan\sqrt{T^{2}+X^{2}}\,,\quad\Phi=\frac{1}{2i}\ln\frac{T+iX}{T-iX}\,.$ (80)
+$\theta=2\tan^{-1}\sqrt{T^{2}+X^{2}}\,,\quad\phi=\frac{1}{2i}\ln\frac{T+iX}{T-iX}\,.$ (64)
 In this coordinate system, the endpoints of $A$ are $(\cos\phi_{a},-\sin\phi_{a})$ and $(\cos\phi_{a},\sin\phi_{a})$, and the line element becomes
@@
 The transformed line element is
-$ds^{2}$ $=\frac{\ell_{\rm dS}^{2}\cosh^{2}t_{c}\sin^{2}\Phi_{a}}{(1+\cos(\phi-\Phi_{a})\cos(\Phi_{a})\sin(\theta))^{2}}(d\theta^{2}+\sin^{2}\theta d\phi^{2})$
+$ds^{2}$ $=\frac{\ell_{\rm dS}^{2}\cosh^{2}t_{c}\sin^{2}\phi_{a}}{(1+\cos(\Phi-\phi_{a})\cos\phi_{a}\sin\Theta)^{2}}(d\Theta^{2}+\sin^{2}\Theta d\Phi^{2})$
 $=\Omega(\Theta,\Phi)^{2}r_{\rm eff}^{2}(d\Theta^{2}+\sin^{2}\Theta d\Phi^{2})\,,$ (67)
 where $r_{\rm eff}=\ell_{\rm dS}\cosh t_{c}\sin\phi_{a}$ is the effective radius of the sphere, and the conformal factor is
-$\Omega(\theta,\phi)=\frac{1}{1+\cos(\phi-\Phi_{a})\cos(\Phi_{a})\sin(\theta)}\,.$ (84)
+$\Omega(\Theta,\Phi)=\frac{1}{1+\cos(\Phi-\phi_{a})\cos\phi_{a}\sin\Theta}\,.$ (68)
 At the antipodal points $\Theta=0,\pi$, the conformal factor $\Omega(\Theta,\Phi)=1$, allowing us to directly use the results in eq. (50) to calculate the entanglement entropy. The final result is
-$S_{A}$ $=-\frac{ic_{\rm dS}}{3}\mathrm{arccoth}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6}$
+$S_{A}$ $=-\frac{ic_{\rm dS}}{3}\coth^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6}$
 $\simeq-\frac{ic_{\rm dS}}{3}\log(2\cosh t_{c}\sin\phi_{a})+\frac{ic_{\rm dS}}{12\cosh^{2}t_{c}\sin^{2}\phi_{a}}+\frac{\pi c_{\rm dS}}{6}.$ (69)
-This result is consistent with the area of the RT surface calculated in Appendix A.2, and also matches the findings of ref. Chen:2023eic when $\Phi_{a}=\frac{\pi}{2}$.
+This result is consistent with the area of the RT surface calculated in Appendix B.2, and also matches the findings of ref. Chen:2023eic when $\phi_{a}=\frac{\pi}{2}$.
 4.3 System with finite temperature
@@
 Integrating the above function with respect to $r$, the entanglement entropy is given by
-$S_{A}=\frac{c_{\rm AdS}}{6}\left[\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{{\rm eff},\theta=0}^{2}}}}\right)+\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{{\rm eff},\theta=\pi}^{2}}}}\right)\right]\,.$ (96)
+$S_{A}=\frac{c_{\rm AdS}}{6}\left[\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{{\rm eff},\Theta=0}^{2}}}}\right)+\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{{\rm eff},\Theta=\pi}^{2}}}}\right)\right]\,.$ (80)
 Substituting the radii $r_{{\rm eff},\Theta}$ into this expression and expanding the result to second order in $R_{0}/R_{c}$,
@@
 where we have replaced $R_{0}$ with $\frac{2\pi\ell_{\rm AdS}^{2}}{\beta}$.
-In the high-temperature limit $\beta\ll 1$, the last term is negligible compared to the second term.
-This result is consistent with the findings in refs.
-Chen:2018eqk ; Apolo:2023ckr by identifying $R_{c}^{2}=\frac{6\ell_{\rm AdS}^{2}}{\mu\pi c_{\rm AdS}}$.
-5 (A)dS/dS case and non-locality
-In this section, we apply the methods from the previous section to the AdS/dS, dS/dS, and half-dS holography cases.
-For all three, the holographic boundary is a timelike hypersurface.
-In dS/dS and half-dS holography, previous studies Kawamoto:2023nki ; Geng:2019ruz suggest that the dual field theory may exhibit non-locality.
-Unlike the examples discussed earlier, the boundary-induced metrics in this section are time-dependent.
-As a result, the properties of entanglement entropy may vary significantly over time, potentially influencing the determination of the strong subadditivity or the boosted strong subadditivity.
-This issue will be discussed in detail in this section.
-5.1 AdS/dS
-For the AdS/dS correspondence, the bulk metric is expressed as
-$ds^{2}=\ell_{\rm AdS}^{2}(dw^{2}+\sinh^{2}w(-dt^{2}+\cosh^{2}tdx^{2})),$ (98)
-where $x\sim x+2\pi$, and the holographic boundary is determined by $w=w_{c}\to\infty$.
-We choose the boundary subregion $A$ as $x\in(-x_{a},x_{a})$ at $t=t_{0}$.
-In the Euclidean signature, after rescaling the time coordinate, the boundary metric can be written as
-$ds^{2}=\ell_{\rm AdS}^{2}\frac{\sinh^{2}w_{c}}{\cosh^{2}\tilde{t}}\left(d\tilde{t}^{2}+dx^{2}\right)\,,$ (99)
-where $\tilde{t}$ is related to $t$ by the relation
-$\tilde{t}=\ln\frac{\cot\frac{t}{2}+1}{\cot\frac{t}{2}-1}\,.$ (100)
-This metric is conformally equivalent to the cylinder. As we have done in the last section, we can map it to the complex plane through the following transformations
-$\tilde{t}\rightarrow\frac{1}{2}\ln(T^{2}+X^{2})\,,\quad x\rightarrow\frac{1}{2i}\ln\left(\frac{T+iX}{T-iX}\right)\,.$ (101)
-and the resulting conformal flat metric is
-$ds^{2}=\ell_{\rm AdS}^{2}\sinh^{2}w_{c}\frac{4\left(dT^{2}+dX^{2}\right)}{\left(1+T^{2}+X^{2}\right)^{2}}\,.$ (102)
-The two endpoints of the subregion $A$ in the $(T,X)$ coordinates are
-$(T_{1},X_{1})=(e^{\tilde{t}_{0}}\cos x_{a},-e^{\tilde{t}_{0}}\sin x_{a}),\quad(T_{2},X_{2})=(e^{\tilde{t}_{0}}\cos x_{a},e^{\tilde{t}_{0}}\sin x_{a})\,.$ (103)
-Using the CHM map,
-$T\rightarrow\frac{e^{\tilde{t}_{0}}\sin x_{a}\sin\theta\sin\phi}{1+\sin\theta\cos\phi}+e^{\tilde{t}_{0}}\cos x_{a}\,,\quad X\rightarrow\frac{e^{\tilde{t}_{0}}\sin x_{a}\cos\theta}{1+\sin\theta\cos\phi}\,,$ (104)
-this complex plane can be mapped onto the sphere, yielding the conformal spherical metric
-$ds^{2}=\frac{4e^{2\tilde{t}_{0}}\ell_{\rm AdS}^{2}\sinh^{2}w_{c}\sin^{2}x_{a}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)}{\left(1+\sin\theta\cos\phi+e^{2\tilde{t}_{0}}\sin\theta\cos(\phi-2x_{a})+e^{2\tilde{t}_{0}}\right)^{2}}\,.$ (105)
-It is easy to verify that the conformal factor at the endpoints $\theta=0,\pi$ is the same constant, which can be absorbed into $r_{\rm eff}$. Therefore, the effective radius is
-$r_{\text{eff}}=\ell_{\rm AdS}\sinh w_{c}\cos t_{0}\sin x_{a}\,,$ (106)
-or, in Lorentz signature, $r_{\text{eff}}=\ell_{\rm AdS}\sinh w_{c}\cosh t_{0}\sin x_{a}$. Finally, the entanglement entropy of the subsystem is
-$S_{A}\simeq\frac{c_{\rm AdS}}{3}\log(2\sinh w_{c}\cosh t_{0}\sin x_{a})+\frac{c_{\rm AdS}}{12\sinh^{2}w_{c}\cosh^{2}t_{0}\sin^{2}x_{a}}\,.$ (107)
-If we choose the UV cutoff as $\sinh w_{c}=1/\epsilon$, we find that eq.(107) agrees with the result in ref.
-Kawamoto:2023nki .
-Alternatively, if we select the subregion as $t=t_{0}$, $x\in[-\frac{\pi}{2},\frac{\pi}{2}]$, the entanglement entropy is $S\simeq\frac{c_{\rm AdS}}{3}w_{c}$, which matches the result in ref.
-Donnelly:2018bef .
-For verification, we holographically calculate the area of the RT surface, which is expressed as
-$D_{12}$ $=\ell_{\rm AdS}\mathrm{arccosh}(1+2\sinh^{2}w_{c}\cosh^{2}t_{0}\sin^{2}x)$
-$\simeq 2\ell_{\rm AdS}\log(2\sinh w_{c}\cosh t_{0}\sin x_{a})+\frac{\ell_{\rm AdS}}{2\sinh^{2}w_{c}\cosh^{2}t_{0}\sin^{2}x_{a}}.$ (108)
-After dividing $D_{12}$ by $4G$ and substituting the relation $G=3\ell_{\rm AdS}/2c_{\rm AdS}$, it is consistent with (107) up to the second-order correction of the $T\bar{T}$ deformation.
-5.2 dS/dS
+In the high-temperature limit $\beta\ll 1$, the last term is negligible compared to the second term [^5] This result is consistent with the findings in refs.
+Chen:2018eqk ; Apolo:2023ckr by identifying $R_{c}^{2}=\frac{6\ell_{\rm AdS}^{2}}{\mu\pi c_{\rm AdS}}$..
+5 dS/dS holography and half-dS holography
+In this section, we apply the replica method to dS/dS holography and half-dS holography.
+In contrast to the previously discussed examples, here, all boundary-induced metrics are dS metrics, and the dual field theory resides on dS spacetime.
+Previous studies on dS/dS holography and half-dS holography Kawamoto:2023nki ; Geng:2019ruz suggest that the dual field theory exhibits non-locality.
+In this section, we will provide a more detailed discussion of entanglement entropy and non-locality, building upon these prior investigations.
+Typically, the locality of a field theory is examined through strong subadditivity (SSA) and boosted strong subadditivity (BSSA).
+In Appendix C, we provide a concise review of the derivation of SSA and BSSA in flat spacetime.
+However, since the dual field theory in dS/dS holography and half-dS holography resides on dS space, we need to account for the effects of curved spacetime.
+The causal structure of dS spacetime exhibits a distinct geometry compared to the Minkowski causal diamond, owing to the metric’s inherent time evolution, as illustrated in figure 12.
+Figure 12: The causal structure of dS spacetime. The interval size we choose is $\Delta\phi=\frac{98\pi}{100}$.
+In dS spacetime, since the SSA does not consider boost transformations, the intervals $A$ and $B$ remain on the $t=0$ slice.
+Consequently, the SSA is unaffected by the curvature of spacetime.
+Thus, the SSA in dS space is
+$S^{\prime\prime}(\Delta\phi)\leq 0.$ (82)
+When deriving the BSSA, we must consider a small boost transformation of intervals $A$ and $B$.
+In figure 12, we represent interval $A$ with red segments, interval $B$ with blue segments, $A\cap B$ with orange segments, and $A\cup B$ with green segments.
+Considering that the entanglement entropy is a function of the interval size, we use the geodesic distance between two points as the interval length.
+We set the length of interval $A\cup B$ as $L_{A\cup B}=\Delta\phi$, and consider a small boost transformation that shrinks $\delta\phi$ in the $\phi$ direction.
+Then, the length of interval $A\cap B$ is
+$L_{A\cap B}=\cos^{-1}\left(1-2\sin^{2}\left(\frac{\Delta\phi}{2}-\delta\phi\right)\cosh\left(2\tanh^{-1}\left(\tan\frac{\delta\phi}{2}\right)\right)\right)\,.$ (83)
+The lengths of intervals $A$ and $B$ are
+$L_{A}=L_{B}=\cos^{-1}\left(\cos(\Delta\phi-\delta\phi)\cosh\left(2\tanh^{-1}\left(\tan\frac{\delta\phi}{2}\right)\right)\right)\,.$ (84)
+Taking the limit $\delta\phi\to 0$ and expanding formula (139) in a series, we obtain the BSSA in dS space:
+$S^{\prime\prime}(\Delta\phi)+\frac{1}{\sin\Delta\phi}S^{\prime}(\Delta\phi)\leq 0\,.$ (85)
+It is important to note that the preceding discussion is applicable only to the constant time slice at $t=0$.
+For cases where $t\neq 0$, we will directly use the entanglement entropy results obtained in the text to verify whether the SSA and BSSA are satisfied.
+5.1 The dS/dS holography
 For the dS/dS correspondence, the bulk metric is given by
@@
 $ds^{2}=\ell_{\rm dS}^{2}\sin^{2}w_{c}(-dt^{2}+\cosh^{2}td\phi^{2})\,.$ (87)
-Comparing this with the metric in eq.
-(98), the only difference between the induced metrics for the dS and AdS cases is the replacement of the conformal factor $\sinh^{2}w$ with $\sin^{2}w$.
-As a result, the subsequent calculations are identical to those in the AdS case, except for replacing $\sinh^{2}w_{c}$ with $\sin^{2}w_{c}$.
-Consequently, the effective radius is given by $r_{\rm eff}=\ell_{\rm dS}\sin w_{c}\cosh t_{0}\sin\phi_{a}$.
-For dS/dS, the entanglement entropy expression for a general spatial interval is given by eq.
-(50).
-In this case, the effective radius must be less than $\ell_{\rm dS}$, which implies
+In the Euclidean signature, after rescaling the time coordinate, the boundary metric can be written as
+$ds^{2}=\ell_{\rm dS}^{2}\frac{\sin^{2}w_{c}}{\cosh^{2}\tilde{t}}\left(d\tilde{t}^{2}+d\phi^{2}\right)\,,$ (88)
+where $\tilde{t}$ is related to $t$ by the relation $\tilde{t}=\ln\frac{\cot\frac{t}{2}+1}{\cot\frac{t}{2}-1}$.
+This metric is conformally equivalent to the cylinder.
+As we have done in the last section, we can map it to the complex plane through the following transformations
+$\tilde{t}\rightarrow\frac{1}{2}\ln(T^{2}+X^{2})\,,\quad\phi\rightarrow\frac{1}{2i}\ln\left(\frac{T+iX}{T-iX}\right)\,.$ (89)
+and the resulting conformal flat metric is
+$ds^{2}=\ell_{\rm dS}^{2}\sin^{2}w_{c}\frac{4\left(dT^{2}+dX^{2}\right)}{\left(1+T^{2}+X^{2}\right)^{2}}\,.$ (90)
+The two endpoints of the subregion $A$ in the $(T,X)$ coordinates are
+$(T_{1},X_{1})=(e^{\tilde{t}_{0}}\cos\phi_{a},-e^{\tilde{t}_{0}}\sin\phi_{a}),\quad(T_{2},X_{2})=(e^{\tilde{t}_{0}}\cos\phi_{a},e^{\tilde{t}_{0}}\sin\phi_{a})\,.$ (91)
+Using the CHM map,
+$T\rightarrow\frac{e^{\tilde{t}_{0}}\sin\phi_{a}\sin\Theta\sin\Phi}{1+\sin\Theta\cos\Phi}+e^{\tilde{t}_{0}}\cos\phi_{a}\,,\quad X\rightarrow\frac{e^{\tilde{t}_{0}}\sin\phi_{a}\cos\Theta}{1+\sin\Theta\cos\Phi}\,,$ (92)
+this complex plane can be mapped onto the sphere, yielding the conformal spherical metric
+$ds^{2}=\frac{4e^{2\tilde{t}_{0}}\ell_{\rm dS}^{2}\sin^{2}w_{c}\sin^{2}\phi_{a}\left(d\Theta^{2}+\sin^{2}\Theta d\Phi^{2}\right)}{\left(1+\sin\Theta\cos\Phi+e^{2\tilde{t}_{0}}\sin\Theta\cos(\Phi-2\phi_{a})+e^{2\tilde{t}_{0}}\right)^{2}}\,.$ (93)
+It is easy to verify that the conformal factor at the endpoints $\Theta=0,\pi$ is the same constant, which can be absorbed into $r_{\rm eff}$.
+Therefore, the effective radius is $r_{\text{eff}}=\ell_{\rm dS}\sin w_{c}\cos t_{0}\sin\phi_{a}$.
+Finally, in Lorentz signature, the effective radius is given by
+$r_{\text{eff}}=\ell_{\rm dS}\sin w_{c}\cosh t_{0}\sin\phi_{a}\,.$ (94)
+As mentioned in Section 3.2, we require that the effective spherical radius be less than $\ell_{\rm dS}$, i.e.,
 $\sin w_{c}\cosh t_{0}\sin\phi_{a}<1.$ (95)
-Finally, we can obtain the entanglement entropy of subregion $A$ as
-$S_{A}=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{1}{(\sin w_{c}\cosh t_{0}\sin\phi_{a})^{2}}-1}}\right)\,.$ (112)
-For the specific case $t=0$, we can derive the relation
-$\frac{1}{(\sin w_{c}\sin\phi_{a})^{2}}-1=\cot^{2}\phi_{a}+\frac{\cot^{2}w_{c}}{\sin^{2}\phi_{a}}\,.$ (113)
-Substituting this into (112), the entanglement entropy at $t_{0}=0$ becomes
-$S_{A}=\frac{c_{\rm dS}}{3}\mathrm{arccot}\left(\sqrt{\cot^{2}\phi_{a}+\frac{\cot^{2}w_{c}}{\sin^{2}\phi_{a}}}\right)\,,$ (114)
+In fact, this ensures that the geodesic length connecting the endpoints of the selected subregion in dS spacetime is real Geng:2020kxh ; Dong:2018cuv .
+Finally, we obtain the entanglement entropy of subregion $A$ as
+$S_{A}=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\rm eff}^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{1}{(\sin w_{c}\cosh t_{0}\sin\phi_{a})^{2}}-1}}\right)\,.$ (96)
+For the specific case $t=0$, the entanglement entropy at $t_{0}=0$ becomes
+$S_{A}=\frac{c_{\rm dS}}{3}\cot^{-1}\left(\sqrt{\cot^{2}\phi_{a}+\frac{\cot^{2}w_{c}}{\sin^{2}\phi_{a}}}\right)\,,$ (97)
 which is consistent with the holographic calculation of the area of the RT surface in ref.
@@
 Gorbenko:2018oov .
-The next task is to determine whether the system’s entanglement entropy satisfies the strong subadditivity inequality.
-To achieve this, we compute the second-order derivative of the entanglement entropy with respect to the subsystem scale $\Delta\phi$, where $\Delta\phi=2\phi_{a}$.
-The result is
-$\partial^{2}_{\Delta\phi}S_{A}(\Delta\phi)=-\frac{c_{\rm dS}}{12\sin^{2}\frac{\Delta\phi}{2}}\frac{\frac{1}{\sin^{2}w_{c}\cosh^{2}t_{0}}-1}{\left(\frac{1}{\sin^{2}\frac{\Delta\phi}{2}\sin^{2}w_{c}\cosh^{2}t_{0}}-1\right)^{3/2}}.$ (115)
-We find that when $\sin w_{c}\cosh t_{0}>1$, the entanglement entropy always violates the strong subadditivity inequality.
-However, when $\sin w_{c}\cosh t_{0}<1$, for any $\Delta\phi$, the condition $r_{\rm eff}<\ell_{\rm dS}$ is always satisfied, and the entanglement entropy satisfies the strong subadditivity inequality, as illustrated in figure 12(a) and figure 12(b).
-(a) $w_{c}=\frac{\pi}{2}$.
-(b) $t_{0}=0$.
-Figure 12: The second-order derivative of the entanglement entropy.
-12(a) we take $w_{c}=\frac{\pi}{2}$, with $r_{\text{eff}}=\ell_{\rm dS}\cosh t_{0}\sin\frac{\Delta\phi}{2}$.
-In this case, we need to always ensure that $\cosh t_{0}\sin\frac{\Delta\phi}{2}<1$.
-12(b) we take $t_{0}=0$, with $r_{\text{eff}}=\ell_{\rm dS}\sin w_{c}\sin\frac{\Delta\phi}{2}<\ell_{\rm dS}$.
-For the two cases described above, the boundary positions are illustrated in figure.
-13.
-The orange region corresponds to $\sin w_{c}\cosh t_{0}>1$, where the dual field theory is nonlocal.
-The blue region corresponds to $\sin w_{c}\cosh t_{0}<1$, where the dual field theory satisfies the strong subadditivity inequality.
-However, it can be proven that in this case, the entanglement entropy does not satisfy the boosted strong subadditivity inequality.
-In the dS2 background, the infinitesimal form of the boosted strong subadditivity inequality is given by Geng:2019ruz
-$\text{BSSA}(\Delta\phi)=\partial^{2}_{\Delta\phi}S_{A}(\Delta\phi)+\frac{1}{\sin\Delta\phi}\partial_{\Delta\phi}S_{A}(\Delta\phi)\leq 0.$ (116)
-Figure 13: The Penrose diagram of dS/dS correspondence.
+The subsequent task is to determine whether the system’s entanglement entropy satisfies both SSA and BSSA.
+We will discuss the two cases of $t_{0}=0$ and $t_{0}\neq 0$ separately.
+5.1.1 The case of $t_{0}=0$
+The second-order derivative of the entanglement entropy with respect to the subsystem scale $\Delta\phi$ ($\Delta\phi=2\phi_{a}$) is
+$\partial^{2}_{\Delta\phi}S_{A}(\Delta\phi)=-\frac{c_{dS}}{12\sin^{2}\frac{\Delta\phi}{2}}\frac{\frac{1}{\sin^{2}w_{c}}-1}{\left(\frac{1}{\sin^{2}\frac{\Delta\phi}{2}\sin^{2}w_{c}}-1\right)^{3/2}}<0.$ (98)
+It can be observed that the SSA is manifestly satisfied, as illustrated in figure 13(a).
+However, the BSSA is violated in this case.
+Substituting Eq.
+(97) into the left-hand side of the BSSA (85), we obtain
+$\text{BSSA}(\Delta\phi)=\frac{c_{dS}}{48}\frac{\sin^{2}\Delta\phi}{\sin^{6}\frac{\Delta\phi}{2}\sin^{2}w_{c}\left(\frac{1}{\sin^{2}\frac{\Delta\phi}{2}\sin^{2}w_{c}}-1\right)^{\frac{3}{2}}}>0,$ (99)
+where $\text{BSSA}(\Delta\phi)$ represents $S^{\prime\prime}(\Delta\phi)+\frac{1}{\sin\Delta\phi}S^{\prime}(\Delta\phi)$.
+This indicates that the dual field theory continues to exhibit non-locality.
+Notably, at the ultraviolet boundary where $w_{c}=\frac{\pi}{2}$, there is a significant violation of the BSSA [^6] In the $T\bar{T}$-deformed AdS/CFT, the BSSA violation also occurs due to the influence of the $T\bar{T}$ deformation Aharony:2023dod .
+However, as the finite truncation $r_{c}$ approaches the boundary, the BSSA tends to saturate, which contrasts with the behavior in the dS spacetime background., as shown by the purple line in figure 13(b).
+(a) The SSA is satisfied.
+(b) The BSSA is violated.
+Figure 13: The SSA and BSSA in dS/dS holography at $t_{0}=0$.
+13(a) The SSA is satisfied at $t_{0}=0$.
+When $w_{c}=\frac{\pi}{2}$, the SSA is saturated.
+13(b) The BSSA is violated at $t_{0}=0$.
+When $w_{c}=\frac{\pi}{2}$, the violation of the BSSA inequality is obvious.
+5.1.2 The case of $t_{0}\neq 0$
+In the case of $t_{0}=0$, $\sin w_{c}\cosh t_{0}<1$, but for $t_{0}\neq 0$, the value of $\sin w_{c}\cosh t_{0}$ can be either greater than 1 or less than 1.
+Since the effective spherical radius $r_{\text{eff}}=\ell_{dS}\sin w_{c}\cosh t_{0}\sin\frac{\Delta\phi}{2}$ is required to be less than $\ell_{dS}$, we must consider both cases, $\sin w_{c}\cosh t_{0}<1$ and $\sin w_{c}\cosh t_{0}>1$, separately.
+(a) The SSA is satisfied.
+(b) The BSSA is violated.
+Figure 14: The SSA and BSSA in dS/dS holography at $\sin w_{c}\cosh t_{0}<1$.
+14(a) The SSA is satisfied at $\sin w_{c}\cosh t_{0}<1$.
+The entanglement entropy, as a function of the subinterval scale $\Delta\phi$, is a concave function, ensuring that subadditivity is always satisfied.
+14(b) The BSSA is violated at $\sin w_{c}\cosh t_{0}<1$.
+To maintain $\sin w_{c}\cosh t_{0}<1$, we set $t_{0}$ to 0.01, 0.03, 0.05, 0.07, and 0.09, respectively, and take $\delta\phi=0.01\Delta\phi$.
+When $\sin w_{c}\cosh t_{0}<1$, the condition $r_{\text{eff}}<\ell_{dS}$ is always satisfied for any $\Delta\phi$.
+Since $t\neq 0$, we directly use the entanglement entropy to verify whether SSA and BSSA are satisfied.
+As shown in figure 14(a), when $\sin w_{c}\cosh t_{0}<1$, the entanglement entropy is a concave function, and the concavity of the function indicates that $\partial^{2}_{\Delta\phi}S(\Delta\phi,w_{c},t_{0})<0$, confirming that the SSA is satisfied.
+However, upon considering the boost transformation, we find that the BSSA is consistently violated, as shown in Figure 14(b).
+(a) The SSA is violated.
+(b) The BSSA is violated.
+Figure 15: The SSA and BSSA in dS/dS holography at $\sin w_{c}\cosh t_{0}>1$.
+15(a) The SSA is violated at $\sin w_{c}\cosh t_{0}>1$.
+The entanglement entropy is a convex function, ensuring that subadditivity is always violated.
+15(b) The BSSA is violated at $\sin w_{c}\cosh t_{0}>1$.
+We set $t_{0}$ to 1.2, 1.4, 1.6, 1.8, and 2.0, respectively, and take $\delta\phi=0.1\Delta\phi$.
+When $\sin w_{c}\cosh t_{0}>1$, to ensure that the effective spherical radius is less than $\ell_{\rm dS}$, the value of $\Delta\phi$ has a maximum value $\Delta\phi_{\text{max}}$ such that $\sin w_{c}\cosh t_{0}\sin\frac{\Delta\phi_{\text{max}}}{2}=1$.
+As shown in figure 15(a), in the interval $[0,\Delta\phi_{\text{max}}]$, the entanglement entropy as a function of $\Delta\phi$ is consistently a convex function, which ensures that the subadditivity is always violated.
+Similarly, considering the boost transformation, it can be observed that the BSSA is still violated when $\sin w_{c}\cosh t_{0}$ is greater than 1, as shown in figure 15(b).
+Figure 16: The Penrose diagram of dS/dS holography.
 The black and blue dashed lines represent the codimension-one hypersurfaces at $t=0$ and $w=\frac{\pi}{2}$, respectively.
 The black curve represents the spacelike hypersurface at $t=t_{0}$, and the orange and blue curves represent the dS slices at $w=w_{c}$, where the dual field theory is located.
-The intersection of these two hypersurfaces corresponds to the critical case that saturates the strong subadditivity inequality, where $\sin w_{c}\cosh t_{0}=1$.
-By substituting eq. (114) into the left-hand side of the boosted strong subadditivity inequality (116), we obtain
-$\text{BSSA}(\Delta\phi)=\frac{c_{\rm dS}}{48}\frac{\sin^{2}\Delta\phi}{\sin^{6}\frac{\Delta\phi}{2}\sin^{2}w_{c}\left(\frac{1}{\sin^{2}\frac{\Delta\phi}{2}\sin^{2}w_{c}}-1\right)^{\frac{3}{2}}}>0,$ (117)
-which indicates that, in this case, the dual field theory remains nonlocal.
-In particular, at the ultraviolet boundary where $w_{c}=\frac{\pi}{2}$, there is a significant violation of the boosted strong subadditivity inequality, as indicated by the purple line in figure 14.
-Figure 14: The boosted strong subadditivity inequality violate in the dS/dS holography.
+The intersection of these two hypersurfaces corresponds to the critical case that saturates the SSA, where $\sin w_{c}\cosh t_{0}=1$.
+For the two cases described above, the critical positions are illustrated in figure 16.
+The orange region corresponds to $\sin w_{c}\cosh t_{0}>1$, where the dual field theory is non-local.
+The blue region corresponds to $\sin w_{c}\cosh t_{0}<1$, where the dual field theory satisfies the SSA, but it still violates the BSSA.
+This demonstrates that the dual field theory still exhibits non-locality in the blue region in figure 16.
 5.2 The half-dS holography
-By introducing a timelike boundary $\Theta=\Theta_{0}$ in dS global spacetime, the half-dS holographic model was obtained in ref.
-Kawamoto:2023nki .
-The dual field theory resides on the timelike boundary, and the bulk spacetime is confined to the region where $0<\Theta<\Theta_{0}$.
-The induced metric on the boundary is given by
+For the half-dS holography, the induced metric on the boundary is given by
 $ds^{2}=\ell_{\rm dS}^{2}(-dt^{2}+\cosh^{2}t\sin^{2}\theta_{0}d\phi^{2}).$ (100)
@@
 $ds^{2}=\ell_{\rm dS}^{2}\frac{\sin^{2}\theta_{0}}{\cosh^{2}\left(\tilde{t}\sin\theta_{0}\right)}\left(d\tilde{t}^{2}+d\phi^{2}\right),$ (101)
-where the coordinate transformation between $\tilde{t}$ and $t$ is
-$\tilde{t}=\frac{1}{\sin\Theta_{0}}\ln\frac{\cot\frac{t}{2}+1}{\cot\frac{t}{2}-1}.$ (120)
-Mapping from the cylinder to the plane as in eq. (101), the metric transforms into a conformally flat form
+where the coordinate transformation between $\tilde{t}$ and $t$ is $\tilde{t}=\frac{1}{\sin\theta_{0}}\ln\frac{\cot\frac{t}{2}+1}{\cot\frac{t}{2}-1}$.
+Mapping from the cylinder to the plane as in eq.
+(89), the metric transforms into a conformally flat form
 $ds^{2}=\frac{\ell_{\rm dS}^{2}\sin^{2}\theta_{0}}{\cosh^{2}\left(\ln(\sqrt{T^{2}+X^{2}})\sin\theta_{0}\right)}\frac{dT^{2}+dX^{2}}{T^{2}+X^{2}}.$ (102)
@@
 $r_{\text{eff}}=\ell_{\rm dS}\cosh t_{0}\sin\theta_{0}\sin\phi_{a}.$ (107)
+In order to ensure the existence of spacelike geodesics, the ref.
+Kawamoto:2023nki requires  [^7] Equation (3.6) of ref.
+Kawamoto:2023nki with the substitution $\phi_{1}-\phi_{2}=2\phi_{a}=\Delta\phi$.
+$\cos^{2}\theta_{0}+\sin^{2}\theta_{0}\cos(\phi_{1}-\phi_{2})\leq 1-\frac{2}{\cosh^{2}t_{0}}.$ (108)
+After simplification and deformation, it can be observed that this ensures that the effective radius of the sphere, $r_{\text{eff}}=\ell_{\rm dS}\cosh t_{0}\sin\theta_{0}\sin\frac{\Delta\phi}{2}$, is less than or equal to $\ell_{\rm dS}$.
 Therefore, the entanglement entropy is
-$S=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{1}{(\cosh t_{0}\sin\Theta_{0}\sin\Phi_{a})^{2}}-1}}\right).$ (127)
+$S=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{1}{(\cosh t_{0}\sin\theta_{0}\sin\phi_{a})^{2}}-1}}\right).$ (109)
 In the case of $t_{0}=0$ and $\theta_{0}=\frac{\pi}{2}$, the radius in the spherical partition function is $r_{\text{eff}}=\ell_{\rm dS}\sin\frac{\Delta\phi}{2}$, where $\Delta\phi=|-\phi_{a}-\phi_{a}|=2\phi_{a}$.
 The entanglement entropy is
-$S=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{1}{\sin^{2}\frac{\Delta\Phi}{2}}-1}}\right)=\frac{c_{\rm dS}}{6}\Delta\Phi.$ (128)
+$S=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}-1}}\right)=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{1}{\sin^{2}\frac{\Delta\phi}{2}}-1}}\right)=\frac{c_{\rm dS}}{6}\Delta\phi.$ (110)
 By replacing the field theory parameter $c_{\rm dS}$ with the gravitational parameter $\frac{3\ell_{\rm dS}}{2G}$, we find that the above result matches the holographic calculation presented in ref.
-Kawamoto:2023nki , confirming the validity of the half-dS holography.
-Additionally, the requirement that the effective radius of the sphere be smaller than the cosmological horizon aligns precisely with the condition given in eq.
-(3.6) of ref.
-Kawamoto:2023nki to ensure the existence of spacelike geodesics:
-$r_{\text{eff}}=\ell_{\rm dS}\cosh t_{0}\sin\Theta_{0}\sin\frac{\Delta\Phi}{2}\leq\ell_{\rm dS}.$ (129)
-In ref.
-Kawamoto:2023nki , the dual field theory in the half-dS holography was proposed to exhibit non-locality.
-It was concluded that when $\cosh t_{0}\leq\frac{1}{\sin\Theta_{0}}$, the entanglement entropy satisfies the strong subadditivity inequality, whereas when $\cosh t_{0}>\frac{1}{\sin\Theta_{0}}$, the strong subadditivity inequality is violated.
-As shown in figure 15, the orange line segment represents the region where the strong subadditivity inequality is violated, while the blue line segment represents the region where the strong subadditivity inequality is satisfied.
-Next, we demonstrate that even in the blue region where the strong subadditivity inequality is satisfied, the boosted strong subadditivity inequality is still violated.
-Figure 15: The half-dS holography.
-The dual field theory resides on a timelike boundary at $\Theta=\Theta_{0}$, while the bulk spacetime covers the region $0<\Theta<\Theta_{0}$.
-Here we set $\cos T=\frac{1}{\cosh t}$.
+Kawamoto:2023nki , confirming the validity of half-dS holography.
+By comparing Eqs.
+(96) and (109), we find that the entanglement entropy results for spatial intervals, calculated independently in dS/dS holography and half-dS holography, reveal a close analogy between the two frameworks.
+Therefore, the discussion of SSA aligns with that in dS/dS holography.
+However, due to the difference in boundary-induced metrics, the BSSA receives corresponding corrections.
+In half-dS holography, the infinitesimal form of the BSSA is
+$\text{BSSA}(\Delta\tilde{\phi})=\sin^{2}\theta_{0}\left(\partial^{2}_{\Delta\tilde{\phi}}S(\Delta\tilde{\phi})+\frac{1}{\sin\Delta\tilde{\phi}}\partial_{\Delta\tilde{\phi}}S(\Delta\tilde{\phi})\right)\leq 0,$ (111)
+where $\Delta\tilde{\phi}=\sin\theta_{0}\Delta\phi$. Substituting the entanglement entropy result into the above expression, we obtain
+$\text{BSSA}(\Delta\phi)=\frac{c_{dS}}{48}\frac{\sin^{2}(\Delta\phi\sin\theta_{0})}{\sin^{6}(\frac{1}{2}\Delta\phi\sin\theta_{0})(\frac{1}{\sin^{2}\theta_{0}\sin^{2}(\frac{1}{2}\Delta\phi\sin\theta_{0})}-1)^{\frac{3}{2}}}>0.$ (112)
+The result is illustrated in figure 17.
+Compared with figure 13(b), it is evident that figure 17 exhibits only slight differences.
+The primary behavior of the function remains nearly identical, and a clear deviation in behavior is observed at $\theta_{0}=\frac{\pi}{2}$, indicating that the boundary dual field theory in half-dS holography continues to exhibit non-locality.
+Figure 17: The boosted strong subadditivity is violated in half-dS holography.
+Similarly, there is a critical position $\sin\theta_{0}\cosh t_{0}=1$ in half-dS holography, as indicated by the red dot in figure 18.
+As depicted in figure 18, the orange line segment represents $\sin\theta_{0}\cosh t_{0}>1$, where the SSA is violated, while the blue line segment represents $\sin\theta_{0}\cosh t_{0}<1$, where the SSA is satisfied, but the BSSA is still violated.
+Figure 18: Half-de Sitter holography.
+The dual field theory resides on a timelike boundary at $\theta=\theta_{0}$.
 The blue intervals indicate $\cosh t\leq\frac{1}{\sin\theta_{0}}$, i.e., $\theta_{0}-\frac{\pi}{2}\leq T\leq\frac{\pi}{2}-\theta_{0}$.
-By observing eqs.
-(112) and (127), we find that the results from calculating the spatial interval entanglement entropy in the dS/dS holography and half-dS holography are similar.
-However, due to the difference in boundary-induced metrics, the boosted strong subadditivity inequality receives corresponding corrections.
-In the half-dS holography, the infinitesimal form of the boosted strong subadditivity inequality is
-$\text{BSSA}(\Delta\tilde{\Phi})=\sin^{2}\Theta_{0}\left(\partial^{2}_{\Delta\tilde{\Phi}}S(\Delta\tilde{\Phi})+\frac{1}{\sin\Delta\tilde{\Phi}}\partial_{\Delta\tilde{\Phi}}S(\Delta\tilde{\Phi})\right)\leq 0,$ (130)
-where $\Delta\tilde{\Phi}=\sin\Theta_{0}\Delta\Phi$. Substituting the result for the entanglement entropy into the above expression, we can get
-$\text{BSSA}(\Delta\phi)=\frac{c_{\rm dS}}{48}\frac{\sin^{2}(\Delta\phi\sin\Theta_{0})}{\sin^{6}(\frac{1}{2}\Delta\phi\sin\Theta_{0})(\frac{1}{\sin^{2}\Theta_{0}\sin^{2}(\frac{1}{2}\Delta\phi\sin\Theta_{0})}-1)^{\frac{3}{2}}}>0.$ (131)
-The result is illustrated in figure 16, which indicates that, in the half-dS holography, the boundary field theory still does not satisfy the boosted strong subadditivity inequality.
-Figure 16: The boosted strong subadditivity inequality violates in the half-dS holography.
-We have calculated the entanglement entropy of subsystem $A$ using the generalized spherical partition function method within the dS/dS correspondence and half-dS holography frameworks.
-Furthermore, we have reaffirmed the conclusions from Section 3.2 by applying the boosted strong subadditivity inequality and the strong subadditivity inequality: in dS3 spacetime, when the effective radius of the sphere is smaller than the cosmological horizon, the dual field theory exhibits non-locality.
+We have computed the entanglement entropy of subsystem $A$ using the generalized spherical replica method within the dS/dS holography and half-dS holography frameworks.
+By verifying the SSA and BSSA, we have found that the entanglement entropy exhibits similar behavior in these two holographic models and that the boundary dual field theory exhibits non-locality.
+Comparing figures 16 and 18, we observe the same critical point, namely, when the region where the dual field theory is located crosses the cosmological horizon in three-dimensional dS3 spacetime, the entanglement entropy violates the SSA, and even within the cosmological horizon, the entanglement entropy still violates the BSSA.
 6 Conclusion and Discussion
-This paper investigated the effects of the $T\bar{T}$ deformation on various holographic dualities, including AdS/CFT, dS/CFT, and dS/dS, by analyzing the entanglement entropy of the boundary field theory.
-When the boundary is spherical, and the subsystem is a semicircular arc, the entanglement entropy can be derived from the classical partition function on the $n$-replica manifold, with the partition function governed by the trace flow equation.
-To study the entanglement entropy in the dS case, we first derived the trace flow equation for the dS/CFT by applying a double Wick rotation to the AdS case.
-Then, we considered a simple scenario where the subsystem is a semicircular arc on the sphere, which corresponds to global coordinates in the dS/CFT, as well as the $t=0$ moment in both AdS/dS and dS/dS.
-Our findings are as follows:
-  • In dS/CFT, the entanglement entropy is complex, indicating that the dual field theory is non-unitary.
-  • In dS/dS, the entanglement entropy is real but violates the strong subadditivity inequality, suggesting that the dual field theory is unitary but non-local.
-We then examined a more complicated scenario where the boundary-induced metric is time-dependent, covering AdS/dS, dS/dS, and half-dS holography.
-In these cases, since the expressions for the strong subadditivity inequality and the boosted strong subadditivity inequality depend on the specific form of the metric, and the boundary-induced metric is time-dependent, the locality of the boundary field theory may vary across different time periods, as discussed in ref.
-Kawamoto:2023nki .
-To address this, we further generalized the replica method on the sphere to accommodate more general manifolds and subsystems.
-We first verified this method in the global coordinates of AdS/CFT, dS/CFT, and the planar BTZ case, where the results are consistent with existing perturbative calculations Chen:2018eqk ; Chen:2023eic .
-We then applies it to the cases where the boundary-induced metrics are time-dependent.
-Our results show that while the subsystem’s entanglement entropy may satisfy the strong subadditivity inequality at certain time periods in dS/dS and half-dS holographies, the boosted strong subadditivity inequality is consistently violated, indicating that the dual field theories remain non-local throughout.
+This paper systematically investigated the impact of the $T\bar{T}$ deformation on various holographic dualities, including AdS/CFT, dS/CFT, dS/dS, and half-dS holography, by analyzing the entanglement entropy of the single general spatial subinterval in the boundary dual field theory.
+When the boundary is spherical, and the subsystem is a semicircular arc, the entanglement entropy can be derived from the spherical partition function on the $n$-replica manifold.
+For general boundary manifolds, we first mapped it onto a conformal spherical manifold via CHM mapping.
+Then, we calculated the entanglement entropy using the corresponding effective spherical radius, and derived the correction after $T\bar{T}$ deformation in both the dS Poincaré spacetime and global dS spacetime.
+In addition, we found that the dual field theory exhibits notable similarities in properties between the dS/dS holography and half-dS holography models.
+Specifically, the dual field theory in both cases exhibited non-local characteristics.
+When the dual field theory lies within the cosmological horizon, it satisfies strong subadditivity but violates boosted strong subadditivity.
+Conversely, when the dual field theory lies outside the cosmological horizon, both strong subadditivity and boosted strong subadditivity are violated.
 Acknowledgements.
-We would like to thank Shan-Ming Ruan for reading our draft and providing comments.
+We would like to thank Shan-Ming Ruan, Yuan Sun for reading our draft and providing comments.
 We also thank Yang He, and Shan-Ping Wu for their helpful discussion.
@@
 20240101326JC).
-Appendix A The gravity calculation
-A.1 The dS/Poincaré case
+7 Data Availability Statement
+The data that support the findings of this article are openly available ChangHe2025Data
+Appendix A Integration over the endpoints of the entanglement region
+For a two-dimensional spherical manifold $\mathcal{M}$, on the slice where $\phi=\text{constant}$, we select $\theta\in[0,\pi]$ as the subregion $A$, the entanglement entropy $S_{A}$ can be obtained through the $C$-function.
+Differentiating $S_{A}$ with respect to $r$, we have
+$r\frac{d}{dr}S_{A}(r)$ $=r\frac{d}{dr}\lim_{n\rightarrow 1}(1-n\partial_{n})\ln Z_{n}=\lim_{n\rightarrow 1}(1-n\partial_{n})r\frac{d}{dr}\ln Z_{n}$
+$=-\lim_{n\rightarrow 1}(1-n\partial_{n})2\pi nr^{2}\int_{0}^{\pi}\sin\theta d\theta\,\text{tr}\,T$
+$=2\pi r^{2}\lim_{n\rightarrow 1}n^{2}\int_{0}^{\pi}\sin\theta d\theta\,\partial_{n}\text{tr}\,T.$ (113)
+Solving the trace flow and conservation equations without assuming spherical symmetry must determine the energy-momentum tensor $T_{ab}$.
+For the $n$-replica sphere (27), the solution for the energy-momentum tensor is known Donnelly:2018bef and is given by
+$T^{\theta}_{\theta}$ $=\frac{1}{\pi\lambda}\left(1-\sqrt{1+\frac{\lambda c}{12r^{2}}+\frac{\lambda c}{12r^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}\right),$
+$T^{\phi}_{\phi}$ $=\frac{1}{\pi\lambda}\left(1-\frac{1+\frac{\lambda c}{12r^{2}}}{\sqrt{1+\frac{\lambda c}{12r^{2}}+\frac{\lambda c}{12r^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\theta}}}\right).$ (114)
+Substituting this into the integral in eq. (113), in the limit $n\rightarrow 1$, the integrand becomes
+$n\partial_{n}\text{tr}\,T=\frac{c}{3}\frac{\lambda c}{48\pi}\frac{1-n^{2}}{n^{4}r^{4}\sin^{4}\theta}\frac{1}{\left[1+\frac{\lambda c}{12r^{2}}\left(1+\frac{2(1-n)}{\sin^{2}\theta}\right)\right]^{3/2}},$ (115)
+which tends to zero for $\theta\neq 0,\pi$.
+However, near the branch points, the integrand has a non-zero contribution.
+Thus, the main contribution to the integral in eq.
+(113) comes from small neighborhoods near the branch points
+$rS^{\prime}(r)=2\pi r^{2}\lim_{n\rightarrow 1}\left(\int_{\theta\in B(0,\epsilon)}+\int_{\theta\in B(\pi,\epsilon)}\right)\sin\theta d\theta\,\partial_{n}\text{tr}\,T,$ (116)
+where $B(p,\epsilon)$ is a small neighborhood centered at $p$ with radius $\epsilon$.
+A.1 plane
+For a plane manifold $\mathcal{M}$, we first map $\mathcal{M}$ to $\mathbb{S}^{2}$ by CHM map, ensuring that the image of the subsystem $A$ under this mapping is a meridian on $\mathbb{S}^{2}$.
+The mapped metric is conformally equivalent to the sphere, denoted as $\Omega\mathbb{S}^{2}$.
+After applying the $n$-replica trick, we assume the metric takes the form [^8] We use capital Greek letters to represent the angular coordinates of the conformal sphere $\Omega\mathbb{S}^{2}$
+$ds^{2}=r^{2}\Omega^{2}(\Theta,\Phi)\left(d\Theta^{2}+n^{2}\sin^{2}\Theta\,d\Phi^{2}\right),$ (117)
+where the conformal factor $\Omega(\Theta,\Phi)$ is constant at the branch points $\Theta=0$ and $\Theta=\pi$, denoted as $\Omega_{0}$ and $\Omega_{\pi}$, respectively.
+For this metric, eqs.
+(29) and (30) remain valid.
+Thus, we have
+$r\frac{d}{dr}S(r)$ $=-\lim_{n\rightarrow 1}(1-n\partial_{n})nr^{2}\int_{\Omega\mathbb{S}_{n}^{2}}\Omega^{2}(\Theta,\Phi)\sin\Theta\,d\Theta d\Phi\,\text{tr}\,T$
+$=r^{2}\lim_{n\rightarrow 1}n^{2}\int_{\Omega\mathbb{S}_{n}^{2}}\Omega^{2}(\Theta,\Phi)\sin\Theta\,d\Theta d\Phi\,\partial_{n}\text{tr}\,T.$ (118)
+Since the conformal factor $\Omega(\Theta,\Phi)$ depends on the azimuthal angle $\Phi$, $T_{ab}$ cannot be assumed to depend solely on $\Theta$.
+Therefore, the $\Phi$ coordinate is not integrated.
+As discussed in ref.
+Lewkowycz:2019xse , for the partition function on the sphere (113), in the limit $n\rightarrow 1$, the main contribution to the integral (113) comes from a small neighborhood near the branch points, with radius $\epsilon\sim\sqrt{n-1}$.
+This property is expected to hold for any manifold $\mathcal{M}$, so the integral (118) is also performed only in small neighborhoods near the branch points.
+In these regions, the conformal factor $\Omega$ is constant, and the metric (117) can be approximated by the $n$-replica sphere with radius $r\Omega$.
+As a result, the conformal transformation can be regarded as an identity transformation999 In standard CFT calculations, a conformal transformation is often applied to map $\Omega\mathbb{S}^{2}$ to the spherical metric, and conformal symmetry ensures that the results for $\Omega\mathbb{S}^{2}$ and $\mathbb{S}^{2}$ differ only by a function of the conformal factor.
+However, for the $T\bar{T}$-deformed field theory, conformal symmetry is absent, and the relationship between quantities before and after the conformal transformation is not determined..
+Therefore, partition functions on any manifold $\mathcal{M}$ can be calculated using the results on $\mathbb{S}^{2}$ for non-conformal field theories.
+Finally, we can get
+$rS^{\prime}(r)=\lim_{n\rightarrow 1}2\pi(r\Omega)^{2}\left(\int_{0}^{\Theta_{0}}+\int_{\pi-\Theta_{0}}^{\pi}\right)\sin\Theta\,d\Theta\,\partial_{n}\text{tr}\,T,$ (119)
+where $\text{tr}\,T$ corresponds to the solution on the sphere with radius $r\Omega$.
+When Poincaré coordinates are chosen in EAdS3, the holographic boundary corresponds to a two-dimensional complex plane with the metric
+$ds_{2}^{2}=\frac{\ell_{\rm AdS}^{2}}{z_{c}^{2}}(dT^{2}+dX^{2}).$ (120)
+The subregion $A$ is defined by $T=0$ and $X\in[-X_{a},X_{a}]$.
+To compute the entanglement entropy for $A$, the complex plane is mapped onto a two-dimensional sphere such that the image of $A$ corresponds to the meridian on the sphere at $\Phi=0$.
+This mapping, known as the CHM mapping, is given by
+$X=\frac{X_{a}\cos\Theta}{1+\sin\Theta\cos\Phi},\quad T=\frac{X_{a}\sin\Theta\sin\Phi}{1+\sin\Theta\cos\Phi}.$ (121)
+Under this transformation, the boundary plane in $(\Theta,\Phi)$ coordinates has the metric
+$ds_{2}^{2}=\frac{\ell_{\rm AdS}^{2}X_{a}^{2}}{z_{c}^{2}(1+\cos\Phi\sin\Theta)^{2}}(d\Theta^{2}+\sin^{2}\Theta\,d\Phi^{2}),$ (122)
+which is conformally equivalent to a sphere with an effective radius $r_{\mathrm{eff}}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}$, and the conformal factor $\Omega(\Theta,\Phi)=(1+\cos\Phi\sin\Theta)^{-1}$.
+As previously noted, when computing the partition function on a general manifold $\mathcal{M}$, the primary contribution to the integral (113) comes from an infinitesimal neighborhood near the branch points.
+In these regions, the conformal factor $\Omega(\Theta,\Phi)$ can be approximated as 1, implying that the metric (122) effectively describes a sphere with radius
+$r_{\mathrm{eff}}=\frac{\ell_{\rm AdS}X_{a}}{z_{c}}\,.$ (123)
+Thus, within the infinitesimal neighborhoods around the endpoints, transforming (120) to the sphere involves only a conformal transformation with a conformal factor of 1, effectively an identity transformation.
+Consequently, the partition function remains the same as that of the sphere, with the radius $r$ replaced by the effective radius $r_{\mathrm{eff}}$.
+Therefore, the solution for the energy-momentum tensor on the $n$-replica sphere is
+$T^{\Theta}_{\Theta}$ $=\frac{1}{\pi\lambda_{\rm AdS}}\left(1-\sqrt{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\Theta}}\right),$
+$T^{\Phi}_{\Phi}$ $=\frac{1}{\pi\lambda_{\rm AdS}}\left(1-\frac{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}}{\sqrt{1+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}+\frac{\lambda_{\rm AdS}c_{\rm AdS}}{12r_{\rm eff}^{2}}\left(\frac{1}{n^{2}}-1\right)\frac{1}{\sin^{2}\Theta}}}\right)\,.$ (124)
+Substituting this into eq. (113) and performing the integration near the endpoints yields the $C$-function as
+$C(r)=\frac{c_{\rm AdS}}{3}\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\rm eff}^{2}}}}\,.$ (125)
+So we can get the entanglement entropy of the subregion A as
+$S_{A}=\frac{c_{\rm AdS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{z_{c}^{2}}{X_{a}^{2}}}}\right)\simeq\frac{c_{\mathrm{AdS}}}{3}\log\frac{2X_{a}}{z_{c}}+\frac{c_{\mathrm{AdS}}}{12}\frac{z_{c}^{2}}{X_{a}^{2}}.$ (126)
+Appendix B The gravity calculation
+B.1 The dS Poincaré spacetime
 In the dS Poincaré spacetime,
@@
 $Y$ $=\frac{\ell_{\rm dS}}{\eta}T,\quad W=\ell_{\rm dS}\cosh(\log(\frac{\ell_{\rm dS}}{\eta}))-\frac{1}{\eta}\frac{X^{2}+T^{2}}{2}.$
-Finally we can get the geodesic distance is
-$D_{12}=\ell_{\rm dS}\arccos\left(\ell_{\rm dS}^{2}\left(1-\frac{2X_{a}^{2}}{\eta_{c}^{2}}\right)\right)\simeq-2i\ell_{\rm dS}\log\frac{2X_{a}}{\eta_{c}}+i\ell_{\rm dS}\frac{\eta_{c}^{2}}{2X_{a}^{2}}+\pi\ell_{\rm dS}.$ (137)
+Finally we can get that the geodesic distance is
+$D_{12}=\ell_{\rm dS}\cos^{-1}\left(\ell_{\rm dS}^{2}\left(1-\frac{2X_{a}^{2}}{\eta_{c}^{2}}\right)\right)\simeq-2i\ell_{\rm dS}\log\frac{2X_{a}}{\eta_{c}}+i\ell_{\rm dS}\frac{\eta_{c}^{2}}{2X_{a}^{2}}+\pi\ell_{\rm dS}.$ (132)
 So the entanglement entropy of the subregion is
@@
 This is consistent with eq. (52).
-A.2 The global dS case
+B.2 The global dS spacetime
 For the dS global coordinates, the coordinate transformation in embedding spacetime is
@@
 We can get the same result as eq. (4.2.2).
+Appendix C Strong subadditivity and boosted strong subadditivity
+For a unitary and Poincaré-invariant field theory, the strong subadditivity is given by:
+$S(A)+S(B)-S(A\cup B)-S(A\cap B)\geq 0,$ (139)
+where the subsystems $A$ and $B$ are shown in figure 19.
+This inequality can be derived from the monotonicity of relative entropy, which comes from the locality of field theory Witten:2018zxz .
+According to the RT formula, the entanglement entropy is quantified by the area of the extremal surface, and thus the entanglement entropy can be expressed as a function of the subsystem size, such as $S(A)\equiv S(L_{A})$.
+If we set $L_{A\cap B}=L$, $L_{A\cup B}=L+2\delta L$, the lengths of $A$ and $B$ are both $L_{A}=L_{B}=L+\delta L$.
+In the limit $\delta L\to 0$, the strong subadditivity can be expressed as
+$L^{2}S^{\prime\prime}(L)\leq 0,$ (140)
+where $L$ represents the size of the subsystems.
+Figure 19: 19 Strong subadditivity. 19 Boosted strong subadditivity.
+A stronger criterion for proving locality, compared to the strong subadditivity, is the boosted strong subadditivity, which can be derived by applying a boost transformation to subsystems $A$ and $B$, as shown in figure 19.
+We also set $L_{A\cap B}=L$, $L_{A\cup B}=L+2\delta L$.
+In Minkowski spacetime, the lengths of $A$ and $B$ are both $L_{A}=L_{B}=\sqrt{L(L+2\delta L)}$.
+In the limit $\delta L\to 0$, the boosted strong subadditivity can be expressed as
+$S^{\prime\prime}(L)+\frac{S^{\prime}(L)}{L}\leq 0.$ (141)